A Bibliography of Publications in The Computer Journal: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.42

Title word cross-reference

(2n − (2p ± 1)) [Hia04]. $103.95/$129.00/£ [Mah06]. $110.00 [Gaz06]. $129.00 [Pol06]. 2 [LCT01, MM00, WTSS01]. 2 × n [CL04]. 3 [Lev07, MM00]. $59.95/ [Gom06]. $79.95/£ [Azi06, Cas06, Smi06]. $89.95/$99.95/£ [Mor06]. A [AND07]. GF (2^{m}) [ZL03]. K [BT06, KW08, Li03, Mar08a, WB04, XUA00]. L(h, k) [Cal06]. λ [BKTV04]. n [Wu02]. O(1) [XU01]. P [Syr06]. T [Kor05, AND07, Kor06b].

-Ary [AND07, Kor06b, XUA00, Kor05]. -calculus [SW01]. -Colorings [BKTV04]. -Coteries [KW08]. -D [Wu02, WTSS01].

-data [Mar08a]. -dimensional [Li03]. -Labelling [Cal06]. -Maximum [BT06]. -Median [LCT01]. -Nearest [WB04]. -order [AND07].

0-262-07261-0 [Pal06a]. 0-262-16228-8 [Car06]. 0-321-30677-5 [BI06]. 0-387-21017-2 [Jas06]. 0-387-23253-2 [Pol06]. 0-387-23759-3 [Cas06]. 0-387-95569-0 [Smi06]. 0-471-73427-6 [Gaz06]. 0-471-74719-X [HM06]. 0-521-43592-7 [Sch07]. 0-7506-6496-7 [Dud06]. 0-7506-7796-1 [Ref09]. 0-7506-7864-X [Bur06b]. 038723795X [Grü07].

1 [Ano06a]. 1-4020-2268-9 [Lev06b]. 1-4020-7894-3 [AR06]. 1-55860-843-5 [Gom06]. 1-59749-076-8 [Har07]. 1-85233-704-4 [Azi06]. 1-85233-836-9 [Mor06].

2003 [BA04]. 2005 [Ano06a]. 2006 [Mai07]. 21 [Ang05, Par05]. 27.99/US$44.95 [Dud06]. 29.00/$50.00 [Sch07]. 2nd [Rei04].

3-540-20854-2 [Poe06]. 3-540-23342-3 [Sch06]. 3-540-23894-8 [Pal06b]. 3-540-23949-9 [Bur06a]. 3-540-25059-X [Mah06]. 3-540-28583-0 [Vak06]. 30.50/£ [Bur06a]. 30/$60 [Gei09]. 31 [Ano06a]. 32.50/US$58.00 [Vak06]. 32.99/$59.95 [Har07]. 34.55/US$117.00 [Pal06b]. 38.50/$59.95 [Mar09]. 39.95/$49.95 [Bur06a].

4-Bit [YSHP08]. 44.94/£ [Vak06]. 44.95/$49.95/£ [Bur06b]. 45.00/£ [Azi06]. 469pp [Poe06]. 47.95/£ [Har07]. 49.95/£ [Sch06].

50.00/ [Cas06]. 512 [AD07]. 5th [ZSP^{+}04].

64.95/ [Gaz06].

7 [AA05, Ang05, GPLF05].

91.70/US [Gaz06]. 978-0-321-26877-8 [Har09a]. 978-0-387-29959-4 [Mar09]. 978-0-470-00765-5 [Har09b]. 978-0-521-43592-5 [Sch07]. 978-0-521-87310-9 [Gei09].

A. [Ano07a, Lev06b]. Abelian [Had04]. Abnormality [GVB^{+}09]. Abstention [JLL02]. Abstract [Bar04, BS03, FFPS06, Roc04]. Abstraction [HSGS09, Tof02]. Academic [AR06, Lev06b]. Access [CT07, CLO01a, CLO01b, GB03, PSS09, SG00, Tho05b, WPMK07]. Accessing [ACL03]. Accurate [ELDJ05, VHE08]. Achieving [SSS04]. Acknowledgement [Ano05a, Ano06a]. Across [CSS04]. Action [Col04]. Action-Practical [Col04]. Active [HSM09, LS04, Pan05]. Activity [CQX^{+}09, SJ05]. Acyclic [Luk01]. Ad [AACP09, ABC^{+}07, BDV07, Bet04, Meg09, WTLS02, LCK09]. Adaptable [AK00]. Adaptation [SA09]. Adapting [PPK08, PCW05, Cas06]. Adaptive [GO06, JL07, Wu02]. Adaptivity [Gel06]. Addison [BI06, Har09a]. Adjusting [BO05]. Admission [CV03]. Adopting [Owe03]. aDORe [VBL^{+}05]. Advanced [Dud06, Lan05, Mor06, Pie05, Car06]. Advances [LZ07]. Advice [GO06]. Affinity [TH05]. Affinity-Based [TH05]. Affordance [CT08]. against [HNZI02, Hie03, KK06]. Age [Kov03]. Ageing [SJ02]. Agent [EY04, Kor06a, Ld01, PKSA06, Seh04, DL04]. Agent-based [EY04]. Agents [TKLH04]. Aggregation [CMV08, Her06, STH^{+}09]. agreed [Ano06a]. Agreement [HS07, Tan07, Tse05, YWW09, Tse07]. Aguado [VR04]. Akenine [Rei04]. Akenine-Möller [Rei04]. Akker [Hol05]. AKTiveSA [SRS^{+}07]. Alberto [VR04]. Aldo [Gom06]. Algebra [BBD01]. Algorithm [ADA00, Bal02b, BDK07, CMTK05, Cha05b, CV03, CLH07, CC00b, GP07, GPPR03, GM09, Kor06b, KP02, LCK09, LIPP01, LPS^{+}08, LCT01, Rin03, SNLZ09, SG00, TC04, WLLW06, YWB07, You00, WN95]. Algorithmic [DH08, GHNS04]. Algorithmics [GK08]. Algorithms [AD07, AAL^{+}01, ABPB06, BT06, Bur06a, CW08, CKM03, CG02, CPChX04, CL04, CV05, DFH^{+}05, DÖ05, GK08, GS03, GS08, GNT08, HNW08, LFG^{+}03, LPC09, Mah06, Mar08b, Mor06, PL02, RSW02, San02a, SAS02, SSL08, ST08, Tad08, XU01, Bur05, DLL06, Hro05, Rot06, TKL05, Smi06].

Aliased [RFM09]. Allies [MH06]. Allocation [CCM05, SAA06]. Allowing [JLL02]. Alpha [GR07]. Alternator [LL07]. American [Lom02]. AMULET [Fur02]. Analyses [GPM02]. Analysis [AD07, BFGW02, BF04, BBD01, Bur06a, Car03, Car08, CG02, CP00, FS07b, HMP^{+}02, JWLQ09, Kor06a, LPS^{+}08, Li08, MSK02, MM00, RF05, SF04, SSL08, TFN07, BS03, GSG07, Hro05, McC07, Har09b, Ref09, Roc04]. Analytical [BB05, ZC02]. Analytics [Ref09]. Analyzing [Bar04, De 04a, HD02, Pal06a, GM05, GKKP04]. Andamudi [Jas06]. Andy [Fit04]. Annotated [Cal06]. Anonymous [HSHI06]. Antoniou [Ger05]. Aperiodic [KL02a, Lev05]. Application [BBQV07, CMCS09, SFC09]. Applications [BCKM01, BCGH05, DH08, GCDB01, HOM08, HiOSG08, HH05, KKK^{+}01, KW03, Lam04, LPC09, Man07, Mor06, Pet07, RMJ04, RFM09, STH^{+}09, SFVFM06, Zea00, vRW08, CFB^{+}03, CVS04, TKL05, Gom06]. Applied [MTO^{+}07]. Applying [Alo04, ZTP03]. Appreciation [Lav09, ca09]. Approach [BF04, BHJ06, Cha05b, CL04, CMV08, De 04a, EKWMM02, For08, HC02a, JM08, KK07, Kon01, LLL09, LN07, MV05, VvdAtH07, Vir03, YLK06, GKKP04, Spe04, AR06]. Approximate [CV05]. Approximation [Mar08b]. Approximations [BKTV04, DFGI01]. Arbitrary [CDPV06, DGT04]. Architecture [ACL03, CV04, GBR01, Gut04, HD02, HU08, HH05, LLL07, LLL09, LZ04b, Pan05, Uzu04, ZL03]. Architectures [KLLB04, ZWP03]. Area [Cha05b]. Arguments [Nob00]. Arithmetic [Sta02, ZL03]. Array [DM05, GPM02, KL04, PL02]. Arrays [Tho05b, TFN07]. Art [For04a, Mar05]. Artificial [GS08, Lev07, LZ07]. Ary [AND07, Kor06b, Kor05, XUA00]. ASN.1 [Rin03]. Aspect [CEK03, Col04, PFT05, RB03, SK03]. Aspect-Oriented [Col04, PFT05, CEK03, RB03, SK03]. AspectJ [Col04]. Aspects [Gaz06, LO03, vWJM06]. Aspectual [LO03]. Assembly [Bar04]. Assignment [Cha05b]. Association [CL05]. Assurance [Che01, CC06, DGPT01, GCDB01, Mis04a, TJ01, XK01, Gal04]. Assuring [OC02]. Asynchronous [Bal02a, BFGW02, EB02, GB03]. ATM [Rus06, Zea00]. Atomic [CNV06, MA06]. Attacks [CCD07, HNZI02, KK06, SF04]. August [BCH^{+}05, Vak06]. Australia [ZSP^{+}04]. Authenticated [Chi08, HS07, Tse07]. Authentication [CJ03, CWY05, HSHI06, HL07]. Authority [JLL02]. Automata [Ipa05, LZ04b]. Automated [DHHG06, JWLQ09, MPXT06, SZ09]. Automatic [CQX^{+}09, Spe04, Tad08, You00, AR06]. Automation [ZCHL09]. Availability [FS07a, FS07b, KW08]. Average [Li08]. Average-Case [Li08]. Award [Ano08]. Aware [BLMM06, HPCC^{+}04, SCJ^{+}05, TGSR07, Zea00]. Awareness [CCCL04, SRS^{+}07]. Axiom [ML06]. Axis [For04a].

B [Lev06b, Mis04b, Sch07]. Backtracking [TC03]. Bad [Pip05]. Balancing [LC07, WTSS01, San02a]. Balsa [EB02]. Bandwidth [CSS04, HLL09, MSK02]. Based [AO09, ADA00, BCGH05, CV03, CZCD09, CQX^{+}09, CT07, CW09, CV05, ED09, ED10, Eom08, GHNS04, HGDD05, Her06, HSM09, Jas07, JWLQ09, JCL05, Kor00, LCBK04, LLL09, LZ04b, MH06, Meg09, MC05, MD09, PSS09, TH05, UA02, VBL^{+}05, VvdAtH07, YLW09, YE05, ASDOK01, CMS02, CSS04, CHG07, CC00a, CC00b, EY04, FWLC04, KP04, LH01, LB01, Ngu01, ÖL07, The03, TYH04, WB04, WKwH03, Wu04, YLK06]. Basic [Geo04]. Basis [GOS07]. Battlespace [GHGJ^{+}07]. Bayer [GGA09]. Bayer-sampled [GGA09]. Bayesian [LSY05, NLGP07, PCRZ09]. BCS [CY08]. bed [ABC^{+}07]. Beginnings [Dav01]. Behavioral [ADA00, Mar08a]. behaviors [KS05]. Belief [Li07, LB01]. Bellahséne [Vak06]. Bellare [Cho07]. Benchmark [The03]. Benjamin [Car06]. Beowulf [Hul05]. Bergman [Sta02]. Berners [Jor07]. Berners-Lee [Jor07]. Bertot [Poe06]. Best [Ano08]. Between [Mar08a, Ano06a, HTT06]. Beyond [HiOSG08, Sch03b, See04]. BGM [Alb04]. Bi [YSHP08]. Bi-Directional [YSHP08]. Bibliography [Cal06]. Bidimensionality [DH08]. Bijection [JV07]. Bill [Pol06]. Binary [Bal02a, Luc04, SM02]. Bincode [CC00b]. Bincode-based [CC00b]. Biobjective [DÖ05]. Biopolymer [CHL^{+}08]. Bit [YSHP08]. Bitstreams [KF05]. Blind [MS07, SFC09]. Block [Bal02a]. Blocking [YLK06]. Blocking-based [YLK06]. Blogs [HM06, SI06]. Bloom [LSS02]. BLRL [ELDJ05, VHE08]. Bluetooth [LTL04]. Boerger [Roc04]. Bon [Hol05]. Bongio [Gom06]. Book [Agg02, AR06, Alh03, Alo04, Ama02a, Ama02b, Amb02, Arc02, Aye03, Azi06, BI06, Bas03, Ber02a, Ber02b, BA04, Bob02, Bon02, Bon03, Bot02, Bot03, Bur06a, Bur06b, Car06, Cas06, hC02b, Chu02, Con04, Dam03, De 04a, De 03, Dec02, Dud06, Fit04, For02, FM02b, Gaz06, Gei09, Ger05, Gla03, Gom06, Gro02, Grü07, HM06, Har07, Har09b, Har09a, Hib03, Hir04, Hol05, Hro03, Hun03a, Hun03b, Iwa02, Jas06, Kel04, Ker04, Kir02, Lam04, Lan03, Leo03, Lev06b, Lho04, Lim02, Lio03a, Lio03b, Loc02, Lom03, Lui03, Mac03, Mah06, Mar09, Mar05, Men02, Meo03a, Meo03b, Mic03, Mis03, Mis04b, Mis04a, Mon02a, Mon02b, Mor06, Mot02, Mun02, Nap02, O’K02, Out02, Pal06a, Pal06b, Phi04, Poe06, Pol06, Pom02, Pon02, Por03a, Por03b, Pre02, Ref09, Roc04]. Book [San02b, Sap03, Sch04, Sch06, Sch07, See04, Seh04, Ser03, Smi02, Smi06, Tar02, Tim03, TVT02, Uzu04, Vak06, Val02, VR04, Weg03, Zim03, Zit02]. Boole [Don04, MM07]. Boolean [WMS01]. Bottom [SJ05]. Bound [TC04]. Boundary [YWB07]. Bounded [BK08, SSS04]. Bounds [RSW02]. Brambilla [Gom06]. Branch [HHF05, TC04]. Branch-and-Bound [TC04]. Branch-Coverage [HHF05]. Breaking [Lev05]. Bressan [Vak06]. Brisbane [ZSP^{+}04]. Broadcast [CNV06, LLS02]. Broker [BBQV07, HGDD05]. Brousentsov [Sta02]. Bruce [See04]. Bruijn [KR02]. BSP [SC01]. BSP-style [SC01]. Buffer [GJ01, MPXT06]. Building [Che04, Gel06]. Bullet [O’L05]. Burgin [Smi06]. Burst [PS05]. Bus [DM05, LLL09, PL02]. Business [Kel04, Pal06a, Pol06, GM05, Gou03]. Businesses [HM06, SI06]. Butterworth [Bur06b]. Bytecode [BLMM06, Coo02]. Byzantine [CNV06, KMMS03]. Byzantine-Resistant [CNV06].

C [Car06, Mor06]. C. [Bre08b, Dow08]. Cache [VHE08, HPCC^{+}04, HSM09]. Cache-Aware [HPCC^{+}04]. Caches [VD04]. Caching [DLL06]. Calculation [RGW02]. Calculus [Geo04, Lho04, Li07, Poe06, BC04, SW01]. Cambridge [Gei09, Sch07, Ano07e]. Can [Ver06]. Cancellation [VvdAtH07]. Cancer [GVB^{+}09]. Candidates [Hie03]. Capabilities [CPK08]. Capsule [Ano02a, Ano02b, Ano02c, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano04a, Ano04b, Ano04c, Ano05b, Ano06b, HSM09, Kam04a, Kam04b, Kam05a, Kam05b, Kam05c, Kam05d, Kam06a, Kam06b, Kam06c, Kam06d, Kam06e, Kam07a, Kam07b, Kam07c, Kam07d, Kam07e, Kam07f, Kam08a, Kam08b, Kam08c, Kam09a, Kam09b, Kam09c]. Capsule-Based [HSM09]. Cardinality [Cai08]. Carlo [NLGP07]. Case [AJI07, CQX^{+}09, Eom08, FS07b, JWLQ09, LPS^{+}08, Li08, PGT01, SZ09, ZLB01]. Cases [SZ09]. Castéran [Poe06]. Categorizers [UA02]. CATV [ZC02]. CDMA [CMTK05]. Ceiling [yLKTL00, WKwH03]. Ceiling-based [WKwH03]. Celebration [CK05]. Cellular [LLZ05, LZ04b]. Centered [Owe03]. Centralized [SAA06]. Century [Hir04, Lom02]. Ceri [Gom06, Vak06]. Chains [GJ01]. Challenge [Hir04, Lom02]. Challenges [HM05, MBI^{+}05, MD09]. Changing [HM06, SI06]. channel [WTLS02]. Checking [DFGI01, Mar00, PSR04, SSL07, Xu04]. Checkpointing [BQR01, KHL03, LS06]. Chess [LBI07]. China [LZ07]. Chip [BHJ06, LLL09]. Choice [CN01, Cha03, ML06]. Choices [LBI07]. Chomicki [Lam04]. Chris [Azi06]. Christopher [Grü07]. Circuits [YSHP08]. Clairvoyance [Sub05]. Clairvoyant [Li08]. Clark [Con04]. Class [AO09, You00, CT07]. Classifying [LLH00]. Client [Che05]. Client/Server [Che05]. Closure [GPPR03]. Cluster [Hul05]. Clustered [Tho05b]. Clustering [LCK09, MV05, MSB00, OFW^{+}05, XZJ01, Tam00]. Clusters [FFPS06]. Coarse [GPPR03]. Coarse-Grained [GPPR03]. Coble [HM06]. Cocoon [For04a]. Code [Bar04, DH05, GOS07, MV05, RS03, TGBS05]. Coded [KF05]. Codes [Bal02a, PSR04, PS05, Tad08, TC03, TC04]. Coding [Arn04, Bal02b, Gei09, Had04, Sof05, Teu09, HL08]. Cognitive [WPMK07, vRW08]. Colin [ca09]. Collaboration [Tab06, Har09a]. Collaborations [LO03]. Collaborative [Fit04, Cra03]. Collecting [KK07]. Collection [ZDL^{+}04]. Collective [ZMSH08]. Colleen [Ref09]. Collusion [HNZI02]. Color [SFC09]. Colorings [BKTV04]. Colouring [PT00]. Comai [Gom06]. Combination [Rus06]. Combinatorial [BK08, XU01]. Combined [LC07]. Combining [Abe04, LO03]. Comment [Coh01, Mäk01, TAC01]. Comments [Bre08b, CY08]. Commerce [Sch04, EP04, O’L05]. Commit [HGDD05]. Common [RGW02]. Commons [For04a]. Communicating [Ipa09]. Communication [AK02, CU02, EHKL05]. communications [Sch03a, Phi04]. Community [HGDD05]. Compact [Luk01]. Comparative [CG02]. Comparing [KLLB04, LBI07]. Comparison [CHL^{+}08, LFB00, NPL07, SM02]. Compiler [TGBS05]. Complemental [KW08]. Completeness [CM08]. Complex [SZ09]. Complexity [CHL^{+}08, Cai08, DFL08, FW02, GK08, GKW08, KF05, Mar08b, vRW08]. Component [GO09, Par02, PFT05]. Components [ITV04, PPK08]. Composite [YE05]. Composition [Hil05, PPK08]. Comprehensive [Sub05]. Compression [GPM02, PSR04, PS05, RS03, SS05, Sof05]. Computation [Adi08, Cha06, CM07, Hir04, Mol02, SC01, Weg04, Lom02]. Computational [LPS^{+}08, Pet07, RTV08]. Computations [Had04, San02a]. Computer [AR06, Ano06a, Ano07d, BF00, Bur06a, CY08, Coh01, DH05, DFL08, Fur08, Gel06, Har08, Mäk01, Pal06b, PS03, Poe06, Smi06, Tad08, TAC98, TAC01, Vak06, ZZ04, JCT02, Spe04, Ano06c, GV07a, Hoa07, HKG08]. Computers [ACL03, Ano07e]. Computing [ABDS01, Ano06c, Azi06, CDM04, DÖ05, Don04, HM05, Hul05, KBX^{+}04, KP04, Kov03, LS04, LS06, MBI^{+}05, Mil06, RMJ04, Sch04, SCJ^{+}05, SC08, ZWP03, EP04].

Concept [FM02a]. Concepts [Pan05]. Conceptual [AO00, Ld01, Owe03]. Concern [WS03]. Concerns [RB03]. Concurrency [KHL03, yLKTL00, LH01, LLS02, Puu01]. Concurrent [FS07a, FS07b, GBR01]. Conditional [BO05]. Conduction [LIPP01]. Conference [ZSP^{+}04]. Configuration [MPXT06, SAA06]. Conflicts [Joh00]. Conformance [DHHG06, LBR00]. connected [CL01]. Connectivity [Bet04]. Conquer [AAL^{+}01]. Conscious [Cha05b, Eom08]. Consensus [CNV06, GR07, KMMS03, Ngu01]. Consensus-based [Ngu01]. Consideration [JLH^{+}03]. Considerations [DBS^{+}06]. Consistency [PSR04]. Consistent [BQR01]. Constant [PL02, YWB07]. Constant-Space [YWB07]. Constant-time [PL02]. Constants [CW08]. Constrained [Cai08, CSS04, DBS^{+}06, EY04, PGT01]. Constraint [GS08]. Constraints [BIM^{+}07, Li08, PNM06, Rin03, Sub05]. Constructing [BSC03]. Construction [CU02, SFVFM06, TNM00, TC03, TC04, You00]. Constructions [Poe06, BC04]. Constructive [BF00]. Consumption [CMV08]. Containment [YMB09]. Content [AA05, For08, SFVFM06, SA09]. Content-Intensive [SFVFM06]. Contents [LLH00]. Context [CCCL04, LKCK09, TGSR07, ZZC01]. Context-aware [TGSR07]. Context-sensitive [ZZC01]. Continuous [Car03, Car08]. Contributions [HKG08]. Control [BLMM06, Cha06, CV03, CT07, CLO01a, CLO01b, Hua06, KHL03, yLKTL00, LH01, LLS02, PSS09, Puu01, WTLS02]. Controlling [RVB06, WH04]. Convergence [Sol08]. Conversations [HM06, SI06]. Conversion [PI00]. Converter [CKM03]. Convertible [Chi08].

Coq’Art [Poe06, BC04]. CORBA [CFP^{+}01]. Correction [BEQ00]. Correctness [Ano07b, Ano07c, Hoa07]. Corrigendum [Car08, ED10, Wu04]. Cosmic [Bre08a]. COSMOS [AA05]. COSMOS-7 [AA05]. Cost [CG02, CV03, LLL07, XZJ01]. Cost-Based [CV03]. Costs [LCBK04, Tho05b]. Coteries [KW08]. COTS [ITV04]. Count [Meg09]. Counter [Tam00]. Counter-clustering [Tam00]. Coupling [AO09]. Coupling-Based [AO09]. Course [Lom02]. Coverage [HHF05, YLW09]. Coverage-Based [YLW09]. CPN [Koc06]. Crabtree [Fit04]. Crime [Ref09, McC07]. Critical [BCKM01, TJ01]. Crosscutting [RB03, WS03]. Crossley [Cas06]. Cryptographic [Gut04, Uzu04]. Cryptography [CDD06, YWC08]. Cryptosystems [LZ04b]. CSCW [Pap02]. CSP [Mar00]. Cube [CL01]. Cube-connected [CL01]. Curry [Cas06, PCW05]. Curse [MSB00]. customers [SI06, HM06]. Customizable [PPK08]. Cyberworld [Pip05]. Cycles [CL01]. Cyclic [LSBB04].

D [De 04a, Hol05, Jas06, Lho04, Mis04a, Lev07, WTSS01, Wu02]. D. [Lho04]. Dalkir [Bur06b]. Data [AAL^{+}01, Bal02a, BMNL00, BMN03, Bre08a, Cal07, Cla03, CMV08, De 04a, Gom06, GRS03, Joh00, LPS^{+}08, LV06, McC07, MS07, PSR04, PS05, Ris06, SAA06, Seg07, SZ09, SAS02, SJ02, TZ06b, TKLH04, CFB^{+}03, GKKP04, Mar08a, Ref09, Con04, De 04a]. Data-flow [SAS02]. Data-Intensive [Gom06, CFB^{+}03]. Data-Parallel [AAL^{+}01]. Database [AK00, AJI09, BCH^{+}05, GS08, GPLF05, LCBK04, OR02, Par05, The03, Vak06, ZDL^{+}04, BCH^{+}05]. Databases [Bal05, EKWMM02, yLKTL00, LH01, LFG^{+}03, Man03, Ngu01, Par02, PF01, TS05, CVS04, Lam04]. Datasets [MV05].

David [Bar05, CK05]. Davies [War01]. DC [Lom02]. DCOBE [YE05]. DDoS [WLLW06]. Deadlock [Cha06, CU02]. Deadlock-free [CU02]. December [Ano06a]. Decision [Xu04]. Declustering [BSC03]. Decoding [KW03]. Decomposable [BNM03]. Decomposition [GPPR03]. Deconvolution [SFC09]. Define [TS05]. Defined [LCBK04]. Definition [Ld01]. Degradable [CG02]. Degrees [KL02b]. Delay [Eom08]. Delayed [GHNS04]. Delayed-Multicast [GHNS04]. Delivery [CSS04, TGSR07]. Demand [GHNS04, HLL09, CC00a, WH04]. Demonstrator [SRS^{+}07]. Denial [CCD07, ÖL07, SF04]. DENSITY [GP07]. Dependable [HMP^{+}02]. Dependencies [BLMM06, LSY05]. Deployed [ED09, ED10]. Depth [CDPV06]. Depth-First [CDPV06]. Describing [AJI07]. Description [Gor02]. Design [ACL03, AACP09, BFGW02, Bur06a, Con04, Dud06, GM09, Gut04, Hro05, HTT06, Hua06, JCL05, LCK09, LPS^{+}08, Len06, LS06, Roc04, Uzu04, ZL03, BS03, Cla03, FJ05, Lan05, Sch07]. Designed [Bal02a]. Designer [PS03]. Designing [CFB^{+}03, Cra03, Har08, Pal06a, ST08, GM05, Gom06, Fit04]. Designs [BFS00, CU02]. Desmond [Gei09]. Destination [EY04]. Detecting [DH05, MV05]. Detection [AD07, CCD07, CZCD09, GVB^{+}09, ÖL07]. Detector [ÖL07]. Detectors [KMMS03]. Determining [GB03, ZLB01]. Deterministic [Her06, Hie03]. Developing [HD02]. Development [BFGW02, Eom08, Kor06a, Ld01, NS02, Owe03, Poe06, BC04, For04a, Mar05]. Devices [HC02a, RF05, Tse07]. DFD [LMR04]. Diagnosing [VBvdA01]. Diagnosis [Alb04]. Diagram [CQX^{+}09]. Diagram-Based [CQX^{+}09]. Diagrams [AJI07, You00]. Different [UA02].

Differential [SS05]. Differentiated

[HC02a]. Differentiation [Tad08]. Digest [CZCD09]. Digit [NPL07]. Digital [CV03, Mai07, SAS02, VBL^{+}05]. Digraphs [GY08]. Dilation [Gav00]. Dilemma [Bun06]. Dimensional [LCT01, MTO^{+}07, TZ06b, Li03]. Dimensionality [MSB00, TZ06b]. Dina [Lav09]. d’Inverno [Seh04]. Directed [Luk01]. Directional [YSHP08]. Dirichlet [SFL03]. Discoveries [Meg09]. Discovery [KP04]. Discussant [HKG08]. Discussion [Ano06c, Ano07a, Ano07b, Ano07c]. Disk [HPCC^{+}04, TFN07, TX08, WKwH03]. Disks [TH05, TYH04]. Distance [CPChX04, CV05, PNM06]. Distance-Based [CV05]. Distinguishable [Ris06]. Distributed [AB03, BCS02, CCD07, EKWMM02, HU08, KBX^{+}04, LCK09, LIPP01, LH01, LSS02, MC05, Ngu01, PF01, STH^{+}09, SAA06, TWN08, WLLW06, YE05, Par05]. Distribution [IZ09, LZ04a, Sol03]. Distributions [Sol08]. Diverse [Har08]. Divide [AAL^{+}01]. Divider [NPL07]. Dividing [MK03]. Divisible [Li03]. Division [CW08]. do [GSG07, Har09b]. DOCSIS [ZC02]. Document [GM05, Pal06a, SFVFM06, Pal06a]. Document-Oriented [SFVFM06]. Documents [LLH00, GM05]. Does [EHKL05]. Domains [DT06]. Driven [CCCL04]. DSPN [BF04]. Dynamic [AK00, CC00a, CPK09, LC07, LKCK09, PFT05, SA09, TZ06a, KS05]. Dynamically [GPLF05]. Dynamics [Eom08, STH^{+}09].

Earley [AH02]. Early [Bre08a]. EATCS [Bur06a, Poe06]. Eberbach [CM07]. Economy [Iqb08]. ed [Car06, Hir04]. Edge [RGW02, Yen07]. Edge-Orienting [Yen07]. edited [De 04a, Lam04]. Edition [Cho08, Hul05, Mah06, Rei04]. Editorial [Ang05, Ano07d, Che01, CDM04, Gel06, GTV07, KM09, LZ07, Mol02, Mur02, Mur05, Mur06, Mur08, RB03, Tho05a, ZCHL09]. Editors [DFL08]. EDS [Vak06, Pol06]. Education [BI06, DNM05, MBI^{+}05]. Effect [HU08, RW06]. Effects [Che04]. Efficiency [GS03, OR02]. Efficient [ABDS01, BBvST01, BBQV07, BSC03, CW08, CPChX04, ELDJ05, VHE08, GP07, HM03, Hor01, HL07, HS07, LTL04, LKCK09, RS03, SSNGS00, TGBS05, TYH04, WLLW06, ZMSH08]. Efficiently [XUA00]. EFIS [CHJ^{+}02]. Egon [Roc04]. Election [JLL02]. Electromagnetic [EHKL05]. Electronic [O’L05]. Electronics [GM09]. Elementary [Cha06]. Elements [For08]. Elliott [Sch04]. Elsevier [Dud06, Ref09]. Embedded [ADL^{+}03, KBX^{+}04]. Embeddings [HM03]. Emerging [GPLF05, Lam04, Seg07, CVS04]. emeritus [Ano07e]. Empirical [XCWG06]. Employed [GPM02]. Employing [TYH04]. Encryption [Abe04, Chi08, Cho08, HSHI06, WN95]. Energy [LKCK09]. Energy-Efficient [LKCK09]. Engine [CZCD09]. Engineering [HOM08, LZ07, Pal06a, Pal06b, ZSP^{+}04, GM05, WS03]. engineers [Dan05, Jas06]. Engines [BI06, LBI07, Lev06a]. Enhanced [SRS^{+}07]. Enhancement [CPK09]. Enid [BA04]. Enterprise [Azi06, FM02a, HL03]. Environment [JCT02, YE05]. Environments [AO07, IZ09, LLS02, TGSR07, WKwH03]. Equation [PI00]. Equivalence [Mar08a]. ER-DFD [LMR04]. Eric [Rei04]. Erol [HKG08]. Error [BEQ00, LSBB04]. Error-Correction [BEQ00]. Errors [AD07, PS05]. Escrow [AK02]. Establishment [Hor01]. Estimating [DGPT01]. Estimation [Eom08, HGS^{+}05, KF05]. ethnography [Cra03, Fit04]. Euclidean [CPChX04]. EUR [Bur06b]. Evaluate [PF09]. Evaluation [AACP09, ABC^{+}07, GCDB01, Iqb08, JL07, Kor06a, LCK09, LIPP01, LZ04a, LS06]. Event [CCCL04, Gor02, LZ04a, MC05]. Event-Based [MC05]. Event-Driven [CCCL04]. Evidence [KK07, WB04]. Evolution [CBW08]. Evolutionary [Adi08, Mah06, Mor06, Rot06, TKL05]. Exact [GPM02]. Example [ED09, ED10]. Example-Based [ED09, ED10]. Exception [De 04b, GBR01]. Exchange [CW09, HH05, PVG03]. Exclusion [CDPV05, DGT04, Mai07]. Executables [CV04]. Executed [YMB09]. Execution [Coo02, DÖ05, JWLQ09, Koc06]. Existing [GS03]. Expanded [Cho08]. Experiences [LZ04a]. Experimental [ABC^{+}07]. Experiments [Bre08a]. Expert [GO06]. Explained [Har09a, Tab06]. Explaining [JMT02]. Explicit [Par02]. Exposure [EHKL05]. Extended [Cha03, Pap02, WB04]. Extending [CFP^{+}01]. Extension [GOS07]. Extensions [HOM08]. External [TS05]. Extraction [VR04, NA02].

F [Ger05, Har09b, Mor06]. faced [WMS01]. Facilitation [Har09a, Tab06]. Facing [Ran00]. Factors [SFC09]. Fail [RF05, SSNGS00]. Fail-Secure [RF05]. Fail-stop [SSNGS00]. Fair [PVG03]. Family [CBW08, Hia04]. Fast [Hia04, NPL07]. Fault [ASDOK01, BEQ00, BFS00, GCDB01, Had04, Jas07, KMMS03, MSK02, Wu02, YMB09]. Fault-Containment [YMB09]. Fault-Tolerance [Jas07]. Fault-Tolerant [BEQ00, GCDB01, Had04, MSK02, ASDOK01, Wu02]. Faults [Ran00]. FDDI [ZLB01]. Fear [See04, Sch03b]. Feature [LWY07, NA02, VR04]. Feijs [Sch07]. File [HH05, SS05]. Files [KW03]. Fill [YWB07]. Filter [HLL09]. Filtering [AA05, LSBB04]. Filters [KP04, LSS02]. Finding [LPC09, CC00b]. Finite [Cha05b, DFGI01, Hie03, Ipa05, Ipa09, KK07, KK09]. Finite-state [Hie03]. First [AO00, CDPV06, Hir04, Lav09, Xu04, Ano06c, Lom02]. First-Order [Xu04]. Fit [CCM05]. Fixed [ABC^{+}07, GK08, GS08, GNT08, HNW08]. Fixed-Parameter [GK08, GS08, GNT08, HNW08]. Flip [Cha05b]. Flip-Flop [Cha05b]. Floating [NPL07]. Floating-Point [NPL07]. Flop [Cha05b]. Flow [Bar04, BF00, RF05, SAS02]. Footprint [Che04]. Forbidden [Sap04]. Ford [Lan06, Mar05]. Foreword [Dow08, DFL08]. Form [Hia04]. Formal [Col06, FJ05, Mol02, OC02, Rus06, Sch07]. Formalism [LV06, ZZC01]. Formalizing [PSS09]. Formally [BFS00]. Formulation [AAL^{+}01]. Foundations [vRW08]. Founder [Lav09]. FPGAs [AD07, Jas07]. FPT [LPS^{+}08]. Fragmentation [SG00]. Framework [AO07, ABC^{+}07, CCD07, Coh01, HD02, Ld01, Mäk01, SF04, Sub05, TAC98, TAC01]. Frameworks [PF09]. Franz [Mah06]. Fraternali [Gom06]. Free [CNV06, DFH^{+}05, PT00, YWW09, CU02]. Fringe [MM00]. Front [SA09]. FSMs [DHHG06]. Full [CMS02]. Fully [KKK^{+}01]. Function [LMR04, LCBK04, Ris06]. Functional [ADA00, De 04a, LSY05, GKKP04]. Functions [VD04, WMS01]. Furber [CY08]. Further [ZZ04]. Fusion [LFB00, MTO^{+}07, MS07, TKLH04]. Future [CY08, Eom08, Fur08, MM07, MD09, RVB06]. Fuzzifying [Syr06]. Fuzzy [Kor00].

G [Ger05, Lam04, Pal06b, Sch07, Vak06]. Galin [Mis04a]. Game [Iqb08]. Games [CL04, Lev07]. Gammerman [Ano07a]. Gathering [Ref09, McC07]. GBP [Ref09]. Gelenbe [HKG08]. Gene [CMTK05]. General [CK00]. General-purpose [CK00].

Generalization [GP07]. Generalizations [JV07, KK09]. Generalized [Geo04, SJ05]. Generating [Hie03, Kor05, SZ09, VD04, XUA00]. Generation [AND07, CQX^{+}09, DHHG06, GO09, GHGJ^{+}07, GOS07, KK07, KK09, KL02b, KL03, KL04, TGBS05, XU01, Sof05]. Generators [Bre08b, SM02]. Generic [KK09]. Genetic [AR06, CKM03, Cha05b, Mah06, OYRW05, SG00, Rot06, Spe04, Pol06]. Genome [Kov03]. Geoffrey [Sch04]. Geometric [GKW08]. Geometry [BB05]. Gerchberg [CMCS09]. Gibson [Har09b]. Glimpse [MM07]. Globally [JL07]. Globe [BBvST01]. Glushko [Pal06a]. Gödel [Jor07]. Good [Pip05]. Gouge [Kel04]. GPRS [Wu04, FWLC04]. Grained [GPPR03]. Grammar [ZZC01]. Grand [Hir04, HM05, MBI^{+}05, Lom02]. Graph [CL04, GJ01, RGW02, ZZC01]. Graph-Partition [CL04]. Graphics [BB05]. Graphs [Alb04, AO00, BKTV04, BK08, Luk01, SAS02, Yen07, Xu04]. Gray [De 04a]. Green [De 04b]. Grid [BCGH05, MAD^{+}05, SCJ^{+}05, Tho05a, vM05]. Grids [Che05]. Group [CDPV05, HSHI06, Tan07, Tse07]. Groups [Had04]. Guest [KM09, LZ07, ZCHL09, DFL08]. Guide [Azi06, Dan05, Fit04, Hol05, HL03, Cra03, PQA^{+}03, Jas06]. Guide-IT [Hol05]. Guided [HSGS09]. Gutmann [Uzu04].

H [De 04a, Mor06, Sch07]. H.264 [Ang05, KF05, Sof05]. Hafiz [Cas06]. Haines [Rei04]. Hamming [Bal05]. Handbook [Har07, Sne06]. Handling [GBR01, GRS03, JCL05]. Handoff [LLZ05]. Handshaking [PT00]. Hanoi [Sap04]. Hardbound [Bur06a, Bur06b, Car06, Cas06, Gaz06, Grü07, Jas06, Mah06, Mar09, Mor06, Pal06a, Pol06, Sch06, Smi06]. Hardcover [Gei09, HM06, Har09b, Lev06b, Poe06].

Hardness [CM08]. Hardware [ABPB06, BFGW02, EB02, Gor02]. Harmelen [Ger05]. Hashing [Bal05, CL05]. Heat [LIPP01]. Hedging [Ano07a, GV07a, GV07b]. Heinemann [Bur06b]. Heterogeneous [BCS02, CV03, De 04a, DÖ05, EKWMM02, HC02a, HLL09, GKKP04]. Heuristic [GPM02, HSGS09]. Heuristic-Guided [HSGS09]. Heuristics [BO05]. Hidden [MD09]. Hiding [BG03]. Hierarchical [BSC03, FS07a, HNZI02, Ipa09, JL07, LN07, LLH00, MSK02]. Hierarchies [For04b]. High [Che01, CC06, CPK09, DGPT01, Don04, GCDB01, HNZI02, Koc06, LLL07, Roc04, TZ06b, TJ01, XK01, ZWP03, BS03]. High-Assurance [DGPT01, TJ01, XK01]. High-Dimensional [TZ06b]. High-Level [Roc04, BS03]. High-Performance [ZWP03]. High-Speed [Koc06, LLL07]. Highly [PF01]. Histogram [OR02]. Histogram-like [OR02]. Historical [Dav01]. History [AFB06]. Hoare [Ano07b, Ano07c]. Hoc [ABC^{+}07, BDV07, Bet04, Meg09, WTLS02, AACP09, LCK09]. HOL [NS02]. Home [IZ09, RVB06]. Honeycomb [BB05]. Hop [EHKL05, Meg09, WTLS02]. Hot [HTT06, RSW02]. Hot-plug [HTT06]. Hot-Potato [RSW02]. House [Lav09]. Howard [Cas06, PCW05]. HR [MD09]. Hromkovic [Bur06a]. HSIP [FS07b]. Huffman [KW03]. Human [BA04, PS03, Mum03]. Human-Computer [PS03]. Humans [LBI07]. Hunt [Azi06, Vak06]. Hybrid [Koc06, LZ04b, ZL03]. Hypercube [HM03]. Hypercube-like [HM03]. Hypercubes [ASDOK01, SAOKM00]. Hyperion [GHGJ^{+}07]. Hypermedia [MA06].

I/O [HC02a, XZJ01]. Ideal [Ano07b, Ano07c, Hoa07]. Identify [WLLW06]. Identity [Her06].

Identity-Based [Her06]. II [FS07b, Lan06, OYRW05, Pol06]. Im [Pal06a]. Image [BG03, CMCS09, CPK09, ED09, ED10, GGA09, GPLF05, MS07, MD09, PCRZ09, VR04, NA02]. Images [CC00b, OFW^{+}05, RFM09, SNLZ09, SFC09, VMMK09]. Imaging [Gre09]. Iman [Cas06]. Impact [Kel04, Gou03]. Impacts [KHL03]. Implementation [AD07, ACL03, ABC^{+}07, Hie03, HTT06, LPS^{+}08, LZ04a, Vir03, Zea00, CK00, Mis04a, Pan05]. Implementations [CBW08]. Implications [CY08, Fur08, GM09, STH^{+}09]. Improved [AFB06, BT06, TNM00]. Improves [Eom08]. Inclusion [Mai07]. Incomputable [Ver06]. Increasing [GS03, TZ06b]. incremental [KS05]. Independent [ABPB06]. Index [BNM03, TZ06a, TZ06b, VD04, ZMSH08]. Indexing [BEQ00, Cha05a]. Indices [KR02]. Induction [Sol08]. Inductive [Grü07, Poe06, BC04]. Indulgent [GR07]. Industry [CY08, Fur08]. Infection [Coh01, TAC98, TAC01, Mäk01]. Inference [FW02, Grü07, Li07]. Influencing [WPMK07]. Infocommunications [Kuz05]. Informatics [Pal06a, GM05]. Information [Ano07e, Bar04, BG03, BF00, Col06, EKWMM02, Eom08, For08, GHGJ^{+}07, HC02a, IVS06, Iqb08, KKK^{+}01, Lev06b, LB01, MM07, Mor06, Pal06b, RF05, TGSR07, WPMK07, ZSP^{+}04]. Information-Dynamics-Conscious [Eom08]. InfoSCI [BA04]. Initiated [San02a]. Initiation [NLGP07]. Inline [CCD07]. Innovative [LPS^{+}08]. Inpainting [CMCS09, FSM09]. Input [AFB06, DHHG06]. Instant [MK03]. Instruction [AFB06, BHJ06]. Instruction-Memory [AFB06]. Integer [CW08, HH06, Mar09, Teu09, PW06]. Integrate [EKWMM02]. Integrated [CHG07, KK07, KK09, KL02a]. Integrating [CK00, De 04a, LKCK09, GKKP04].

Integration [AO09, GO09]. Intelligence [GS08, Lev07, LZ07, Pet07, Ref09, McC07]. Intelligent [EY04, Gaz06, vWJM06]. Intensive [Gom06, SFVFM06, CFB^{+}03]. Interacting [FS07a]. Interaction [AJI09, PS03]. Interactions [GT08, WH04]. Interactive [BC04, Lev07, PS03, WH04, HNZI02, Poe06]. Interconnected [BQ03, KKK^{+}01]. Interconnecting [ABC^{+}07]. Interface [CK00, FFPS06, PI00]. Interface/implementation [CK00]. Interfaces [CT08, CFP^{+}01, Har08]. Interleaved [Joh00]. Intermediate [RS03]. International [BCH^{+}05, Vak06, ZSP^{+}04]. Internet [Con04, ACL03, ABC^{+}07, Cla03, JLHD01, JLH^{+}03]. InternetBeans [For04a]. Interoperation [Col06]. Interpolated [GRS03]. Interpretation [Bar04]. Interval [Gav00, LMOP00]. Intervals [MH06]. Intractability [CM08]. Intriguing [MM07]. Introduction [BI06, Bur06a, GTV07, Man03, Ric08, War01, HL08, Hro05, Lev06a, Gei09]. Introductory [Rüd07]. Intrusion [CZCD09, CDW09]. Invariants [VvdAtH07]. Inverse [LIPP01]. Inversion [Arn04, Car03, Car08]. Inverting [SFL03]. Investigation [SC01]. Invitation [GK08]. IP [Con04, Set05, CHG07, Cla03, HL07, Koc06]. ISBN [AR06, Azi06, BI06, Bur06a, Bur06b, Car06, Cas06, Dud06, Gaz06, Gei09, Gom06, Grü07, HM06, Har07, Har09b, Har09a, Jas06, Lev06b, Mah06, Mar09, Mor06, Pal06a, Pal06b, Poe06, Pol06, Ref09, Sch06, Sch07, Smi06, Vak06]. ISBN-10 [Sch07]. ISBN-13 [Gei09, Har09b, Har09a, Mar09, Sch07]. Israel [HM06]. Issue [Che01, DFL08, GTV07, LZ07, ZCHL09]. Ivarama [Jas06]. Ives [Vak06].

J [De 04a, Gaz06, Hir04, Hol05, Lev06b, Pal06a]. J. [Lam04]. J2EE [Azi06, HL03]. Jansen [Lev06b]. January [Ano06a, Lom02]. Java [Azi06, Dud06, Mar05, BCS02, CQX^{+}09, CCM05, Coo02, For04a, HL03, Lan05, LBR00, Vir03, Zea00]. Jean [Har09a]. Jeffery [Pal06b]. Jochen [Phi04]. John [Cas06, Gaz06, HM06, Azi06, Har09b]. Johnston [Lav09]. Join [PNM06, XZJ01]. joins [VvdAtH07]. Joint [MS07]. Jones [Ano07e]. Jonkers [Sch07]. Jorna [Gaz06]. Journal [Ano06c, GV07a, Hoa07, HKG08, Ano06a, Ano07d, DFL08]. JPEG [KW03]. Jr [Hir04]. Jumping [CMTK05]. JUnit [For04a]. Juraj [Bur06a]. Just [HSM09]. Just-In-Time [HSM09].

K. [Mor06]. Kalman [HLL09]. Kang [Ker04]. Karen [Ano07e]. Kaufmann [Gom06]. Keith [Pal06b]. Kernel [CV04, CPK08]. Kernel-Level [CV04]. Kerschberg [De 04a]. Kevin [Dud06]. Key [AK02, CW09, DBS^{+}06, HNZI02, Hor01, HS07, Tan07, Tse05, Tse07]. Key-Sharing [HNZI02]. Khor [Mor06]. Kimiz [Bur06b]. Kinds [Sol08]. King [De 04a]. Kluwer [AR06, Lev06b]. Knowledge [Abe04, Bur06b, Dal05, Kor00, Lev06b, LZ07, Mor06, O’L05, SSL07]. Knowledge-Based [Kor00]. Kolmogorov [Sol03]. Kratochvíl [Mis04b]. Kurki [Sch06]. Kurki-Suonio [Sch06].

L [De 04a]. L. [Sch07]. Labelling [Cal06]. Laddad [Col04]. Language [De 04b, EB02, HC02a, Owe03]. Languages [CK00, Gor02, Ipa05, Rüd07, ZZC01, Pie05, Car06]. Lano [Dud06]. Laplace [Car03, Car08]. Large [BQ03, Cha05a, FFPS06, OFW^{+}05]. Large-Scale [BQ03, FFPS06, OFW^{+}05]. LARPBS [DM05]. Latches [Ric08]. Latency [GJ01, LLL07, LS06, SAOKM00]. Laurence [Mar09]. Layer [ABC^{+}07, Koc06, LLL07]. Layer-2 [ABC^{+}07]. Layout [CL01]. Layouts [TYH04]. Lazy [CCM05, MM00]. leaders [Tab06, Har09a]. Learning [Ano07a, GV07a, GV07b, GA05, JM08, LWY07, Sol03]. Lecture [Mai07, Pal06b, Vak06, Ano06c, CY08, Don04, Don01, GV07a, HKG08, Hoa07, MM07, Pip05, Ran00, Sol03]. Lee [Mor06, AR06, Jor07]. Legality [For04b]. Length [FW02, PSR04, PS05, TC03, TC04, Grü07]. Leo [Lan06]. Level [BHJ06, CV04, GBR01, Roc04, BS03]. Levene [BI06]. Liberal [DFH^{+}05]. Library [CV03]. Light [Hua06, Mis04b, KM02]. Like [WH04, HM03, OR02]. Likelihood [ÖL07]. Limited [AK02, Tse07]. Line [Cho08, BCS02]. Linear [DM05, HH06, SSL08, TVM00, YWB07]. Linear-Time [YWB07]. Link [Eom08, LLL07, LTL04]. Link-Delay-Information [Eom08]. Link-State [Eom08]. Linked [Kor05]. Linux [Hul05]. Lists [AO07]. Lists-on-Lists [AO07]. Liveness [CN01]. Load [LC07, San02a, WTSS01]. Load-balancing [San02a]. Loads [Li03]. Local [BDV07, Kon01]. Locality [AO07]. Locating [BIM^{+}07, KKK^{+}01]. Location [CG02, LLZ05, Meg09, PF01, The03, TKLH04]. Location-based [The03]. Loftus [Azi06]. Logarithmic [BNM03]. Logging [SSS04]. Logic [AO00, BF00, CMS02, CLO01a, CLO01b, LB01, Pap02, Xu04]. Logical [LB01]. Logics [CVS04, SSL07, Lam04]. Lomonaco [Hir04]. Long [CL01]. Lookup [LSS02]. looping [DLL06]. Loopless [KL02b, KL03, KL04]. LOTOS [CMS02]. Low [HNZI02, LLL07, LS06]. Low-Latency [LS06]. Lower [RSW02]. LR [BNM03]. LR-tree [BNM03]. Luck [Seh04]. Lun [Gei09].

M [Alo04, De 04a, Gaz06, Sch07, Vak06]. M. [Con04, Mis04b]. MAC [WTLS02]. Machine [Ano07a, Cha05b, GV07a, GV07b, Hie03, LWY07, Sol03]. Machines [Ipa09, KK07, KK09, PI00, Roc04, BS03, Ipa03, Ipa05]. Made [LBI07]. Maintaining [ZDL^{+}04]. Maintenance [BF04]. Mäkinen [Coh01, TAC01]. Making [MAD^{+}05]. Malicious [Tse05]. Management [Bur06b, CMTK05, CG02, CHG07, CC00a, CK00, De 04a, FM02a, Har07, HLL09, KP02, LLZ05, LCBK04, Len06, LV06, O’L05, Seg07, Sne06, SCJ^{+}05, ZDL^{+}04, vM05, Dal05, GKKP04, PQA^{+}03, Lev06b, Hol05]. Management-Modeling [De 04a]. Managing [GJ01, SJ02]. MANET [Len06]. Manipulating [WPMK07]. Marco [Gom06]. Maria [Pal06b]. Maristella [Gom06]. Mark [HC02a, Seh04, BI06, Smi06, VR04]. Mark-up [HC02a]. Markov [Car03, Car08, MD09]. Martin [Cas06]. Masked [LMOP00]. Mastermind [CL04, CLH07]. Matching [CZCD09, LTL04]. Matching-Based [CZCD09]. Matera [Gom06]. Material [Col06]. Materialization [JL07]. Mathematical [Bun06, Hir04, Lom02]. Matrix [SAOKM00]. Matrix-Transpose [SAOKM00]. MAX [GP07]. MAX-DENSITY [GP07]. Maximal [BIM^{+}07]. Maximum [BT06, ÖL07, RGW02]. May [Pol06]. Mccue [Ref09]. McGibbon [Mis04b]. McGrath [Pal06a]. MDA [Dud06, Lan05]. MDGs [Xu04]. Mean [Ver06]. Means [MSB00]. Measurements [ACL03]. Measuring [LMR04]. Mechanism [CK00, FWLC04, HP00, SJ02, Wu04]. Media [Pol06]. Median [LCT01]. Mediator [JL07]. Medical [Gre09, RMJ04]. Medium [CT07, GB03, RW06]. Medium-Term [RW06]. Memoir [Bar05]. Memorial [Don01, Ran00]. Memoriam [Mur03]. Memories [Joh00]. Memory [AFB06, CCM05, DNM05, HNZI02, LLL07, MSK02]. Merging [EM00]. Mesh [WTSS01]. Meshes [CC00b, LCT01, Li03, Wu02]. Message [CW09, FW02, Grü07, SAOKM00, ZLB01]. Messages [KBX^{+}04]. Meta [GBR01]. Meta-Level [GBR01]. Metamodel [GOS07]. Metamodelling [CEK03]. Method [BNM03, Cha05a, CMCS09, GHG00, JL07, KK09, LC07, MD09, Roc04, BS03]. Methodology [ADL^{+}03, CDW09, Owe03, TS05]. Methods [Gaz06, LPS^{+}08, Mol02, NLGP07, OC02, PCRZ09, TNM00, Vir03, vWJM06]. Metric [Bal05, RVB06]. Metrics [PF09]. Meyden [Lam04]. Meystel [Gaz06]. Michael [Seh04]. Microarray [OFW^{+}05]. Micropipeline [BFGW02]. Microprocessors [Fur02]. Microsoft [Cal07]. Microthreading [BHJ06]. Middlesex [ca09]. Middleware [CCCL04]. Mike [Pal06b]. Millennium [Hir04, Lom02]. Milner [Ano06c]. Minimal [DFH^{+}05, Wu02]. Minimality [Ipa03, Ipa05]. Minimization [WMS01]. Minimize [Meg09]. Minimum [FW02, Grü07, PL02]. Mining [CL05, Ref09, McC07]. Mirror [Sta02]. Mirror-symmetrical [Sta02]. Mirrored [TH05, TX08]. Missing [MS07]. Mixed [yLKTL00, Mar09, PW06]. Mobile [ABDS01, ACL03, AB03, BBvST01, CCCL04, CDM04, CHG07, EP04, Kor06a, LCK09, Lho04, LS06, Meg09, PF01, Sch03a, TGSR07, TKLH04, WTLS02, SW01, Tse07, Sch04, Phi04]. Mobility [CHG07, IZ09, KP02]. MOD [WH04]. Modal [CMS02]. Model [Abe04, Alb04, Chi08, Cho07, DFGI01, IVS06, LB01, SSL07, Xu04, ZC02, Gou03, Lan06, Kel04]. Modeling [Adi08, AJI09, AFB06, BF04, De 04a, Kor00, LCBK04, MD09, RS03, RTV08, TJ01, vRW08, GKKP04, KS05]. Modelling [AA05, Coh01, Mäk01, MAD^{+}05, SZ09, TAC98, TAC01, Tof02]. Models [Car03, Car08, CM07, HMP^{+}02, Kor06a, Ris06, Weg04]. Modern [BCKM01]. Modes [JWLQ09]. MODIS [SNLZ09]. Modular [CP00, VBL^{+}05]. Modules [LO03]. Moduli [Hia04]. MOF [GOS07, PQA^{+}03, Hol05]. Möller [Rei04]. Monitoring [AK02, BCS02, LZ04a, RMJ04]. Monographs [Smi06]. Monte [NLGP07]. Morgan [Gom06]. Morphic [Luk01]. Mosaics [BG03]. Motivations [Gaz06, vWJM06]. Move [LBI07]. Moves [Sap04]. Moving [LFG^{+}03]. MPEG [Par05, Sof05, AA05, Ang05, GPLF05, Par05]. MPEG-21 [Ang05, Par05]. MPEG-4 [Sof05]. MPEG-7 [Par05, AA05, Ang05, GPLF05]. MPLS [Set05]. Multi [BDK07, CT07, EHKL05, Hor01, HS07, JCT02, JLL02, Kor06a, MTO^{+}07, PKSA06, TYH04, Tse05, WTSS01, WTLS02]. Multi-Agent [Kor06a, PKSA06]. Multi-Authority [JLL02]. Multi-channel [WTLS02]. Multi-class [CT07]. Multi-computer [JCT02]. Multi-Dimensional [MTO^{+}07]. Multi-Hop [EHKL05, WTLS02]. Multi-Node [WTSS01]. Multi-Party [Tse05, Hor01]. Multi-Servers [HS07]. Multi-Wave [BDK07]. Multi-zone [TYH04]. Multicast [ABDS01, AB03, GHNS04, WTSS01]. Multidimensional [Cha05a]. Multimedia [Ang05, BBD01, CV03, CSS04, CC00a, HOM08, HL07, Par05, RVB06, WH04, Sof05]. Multimedia-on-Demand [WH04]. Multiobjective [CMTK05, Mor06, TKL05]. Multiple [BIM^{+}07, EY04, LSBB04, MPXT06, MS07, WKwH03]. Multiple-disk [WKwH03]. Multiples [NPL07]. Multiplication [CW08]. Multiplier [CBW08, Hia04]. Multiprocessor [BQ03, GCDB01]. Multiprocessors [BHJ06, MSK02]. Multirepeats [BIM^{+}07]. Multiset [KL04]. Multispectral [VMMK09]. Multitarget [NLGP07]. Multithreaded [TGBS05, URŠ02]. Multiway [Xu04]. Mumford [BA04]. Mutation [SZ09]. Mutual [CDPV05, DGT04].

N [Mar05]. Naked [SI06, HM06]. Name [LSS02]. Navigation [BI06, Lev06a]. Nearest [WB04]. Needham [Mur03]. Negotiation [DBS^{+}06]. Neighbor [CC00b]. Neighbor-finding [CC00b]. Neighbours [WB04]. Net [JCL05, Kor00]. Net-Based [JCL05]. Nets [CN01, Cha03, Cha06, CP00]. Network [ACL03, BQ03, HL08, Koc06, LC07, LPC09, ÖL07, RVB06, TKLH04, XK01, ZC02, Gei09, Set05]. Networks [AACP09, ABC^{+}07, BDV07, Bet04, BF00, CKM03, CG02, CSS04, CT07, Con04, CDPV06, CMV08, DBS^{+}06, Gel06, GT08, HTT06, HSM09, HLL09, KBX^{+}04, KKK^{+}01, Kor06a, KP02, LCK09, LLZ05, LS04, LSY05, LKCK09, MSK02, Meg09, STH^{+}09, SC08, TWN08, WTLS02, YWW09, Zea00, ZLB01, Cla03]. Neural [ÖL07, SC08]. Neuronal [GT08]. Newsome [Cas06]. Next [GHGJ^{+}07, Sof05]. Next-Generation [GHGJ^{+}07, Sof05]. Nigel [Sch04]. Nixon [VR04]. Node [FWLC04, WTSS01, Wu04]. Nodes [LSY05]. Non [DFGI01, HNZI02, Hie03, Li08, XZJ01]. Non-Clairvoyant [Li08]. Non-deterministic [Hie03]. Non-finite-state [DFGI01]. Non-interactive [HNZI02]. Non-uniform-sized [XZJ01]. Normalizing [PSS09]. Norway [BCH^{+}05, Vak06]. Note [CWY05]. Notes [Pal06b, Vak06]. Notifications [LZ04a]. Novel [AO07, CMV08]. November [ZSP^{+}04]. Novices [GHG00]. NSL [VHE08]. Number [Bre08b, Meg09, Sta02]. Numbers [GHG00].

NWS [BCGH05].

O [HC02a, XZJ01]. O. [Alo04]. OBDDs [SSL07]. Object [AK00, Ala00, ADL^{+}03, CSS04, CK00, JCT02, LZ04a, LH01, LCBK04, PF09, VBL^{+}05, YE05]. Object-Based [YE05, CSS04, LH01]. Object-Oriented [Ala00, AK00, ADL^{+}03, CK00, JCT02]. Object-Relational [LCBK04]. Objective [DT06]. Objects [BBvST01, CV03, CCM05, LFG^{+}03, PF01, XZJ01]. Observations [CU02]. Occasion [Ano07d]. ODMG [TS05]. ODP [RTV08]. off [DÖ05]. Omega [BQ03]. On-Chip [LLL09]. On-demand [CC00a]. On-line [BCS02]. One [AK02, RSW02]. One-to-one [RSW02]. One-way [AK02]. Online [Li08]. Ontological [GOS07]. Ontologies [Col06, O’L05]. Ontology [FM02a]. Opening [LBI07]. OpenMP [TGBS05]. Operation [Par02, Cla03, Con04]. Operations [Hol05, JLH^{+}03, PQA^{+}03]. Operators [BMNL00]. Opportunities [vRW08]. Optical [Set05]. Optimal [BDK07, CL04, CW09, JLHD01, KLLB04, LPC09, LL07, RFM09, SAS02]. Optimization [Adi08, BK08, Cai08, CLH07, LLL09, OR02, ZDL^{+}04]. Optimized [JL07]. Optimizer [LCBK04]. Optimizing [CSS04]. OR-joins [VvdAtH07]. Oracle [Abe04, Chi08]. Order [AO09, AO00, ABPB06, Xu04, AND07]. Order-Independent [ABPB06]. Ordered [TZ06a]. Ordering [LLS02, Tad08]. O’Reilly [Pol06]. Organization [Kel04, Gou03]. Organizations [TX08]. Organized [STH^{+}09]. Organizing [AO07, BBQV07]. Oriented [AA05, Ala00, ABPB06, Col04, LWY07, PF09, PFT05, SFVFM06, AK00, ADL^{+}03, CK00, CEK03, JCT02, RB03, SK03]. Orienting [Yen07]. Orlowska [Pal06b]. OSGi [IZ09]. Output [Cha05b, DHHG06].

Outsourcing [Kel04, Gou03]. Ovarian [GVB^{+}09]. Overcoming [MSB00]. Overlapping [TVM00]. Overlay [BBQV07, STH^{+}09]. Overview [Rüd07, ST08].

P [Con04, De 04a, Hol05, Jas06, Pal06b, DM05, PGT01]. P. [De 04a]. P2P [Che05, Man07]. Packet [Dav01, RSW02]. Page [Cha05a]. Papazoglou [Pal06b]. Paper [Ano08, War01]. Paperback [Ref09]. Papoulis [CMCS09]. Paradigm [SFVFM06]. Paradigms [Bur06a, Hro05]. Parallel [AND07, AAL^{+}01, CPChX04, GPPR03, Had04, JCT02, KW03, KLLB04, Li03, Li08]. Parallelism [BHJ06]. Parameter [GK08, GS08, GNT08, HNW08]. Parameterized [CHL^{+}08, Cai08, CM08, DM05, DFL08, GK08, GKW08, GY08, Mar08b, ST08, vRW08]. Parameters [HiOSG08]. Parametric [Vir03]. Parent [MM00]. Pareto [SA09]. Parsers [SJ05]. Parsing [AH02]. Part [FS07a, FS07b]. Partial [Her06]. Participants [Tse05]. Partition [CL04, TZ06a]. Partitioned [KL02a]. Party [CW09, Tse05, Hor01]. Passing [Vir03]. Password [CPK08, CW09]. Password-Based [CW09]. Password-Capabilities [CPK08]. Path [GPPR03, Meg09, SG00, SAS02]. Pattern [CZCD09, LTL04, Tam00]. patterns [DLL06]. PDetect [MV05]. Pearson [BI06]. Peer [HH05, TWN08]. Peer-to-Peer [HH05, TWN08]. Peers [Che05]. Pen [Man07]. PEPA [BCGH05]. Perfect [CL05, EM00, Iqb08]. Performability [Tho05a]. Performance [AACP09, AFB06, Don04, HGDD05, Hil05, KLLB04, Kor06a, LIPP01, Li08, MAD^{+}05, SCJ^{+}05, TX08, UA02, XCWG06, Zea00, ZC02, ZWP03]. Performance-Aware [SCJ^{+}05]. Periodic [KL02a]. Periodicity [Sub05]. Permutations [KL04]. Persistent [SJ02, TZ06a]. Personal [Bar05]. Perspectives [ZTP03, Alo04]. Pervasive [CDM04, KP04, LS04]. Peter [Uzu04]. Petri [CN01, Cha06, CP00, JCL05, Kor00]. Peuser [Alo04]. Phase [CLH07, HGDD05, LLL09]. Phase-Based [LLL09]. Phillips [Sch04]. Phishing [KK06]. Phylogenetics [GNT08]. Pi [Lho04, SW01]. Pi-Calculus [Lho04]. Piconets [LTL04]. Pierce [Car06]. Piero [Gom06]. Pierre [Poe06]. Piggyback [ZDL^{+}04]. Pipelined [DM05]. Pipelines [ADA00]. Place [SS05]. Placement [CKM03, JLHD01, JLH^{+}03, Joh00]. Plagiarism [DH05, MV05]. Planar [Yen07]. Planning [Gaz06, Mar09, vWJM06, PW06]. Platform [IZ09, KBX^{+}04, PFT05]. plug [HTT06]. Pochet [Mar09]. Pocket [Hol05, PQA^{+}03]. POEMS [JCT02]. Poernomo [Cas06]. Point [MH06, NPL07]. Point-Based [MH06]. Points [LMR04]. Polarity [Cha05b]. Policy [AK00, LTL04]. Politics [Cho08]. Polling [LTL04]. Polymorphic [JMT02]. Pools [MPXT06]. Potato [RSW02]. Poulovassilis [De 04a]. Power [ACL03, CMV08, DBS^{+}06, HU08, WTLS02]. Power-Saving [ACL03]. pp [BI06, Bur06a, Bur06b, Car06, Cas06, Dud06, Gaz06, Gei09, Gom06, Grü07, HM06, Har07, Har09b, Har09a, Jas06, Lev06b, Mah06, Mar09, Mor06, Pal06a, Pal06b, Pol06, Ref09, Sch06, Smi06, Vak06, AR06, Azi06]. Practical [AH02, Col04, Fit04, HNW08, LPS^{+}08, Sch06, TGBS05, Cra03, KS05]. Practice [Bur06b, GK08, Pol06, ca09, Dal05, OYRW05, Zha03, Ker04]. Practices [Seg07]. Pragmatic [CDW09]. Precedence [Li08, PGT01]. Predating [GVB^{+}09]. Predictable [MAD^{+}05, SSS04]. Prediction [Meg09, XK01]. Prediction-Based [Meg09]. Predictions [Ano07a, Che04, GV07a, GV07b]. Predictive [HLL09, Ref09, McC07].

Prefetch [AK00]. prefetching [DLL06]. Prefix [Had04]. Presence [SAOKM00]. Presentations [HC02a]. Preserving [JCL05, YMB09]. Press [Car06, Gei09, Pal06a, Sch07]. Prevention [CCD07, CDW09]. primer [AV04, Ger05]. Primitive [MH06]. Principle [Sta02]. Principles [Pet07, UA02]. Privacy [Cho08]. Probabilistic [CDD06, HGS^{+}05, IVS06, Sol08, Tof02]. Probability [ASDOK01]. Probability-based [ASDOK01]. Problem [AO09, BT06, Cal06, CKM03, JCL05, LIPP01, LCT01, YWW09, YWB07, ML06]. Problems [Cai08, EY04, GKW08, GS08, GY08, HH06, PL02]. Procedures [KHL03]. Proceedings [Pal06b, Vak06, ZSP^{+}04, BCH^{+}05]. Process [BBD01, Tof02]. Processes [DFGI01, FS07a, VBvdA01, SW01, Lho04]. Processing [Bre08a, GJ01, LZ04a, Li03, Mor06, PNM06, SAS02, TFN07, TVM00, XZJ01, ZMSH08, NA02, VR04]. Processor [ELDJ05, VHE08, Koc06, PL02]. Processors [BCKM01, Jas06, URŠ02, Dan05]. Production [PW06, Mar09]. Professional [Azi06, Man07]. Professor [Ano07e, CY08, ca09]. Program [Ano07b, Ano07c, DFH^{+}05, Hoa07, JWLQ09, JM08, Poe06, BC04]. Programmer [Man07]. Programmers [Jas06, Dan05]. Programming [AR06, Ala00, Car06, CK00, Col04, GHG00, GA05, HH06, MC05, Pol06, RB03, Rüd07, SK03, OYRW05, Pie05, PW06, Spe04, Mar09]. Programs [AFB06, Cas06, CQX^{+}09, HHF05, TGBS05, DLL06, PCW05]. Progressive [ZMSH08]. Project [Har07, Har09a, Sne06, Tab06]. Projection [RFM09]. projects [ZTP03, Alo04]. Proof [Abe04, Cho07]. Proofs [Cas06, PCW05]. Proofs-as-Programs [Cas06, PCW05]. Propagation [Che04]. Properties [Bal05, Bar04, TJ01]. Property [JCL05, Mar00]. Property-Preserving [JCL05]. Propositional [LB01]. Prospects [GK08]. Protecting [KK06]. Protection [Set05]. Proteomic [GVB^{+}09]. Protocol [ABDS01, AACP09, CT07, CJ03, CWY05, Cho07, GB03, HGDD05, Hor01, HS07, yLKTL00, Meg09, PI00, Rus06, Tan07, Tse05, WTLS02, WKwH03, YLK06, PCW05, Tse07, Cas06]. Protocols [AB03, CFP^{+}01, CU02, Con04, CNV06, Koc06, LH01, OC02, YMB09, Cla03]. Provably [CW09, HL07, HS07]. Proving [Poe06, BC04]. Proxies [JLHD01, JLH^{+}03]. Proxy [CHG07]. Proxy-based [CHG07]. PSIM [XCWG06]. PSQA [RVB06]. Public [Lev06b, SJ04]. Publish [BBQV07]. Publish/Subscribe [BBQV07]. Publishers [AR06, Gom06, Lev06b]. Pultorak [Hol05]. Pure [KR02]. purpose [CK00]. Push [FWLC04, Wu04]. Pyramidic [OFW^{+}05].

QoS [RVB06, Zea00]. QoS-aware [Zea00]. Quadtrees [TVM00]. Quagliarello [Hol05]. Qualitative [LN07]. Quality [Mis04a, XK01, Gal04]. Quantitative [HMP^{+}02]. Quantum [AR06, Lom02, MM07, Rüd07, Spe04, Hir04]. Queries [BSC03, Cha05a, CV05, PNM06, The03, ZMSH08]. Query [JL07, LCBK04, OR02, Owe03, TVM00, ZDL^{+}04]. Quotient [NPL07, PI00].

R

[Gaz06, Hol05, Lam04, Vak06, CV05, Li07]. R-Calculus [Li07]. R-trees [CV05]. R. [Sch06]. R.C.P.C.P. [WKwH03]. Radio [Che04]. Radix [NPL07]. Radix-4 [NPL07]. RAID [Tho05b]. RAID5 [TFN07]. Rainbow [PT00, BFGW02]. Ramnivas [Col04]. Random [Abe04, Bre08b, Chi08, ÖL07, SM02]. Randomization [Car03, Car08]. Randomized [BO05, Bur06a, GS03, San02a, VD04, Hro05]. Range [BSC03, CPK09]. RASCAL [RGW02]. Raster [BB05]. Rating [TWN08]. Rational [SFC09]. Ray [Bre08a]. Re [WS03, Dow08]. Re-engineering [WS03]. Reactive [Meg09, Sch06, KS05]. Read [JLH^{+}03]. Real [Alo04, ADL^{+}03, HPCC^{+}04, KL02a, Kon01, KHL03, yLKTL00, SSS04, Sub05, ZTP03, Rei04]. Real-Time [HPCC^{+}04, KHL03, SSS04, Sub05, ADL^{+}03, KL02a, Kon01, yLKTL00, Rei04]. Real-World [Alo04, ZTP03]. Reasoning [LN07, PS03]. Rebuild [TFN07]. Receiver [San02a]. Reconfigurable [CC00b, DM05, LZ04b, PL02, ZL03]. Reconstruction [CPK09, ED09, ED10, SNLZ09]. Recovery [HP00, LH01, Set05, SSS04]. recursive [Bur05, Smi06]. Redesigning [Mum03, BA04]. Reduce [EHKL05, Joh00, XZJ01]. Reduced [yLKTL00, SJ05]. Reducing [CMV08]. Reduction [JCL05, LLL07]. Referees [Ano05a, Ano06a]. Reference [AO07, DLL06]. Refinement [HSGS09]. Regenerative [Car03, Car08]. Region [TYH04]. Region-based [TYH04]. Regional [CHG07]. Regions [VvdAtH07]. Registration [CHG07, RFM09]. Regression [LCBK04]. Regression-Based [LCBK04]. Regular [XUA00]. Regularization [ED09, ED10]. Rejoinder [GV07b]. Rejuvenation [DGPT01]. Related [PL02]. Relating [Gor02]. Relational [LCBK04, LV06]. Relationships [AJI07]. Relaxed [VvdAtH07]. Reliability [DÖ05, SAA06, TX08]. Reliable [AB03, BQ03]. Remote [SNLZ09]. Rendering [ABPB06, Rei04]. Replicated [JLHD01]. Reply [CM07, Coh01, TAC01]. Report [CHJ^{+}02]. Repository [VBL^{+}05]. Representation [BMN03, Kor05, LN07].

Representations

[FSM09, Mah06, PS03, Rot06].

Representing [MK03]. Requirements [GJ01, KM02, Mis04b]. Research [HM05]. Reservations [MAD^{+}05]. Residue [Hia04]. Resistance [HNZI02]. Resistant [CNV06, Tse05]. Resolution [CMCS09, CPK09, ED09, ED10, Gre09, MD09, PCRZ09, RFM09, SNLZ09, SFC09, VMMK09]. Resource [CMTK05, CC00a, DT06, HGDD05, HH05, Tse07]. resource-limited [Tse07]. Response [ZLB01]. Responses [Ano07c]. Restoration [GGA09, MS07, Set05]. Results [Nob00, ZZ04]. Retrieval [KKK^{+}01, LB01]. Retrieve [EKWMM02]. Reverse [Coo02]. Reversible [Ric08, TC03, TC04, YSHP08]. Reversing [BEQ00, YWC08]. Review [Agg02, AR06, Alh03, Alo04, Ama02a, Ama02b, Amb02, Arc02, Aye03, Azi06, BI06, Bas03, Ber02a, Ber02b, BA04, Bob02, Bon02, Bon03, Bot02, Bot03, Bur06a, Bur06b, Car06, Cas06, hC02b, Che05, Chu02, Con04, Dam03, De 04a, De 03, Dec02, Dud06, Fit04, For02, FM02b, Gaz06, Gei09, Ger05, Gla03, Gom06, Gro02, Grü07, HM06, Har07, Har09b, Har09a, HOM08, Hib03, Hir04, Hol05, Hro03, Hul05, Hun03a, Hun03b, Iwa02, Jas06, Kel04, Ker04, Kir02, Lam04, Lan03, Leo03, Lev06b, Lho04, Lim02, Lio03a, Lio03b, Loc02, Lom03, Lui03, Mac03, Mah06, Mar09, Mar05, Men02, Meo03a, Meo03b, Mic03, Mis03, Mis04b, Mis04a, MD09, Mon02a, Mon02b, Mor06, Mot02, Mun02, Nap02, O’K02, O’L05, Out02, Pal06a, Pal06b, Pan05, Par05, Phi04, Poe06, Pol06]. Review [Pom02, Pon02, Por03a, Por03b, Pre02, Ref09, Roc04, San02b, Sap03, Sch04, Sch06, Sch07, See04, Seh04, Ser03, Set05, Smi02, Smi06, Sof05, Tar02, Tim03, Uzu04, Vak06, Val02, VR04, Weg03, Zim03, Zit02, Ano06a]. Reviewers [Ano06a]. Reviews [Ano02a, Ano02b, Ano02c, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano04a, Ano04b, Ano04c, Ano05b, Ano06b, Kam04a, Kam04b, Kam05a, Kam05b, Kam05c, Kam05d, Kam06a, Kam06b, Kam06c, Kam06d, Kam06e, Kam07a, Kam07b, Kam07c, Kam07d, Kam07e, Kam08a, Kam08b, Kam08c, Kam09a, Kam09b, Kam09c, TVT02, Kam07f]. Revised [Cho07]. Revision [Li07, LB01]. Revisited [GP07]. Rewarded [Car03, Car08]. Rewriting [Jou04]. RFID [LKCK09]. Rick [Pol06]. Ring [CKM03, YMB09]. Rings [CDPV05, LL07]. Riolo [Pol06]. RISC [Jas06, Dan05]. Rivals [MH06]. Robert [HM06, Pal06a, Roc04]. Robin [LC07, Ano06c]. Robust [CJ03, CWY05, PSR04, PS05, Tse05]. Rogaway [Cho07]. Roger [Mur03]. Role [PSS09]. Role-Based [PSS09]. Rotations [BO05, Luc04]. Rothlauf [Mah06]. Round [CC00a, LC07]. Round-based [CC00a]. Round-Robin [LC07]. Route [Meg09]. Routing [ASDOK01, BDV07, EY04, Eom08, Gav00, GO06, HGS^{+}05, LMOP00, Meg09, RSW02, TH05, Wu02]. Rules [CL05]. Rummikub [HH06]. Run [CV04]. Rys [Vak06].

S [Alo04, Bre08b, Dow08, Gei09, Grü07, HM06, Vak06, VR04]. S. [Jas06, Vak06]. Saake [Lam04]. Safe [LBR00]. Safety [BCKM01, TJ01]. Safety-Critical [BCKM01, TJ01]. SAM [HD02]. Sampled [ELDJ05, VHE08, GGA09]. Samuel [Hir04]. Sangiorgi [Lho04]. Sara [Gom06]. sat [Mar00]. Satisfaction [GS08]. Saving [ACL03]. Scalable [BHJ06]. Scale [BQ03, FFPS06, OFW^{+}05, YWW09]. Scale-Free [YWW09]. Scenarios [MS07]. Scheduled [BF04]. Scheduler [RW06]. Scheduler* [DGT04]. Schedules [DGPT01]. Scheduling [ADA00, BCGH05, CT07, DT06, DÖ05, HPCC^{+}04, KL02a, Li08, PGT01, SAS02, Sub05]. Scheduling-Based [CT07]. Schema [For04b, For08]. Schemas [DFH^{+}05, TS05]. Scheme [AK02, AA05, BEQ00, HNZI02, LS06, LMOP00, SSNGS00, Wu02]. Schemes [BSC03, CL05, CDD06, Joh00, WH04, YWC08]. Scherer [Har09b]. Schiller [Phi04]. Schmidt [JV07]. Schneier [See04]. Schröder [KL03]. Science [Bur06a, Kuz05, Lev06b, Pal06b, Poe06, Pol06, Smi06, Vak06]. Scientific [HOM08]. Screen [GVB^{+}09]. SCTP [HTT06]. SDH [Set05]. SDI [TZ06b]. Search [ADA00, BO05, BI06, CDPV06, Lev06b, Lev06a, SJ04]. Searching [Lev06b, SJ04]. Second [Hul05, Mah06, GV07a]. Secret [JLL02]. Section [Man03]. Secure [CW09, HSHI06, Hor01, HL07, RF05, Tse07]. Security [DBS^{+}06, Gut04, Har07, OC02, Pip05, See04, Sne06, Tan07, Uzu04, WS03, Sch03b]. Segment [Cha05a]. Segment-Page [Cha05a]. Selection [Cha05b, Ipa09, LWY07, NPL07, SA09, SG00, Tam00]. Selectively [Chi08]. Self [AO07, BO05, BBQV07, BDV07, DGT04, LCBK04, PPK08, STH^{+}09, ZDL^{+}04, vM05]. Self-Adapting [PPK08]. Self-Adjusting [BO05]. Self-Maintaining [ZDL^{+}04]. Self-Management [vM05]. Self-Organized [STH^{+}09]. Self-Organizing [AO07, BBQV07]. Self-Stabilizing [BDV07, DGT04]. Self-Tuning [LCBK04]. Semantic [AA05, Ger05, LFB00, SC01, AV04, Col06]. Semantics [Gor02, GPLF05]. Semi [TZ06a]. Semi-Dynamic [TZ06a]. Sensibly [See04, Sch03b]. Sensing [SNLZ09]. sensitive [ZZC01]. Sensor [CMV08, DBS^{+}06, IVS06, SC08, TKLH04]. Sensors [LKCK09]. SEOF [AK00]. Separation [CK00, RB03]. Sequence [AJI07, CHL^{+}08, DHHG06, GO09]. Sequences [GGA09, Teu09]. Sequential [GBR01]. Series [Azi06, Bur06a, Lev07, Man07, Mor06, Pal06b, Poe06, Smi06, Vak06]. Serum [GVB^{+}09]. Server [Che05, JLH^{+}03]. Servers [CC00a, HS07, JLHD01, SAA06, TYH04]. Service [CCD07, CHG07, FS07a, FS07b, Hol05, IZ09, ITV04, KP04, LZ04a, ÖL07, SAA06, SF04, PQA^{+}03]. Services [Alo04, BCS02, BF00, Che05, CT07, FS07a, GHGJ^{+}07, Pan05, The03, GM05, ZTP03, Pal06a]. Set [Rin03, VD04]. SHA [AD07]. SHA-512 [AD07]. Shall [Ano06c, Mil06]. shaped [San02a]. Shaping [Gou03, Kel04]. Shared [MSK02]. Sharing [HNZI02, HH05]. Shel [HM06]. Shells [HP00]. Short [Lom02, MD09]. Shortest [SAS02]. Shortest-path [SAS02]. Shuffle [EM00]. SIENA [BBQV07]. Signal [SAS02]. Signature [SSNGS00, TNM00]. Signature-Tree [TNM00]. Signatures [CNV06, Her06]. Silver [O’L05]. SIMD [HOM08]. Similarity [FM02a, For08, RGW02]. Simion [JV07]. Simple [CJ03, CWY05, GPM02]. Simplification [Che04]. Simulation [BFGW02, ELDJ05, VHE08]. Single [ED09, ED10, HC02a, WMS01]. Single-faced [WMS01]. Siphons [Cha06]. situ [EM00]. Situation [SRS^{+}07]. Size [HNZI02]. sized [XZJ01]. Skarbek [Kor06b]. Skeleton [BCGH05]. Skeleton-Based [BCGH05]. Skills [Har09a, Tab06]. Slicing [DFH^{+}05]. Small [CCM05]. Smart [KBX^{+}04]. Smooth [LCBK04]. Snap [BDK07, CDPV06, HGDD05]. SNAP-Based [HGDD05]. Snap-Stabilizing [BDK07, CDPV06]. Snedaker [Har07]. Snob [JM08]. SOAP [Alo04, ZTP03]. Society [Lom02]. Softbound [BI06, Dud06, Gom06, Har07, Har09a, Pal06b, Sch07, Vak06]. Software [CC06, DGPT01, Eom08, FFPS06, Gal04, GO09, GBR01, GA05, Har09a, HD02, Lav09, Mis04b, XK01, Zha03, ZCHL09, ca09, KM02, Tab06, Ker04, Mis04a]. Solution [RW06]. Solving [CKM03, HH06, KMMS03]. Some [Bre08b, GY08, JV07, ZZ04]. SONET [Set05]. SONET-SDH [Set05]. Sons [Gaz06, HM06]. Sorting [GS03, SSL08]. Soundness [VvdAtH07]. Source [MV05]. Space [Bal05, BLMM06, HGS^{+}05, YWB07]. Space-Aware [BLMM06]. Spanning [PL02]. Spärck [Ano07e]. Sparse [FSM09]. Spatial [BNM03, CPK09, LN07, LV06, XZJ01]. Spatial-join [XZJ01]. Spatio [TVM00]. Spatio-Temporal [TVM00]. Spatiotemporal [Man03]. Special [Che01, DFL08, GTV07, LZ07, Man03, ZCHL09]. Specification [LMR04, Par02, Sch07, ZZC01, FJ05]. Specifications [HD02, Rin03, RTV08]. Specified [KL02b]. Specify [Mis04b, KM02]. Specifying [Sub05]. Spector [AR06]. Speed [Koc06, LLL07, Li03, LSS02]. Speed-up [Li03, LSS02]. Spiked [GT08]. Spink [Lev06b]. Splay [Geo04]. Split [MM00, Yen07]. Springer [Azi06, Bur06a, Cas06, Grü07, Jas06, Mah06, Mar09, Mor06, Pal06b, Pol06, Sch06, Smi06, Vak06]. Springer-Proceedings [Pal06b]. Springer-Verlag [Jas06, Mah06, Mor06, Sch06, Smi06, Vak06]. SRAM [Jas07]. SRAM-Based [Jas07]. St [Lav09]. Stabilizing [BDK07, BDV07, CDPV06, DGT04]. Stable [EM00]. Stack [SJ05]. Staerk [Roc04]. Standards [Ang05, VBL^{+}05]. Standards-Based [VBL^{+}05]. Stanley [Pal06b]. State [Cha05b, Eom08, Ipa09, KK07, KK09, LL07, Roc04, BS03, DFGI01, Hie03]. State-Optimal [LL07]. Statechart [HMP^{+}02]. Statecharts [Hua06]. Static [DFH^{+}05, DT06]. Statistical [RS03, Grü07]. Statistics [LLZ05, Mur02, ZDL^{+}04].

Stefano [Gom06]. Steps [CW09]. Stepwise [TZ06b]. Steve [CY08]. Stochastic [BBD01]. stop [SSNGS00]. Storing [CV03]. Strange [MM07]. Strategies [TKLH04]. Strategy [EY04, SA09, WPMK07]. Stream [BBD01, Ipa03, Ipa05]. Streaming [HTT06]. Strengthen [UA02]. Strict [FW02]. Strings [BIM^{+}07]. Strongly [KL02a]. Structural [For08, LBR00]. Structurally [SZ09]. Structure [LLH00, Ris06]. Structured [RS03]. Structures [Mar08a]. Structuring [AB03]. Struts [For04a]. Studies [FS07b]. Study [Dav01, HGDD05, LPS^{+}08, SZ09, XCWG06]. style [SC01]. Subarray [BT06]. Subgraphs [RGW02]. Subscribe [BBQV07]. Subsystem [HL07]. Suggestion [Hia04]. sum [Iqb08]. Summary [MBI^{+}05]. Suonio [Sch06]. Super [Bur05, CMCS09, ED09, ED10, Gre09, MD09, PCRZ09, RFM09, SNLZ09, SFC09, VMMK09, Smi06]. Super-recursive [Bur05, Smi06]. Super-Resolution [CMCS09, ED09, ED10, Gre09, MD09, PCRZ09, RFM09, SNLZ09, SFC09, VMMK09]. Superimpositions [SK03]. Support [FWLC04, GHNS04, Koc06, PS03, Wu04]. Supported [Par05]. Supporting [CT07, GA05, IZ09]. Survey [Cal06, GA05, YLW09]. Susan [Har07]. Switching [Dav01]. Sydney [Bre08a]. Symbolic [CMS02]. Symmetric [Alb04]. Symmetrical [TC03, Sta02]. Symmetry [CT08, Lev05]. SymOntos [FM02a]. Symposium [BCH^{+}05, Vak06]. Synchronization [JCL05, PSR04, WH04]. Synchronized [CN01, Cha03, GT08]. Synchronous [LL07, ZLB01]. Syngress [Har07, Sne06]. Synthesis [ADA00, Cha05b, EB02, YSHP08]. Syntropy [LFB00]. System [BFS00, BF04, DM05, De 04b, DNM05, FFPS06, JCL05, JL07, LMR04, LCBK04, Len06, Li07, Roc04, SRS^{+}07, Sta02, ZDL^{+}04, BS03]. Systems [AK00, ADL^{+}03, AB03, BQR01, BQ03, BCS02, BA04, CMS02, CMTK05, Che01, CV03, CC06, CHG07, CDW09, CK00, Col06, DÖ05, DGPT01, Dud06, FS07a, FS07b, Fit04, Gaz06, GCDB01, Har09b, Hil05, Hua06, HMP^{+}02, JCT02, Jou04, KR02, KBX^{+}04, KL02a, Kon01, Kor00, LSBB04, LS06, LSS02, MC05, Pal06b, PL02, PPK08, PS03, PKSA06, PSS09, Syr06, TWN08, TJ01, WH04, XK01, ZSP^{+}04, Cra03, DL04, EP04, GSG07, KS05, Lan05, Mum03, vWJM06, Sch04, Sch06, Seh04].

T [Har09b, Lan06, Mor06, Pal06a]. Tabaka [Har09a]. Tabu [ADA00]. Tags [LKCK09]. Talk [HM06, SI06]. Tan [Mor06]. Tapestry [For04a]. Task [DT06, KL02a]. Tasks [Li08]. Taxonomy [MC05]. TEA [WN95]. Teaching [GA05]. Technical [SRS^{+}07]. Technique [BSC03, DH05, GPPR03, ZMSH08]. Techniques [Che04, DBS^{+}06, Had04, HNW08, Jas07, OR02, Pet07, Rus06, Seg07, ST08, TGBS05]. Technologies [Par05, Seg07, Vak06, BCH^{+}05]. Technology [CY08, Fur08, GM09, Lev07]. Telecommunications [XK01]. Template [Geo04]. Temporal [Ala00, LSY05, Ngu01, Pap02, SSL07, TVM00, Xu04]. Ten [The03]. Terese [Jou04]. Term [RW06, Jou04]. Termination [NLGP07]. Ternary [BO05, Sta02]. Tessellations [SFL03]. Test [AO09, ABC^{+}07, CQX^{+}09, GO09, HU08, Ipa09, KK07, KK09, SZ09]. Test-bed [ABC^{+}07]. Testability [HHF05]. Testbed [Len06]. Testing [CDW09, DHHG06, GO09, Hie03, HU08, KK07, Kon01, Man07, SZ09, YLW09, ZCHL09]. Texts [Bur06a, Poe06]. Their [Che04, STH^{+}09, Tho05b, HOM08, HiOSG08, MSK02, WMS01]. Theorem [Poe06, BC04]. Theorems [Sol08]. Theoretical [Bur06a, Poe06, ZZ04]. Theory [Bur06b, DH08, FS07a, GK08, Geo04, Ker04, Lho04, Mis04a, Pol06, Sch06, WB04, Dal05, KS05, OYRW05, SW01, Zha03]. There [CM07]. Thinking [Jor07, See04, Sch03b]. Third [BCH^{+}05, Vak06, Hoa07]. Thrashing [RW06]. Thread [NS02]. Three [CW09, HGDD05, KK09, Sol08]. Three-Party [CW09]. Three-Phase [HGDD05]. Threshold [GB03]. Tilings [Lev05]. Time [AK02, Car03, Car08, CV04, CNV06, DT06, DÖ05, HPCC^{+}04, HSM09, JWLQ09, KHL03, SSS04, SSL08, Sub05, SC08, XU01, YWB07, ZLB01, ADL^{+}03, KL02a, Kon01, yLKTL00, PL02, Rei04]. Time-Free [CNV06]. Time-Limited [AK02]. Timed [GB03]. Timely [TGSR07]. Timestamp [LLS02]. Timestamps [Ngu01]. Tina [Pol06]. tiny [WN95]. TLS [ADA00]. Token [CDPV05, GB03]. Tolerance [Jas07]. Tolerant [BEQ00, BFS00, GCDB01, Had04, MSK02, ASDOK01, Wu02]. Tomas [Rei04]. Tomlinson [Alo04]. Tony [Ano07b, Ano07c]. Tool [FM02a, Kor00, Mar00, SZ09]. Toolkit [Cal07]. Tools [BFGW02, GA05, YLW09]. Topics [Car06, Pie05]. Topologies [HM03]. Tori [Li03]. Torus [WTSS01]. Torus/Mesh [WTSS01]. Tournament [Teu09]. Tower [Sap04]. Trace [Gor02]. Traceback [WLLW06]. Tracey [Gei09]. Tracking [BBvST01, KP02, NLGP07, PF01, TKLH04]. Trade [DÖ05]. Trade-off [DÖ05]. Trading [ITV04]. Traffic [Hua06, SAOKM00, Tof02, WTSS01, WLLW06]. Training [Tam00]. Transaction [BQR01]. Transactions [yLKTL00]. Transcriptomic [LPS^{+}08]. Transform [Car03, Car08, CPChX04, MSB00]. Transformation [HHF05]. Transient [Car03, Car08]. Transition [CMS02, JCL05]. Transition-Reduction [JCL05]. Transitive [GPPR03]. Translational [Lev05]. Transmission [Bal02a].

Transparency [ABPB06]. Transport [AACP09]. Transpose [SAOKM00]. Tree [HiOSG08, LPC09, PL02, RS03, San02a, SM02, TNM00, BNM03]. Tree-shaped [San02a]. Tree-width [HiOSG08]. Treelike [Yen07]. Trees [AND07, Bal02a, Bal02b, KL02b, KL03, Kor05, Kor06b, Luc04, MM00, RSW02, XUA00, YMB09, CV05]. Treewidth [BK08]. Trends [Don04, GM09, Mur02]. Tries [BO05]. Trojans [Coh01, Mäk01, TAC01, TAC98]. Trondheim [BCH^{+}05, Vak06]. Trunk [LPC09]. Tully [ca09]. Tuning [Hil05, LCBK04]. Turing [Don01, Jor07, Mai07, Pip05, Ran00]. Twenty [Hir04, Lom02]. Twenty-First [Hir04, Lom02]. Two [CLH07, LCT01]. Two-Dimensional [LCT01]. Two-Phase [CLH07]. Type [For04b, KR02, Vir03]. Type-Passing [Vir03]. Types [Car06, JMT02, Pie05]. Typical [LBI07].

Ubiquitous [Ano06c, IZ09, Mil06]. UDDI [Alo04, ZTP03]. Ugly [Pip05]. UK [Lav09]. UML [Dud06, AJI09, CQX^{+}09, HMP^{+}02, KM02, Lan05, RTV08, Mis04b]. UMTS [HL07]. Una-May [Pol06]. Unbreakable [Ver06]. Uncertain [See04, Sch03b]. Uncertainty [KKK^{+}01]. Unconditionally [HSHI06]. Understand [Ano06c, Mil06]. Understanding [DL04, Seh04]. Unified [DT06, GBR01, LH01]. Uniform [LL07, XZJ01]. Unique [DHHG06]. Universal [Har08, Sol03, Sol08]. University [Ano07e, Gei09, Sch07, ca09]. Unland [Vak06]. Unreliable [YWW09]. Unstructured [HHF05]. Unsupervised [JM08]. Untangling [Luc04]. Update [JLH^{+}03]. Updated [Cho08]. Upon [UA02]. Upper [Koc06]. US$46.95 [Jas06]. US$59.95 [Jas06]. Usability [Har08]. Usage [HOM08]. Use [AJI07, Bal05, BCKM01, LPS^{+}08, PF09]. Use-Case [AJI07]. User [AJI09, CT08, KP02, LCBK04, LWY07, Owe03, WH04]. User-Centered [Owe03]. User-Defined [LCBK04]. User-Oriented [LWY07]. Users [Har08, KK06]. Using [AO09, BO05, BCGH05, BLMM06, CW08, CMV08, EKWMM02, FSM09, FS07b, GS03, GO06, HGDD05, HTT06, Hua06, HLL09, LLZ05, LLS02, LSS02, Mar09, NLGP07, PT00, RFM09, RTV08, RVB06, Rus06, TC03, Xu04, BBD01, CKM03, CV05, HC02a, RGW02, Rin03, RSW02, SAS02, SSS04, VBvdA01].

V. [Ano07a]. Validating [Fur02, Rin03]. Validation [IVS06, YLK06]. Values [GB03]. Variable [PSR04, PS05, RFM09, TC03, TC04]. Various [BIM^{+}07]. VCR [WH04]. VCR-Like [WH04]. Velocity [LLZ05, For04a]. Verifiable [JLL02]. Verification [ABDS01, BLMM06, CV04, Gut04, Kor00, TJ01, Uzu04]. Verifying [BFS00, Rus06, VvdAtH07]. Verlag [Jas06, Mah06, Mor06, Sch06, Smi06, Vak06]. versus [Col06, MH06]. Vertex [Tad08]. Vertex-ordering [Tad08]. Vertical [SG00]. Very [Bun06]. Via [SSL07, Luc04]. Viable [ZWP03]. Video [AA05, GHNS04, HTT06, HLL09, KF05, LLH00, Sof05, SA09, TYH04]. Video-on-Demand [GHNS04, HLL09]. Video-Oriented [AA05]. Views [SJ02]. Virus [Coh01, Mäk01, TAC98, TAC01]. Viruses [ZZ04]. Visual [CDD06, YWC08, ZZC01]. Visualization [BMNL00, Ker04, Zha03]. VLSI [CBW08, ZL03]. Vol [Lev06b, Pal06b, Vak06]. Volume [Ano07d, Ano08, BB05]. Voting [JLL02, PKSA06]. Vovk [Ano07a].

W. [Gaz06]. Wait [PT00]. Wait-Free [PT00]. Walker [Lho04]. Wallace [Grü07, Bre08b, Dow08, JM08]. Walnut [CPK08]. Warehouse [BMN03, Cal07].

Warmup [ELDJ05, VHE08]. was [ML06]. Washington [Lom02]. Wave [BDK07]. Wavelet [MSB00]. Way [EM00, HM06, AK02, SI06]. WDM [CKM03]. Weak [UA02]. Wearable [RMJ04]. Web [Alo04, BI06, Che05, Ger05, Gom06, Lev06b, Mar05, Pal06a, Pal06b, ZSP^{+}04, AV04, BMN03, CFB^{+}03, Col06, For04a, GM05, JLHD01, JLH^{+}03, Jor07, KKK^{+}01, Lev06b, Lev06a, Man07, Seg07, SJ04, ZSP^{+}04, ZTP03]. Web-Server [JLH^{+}03]. WebWorks [For04a]. Wegner [CM07]. Weighted [LC07]. Weights [AO09]. Wesley [BI06, Har09a]. Wezel [Gaz06]. WGSN [Wu04, FWLC04]. Wheeler [Bar05, CK05]. who [Ano06a]. WHOWEDA [BMNL00]. Wide [KKK^{+}01]. Wideband [CMTK05]. Width [HiOSG08]. Wiley [Gaz06, HM06, Har09b]. Wilkes [Ano08]. William [Har09b]. Wired [HTT06]. Wireless [CG02, CT07, CMV08, HTT06, KP02, LLZ05, LKCK09, Sch04, TKLH04, EP04]. Wires [CL01]. Wiretapping [Cho08]. Wirsing [Cas06]. WISE [Pal06b, ZSP^{+}04]. without [CL01, CNV06]. WLAN [Wu04, FWLC04]. WLAN-based [Wu04, FWLC04]. Woflan [VBvdA01]. Wolsey [Mar09]. Word [Luk01]. Workflow [Puu01, SCJ^{+}05, VBvdA01]. Workflows [VvdAtH07]. Working [Ang05]. Workshop [CHJ^{+}02]. World [Alo04, Mai07, Sch03b, ZTP03, KKK^{+}01, Che05, See04]. Wormhole [WTSS01]. Worst [CCM05, JWLQ09, ZLB01]. Worst-Case [JWLQ09, ZLB01]. Worzel [Pol06]. Wrappers [GS03]. Write [KHL03]. Write-Through [KHL03]. WSDL [Alo04, ZTP03].

X [Bur06b, HM06, Mah06, Ipa03, Ipa05]. X-machines [Ipa03, Ipa05]. X.680 [Rin03]. Xiaofang [Pal06b]. XML [BCH^{+}05, Vak06, For04b, For08, KP04].

XML-based [KP04]. XML-Schema [For04b, For08]. XSym [BCH^{+}05, Vak06]. Xtra [Mis04b, KM02]. Xtra-Light [Mis04b, KM02].

Yahalom [Cho07]. years [ML06]. Yu [Pol06]. Yves [Mar09, Poe06].

Z [Vak06]. Zermelo [ML06]. Zero [Iqb08]. Zero-sum [Iqb08]. Zhang [Ker04]. Zhou [Pal06b]. Zimmermann [Alo04]. zone [TYH04]. Zoned [TH05]. Zooming [FSM09].

References

Agius:2005:CVO

[AA05] Harry Agius and Marios C. Angelides. COSMOS-7: Videooriented MPEG-7 scheme for modelling and filtering of semantic content. The Computer Journal, 48(5): 545–562, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/5/545;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/5/545.

Anastasi:2009:DPE

[AACP09] Giuseppe Anastasi, Emilio Ancillotti, Marco Conti, and Andrea Passarella. Design and performance evaluation of a transport protocol for ad hoc networks. The Computer Journal, 52(2):186–209, March 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/2/186;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/2/186;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/2/186.

Amor:2001:DPF

[AAL^{+}01] M. Amor, F. Argüello, J. López, O. Plata, and E. L. Zapata. A data-parallel formulation for divide and conquer algorithms. The Computer Journal, 44(4): 303–320, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_04/440303.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_04/pdf/
440303.pdf.

Anastasi:2003:SRM

[AB03] Giuseppe Anastasi and Alberto Bartoli. On the structuring of reliable multicast protocols for distributed mobile systems. The Computer Journal, 46(2):146–160, February 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_02/460146.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_02/pdf/
460146.pdf.

Ancillotti:2007:LFI

[ABC^{+}07] E. Ancillotti, R. Bruno, M. Conti, E. Gregori, and A. Pinizzotto. A layer-2 framework for interconnecting ad hoc networks to fixed Internet: Test-bed implementation and experimental evaluation. The Computer Journal, 50(4):478–499, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/4/478;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/4/478;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/4/478.

Anastasi:2001:EVM

[ABDS01] Giuseppe Anastasi, Alberto Bartoli, Nicoletta De Francesco, and Antonella Santone. Efficient verification of a multicast protocol for mobile computing. The Computer Journal, 44(1): 21–30, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_01/440021.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_01/pdf/
440021.pdf.

Abe:2004:CEP

[Abe04]  Masayuki Abe.  Combin-

ing encryption and proof of knowledge in the random oracle model. The Computer Journal, 47(1):58– ??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_01/470058.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_01/pdf/
470058.pdf.

Amor:2006:HOA

[ABPB06] M. Amor, M. Bóo, E. J. Padrón, and D. Bartz. Hardware oriented algorithms for rendering order-independent transparency. The Computer Journal, 49(2):201–210, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/2/201;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/2/201;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/2/201.

Anastasi:2003:PSN

[ACL03] Giuseppe Anastasi, Marco Conti, and Willy Lapenna. A power-saving network architecture for accessing the Internet from mobile computers: Design, implementation and measurements. The Computer Journal, 46(1):3– 15, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460003.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460003.pdf.

Ahmad:2007:ADE

[AD07] Imtiaz Ahmad and A. Shoba Das. Analysis and detection of errors in implementation of SHA-512 algorithms on FP-GAs. The Computer Journal, 50(6):728–738, November 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/6/728;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/6/728;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/6/728.

Ahmad:2000:TTS

[ADA00] Imtiaz Ahmad, Muhammad K.

Dhodhi, and Faridah M. Ali. TLS: a tabu search based scheduling algorithm for behavioral synthesis of functional pipelines. The Computer Journal, 43(2):152–166, ???? 2000. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_02/430152.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/

                                        hdb/Volume_43/Issue_02/pdf/
                                        430152.pdf.

Adil:2008:ECM

[Adi08] Baykasoglu Adil. Evolutionary computation for modeling and optimization. The Computer Journal, 51(6):743, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/51/
                                        6/743;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/51/6/743.

Alvarez:2003:OOM

[ADL^{+}03] J. M. Alvarez, M. Diaz, L. Llopis, E. Pimentel, and J. M. Troya. An objectoriented methodology for embedded real-time systems. The Computer Journal, 46(2):123– 145, February 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_02/460123.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_02/pdf/
                                        460123.pdf.

Assis:2006:MIH

[AFB06] Carlos A. G. Assis, Edil S. T. Fernandes, and Valmir C. Barbosa. Modeling the input history of programs for improved instruction-memory performance. The Computer Journal, 49(6):744–761, November 2006. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/6/744;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/6/744;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/6/744.

Aggarwal:2002:BR

[Agg02] Arvind Aggarwal. Book review. The Computer Journal, 45(3):375, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450375.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450364.pdf.

Aycock:2002:PEP

[AH02] John Aycock and R. Nigel Horspool. Practical Earley parsing. The Computer Journal, 45(6): 620–630, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_06/450620.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_06/pdf/
450620.pdf.

Almendros-Jimenez:2007:DUC

[AJI07] Jesús M. Almendros-Jiménez and Luis Iribarne. Describing use-case relationships with sequence diagrams. The Computer Journal, 50 (1):116–128, January 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/1/116;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/1/116;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/1/116.

Almendros-Jimenez:2009:UMU

[AJI09] Jesús M. Almendros-Jiménez and Luis Iribarne. UML modeling of user and database interaction. The Computer Journal, 52(3):348–367, May 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/3/348;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/3/348;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/3/348.

Ahn:2000:DSA

[AK00] Jung-Ho Ahn and Hyoung-Joo Kim. Dynamic SEOF: An adaptable object prefetch policy for object-oriented database systems. The Computer Journal, 43(6):524–537, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_43/Issue_

06/430524.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_43/Issue_
06/pdf/430524.pdf.

Abe:2002:KES

[AK02] Masayuki Abe and Masayuki Kanda. A key escrow scheme with time-limited monitoring for one-way communication. The Computer Journal, 45(6): 661–671, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_06/450661.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_06/pdf/
450661.pdf.

Alagic:2000:TOO

[Ala00] Suad Alagić. Temporal objectoriented programming. The Computer Journal, 43(6):491– 511, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_06/430491.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_06/pdf/
430491.pdf.

Albini:2004:DSG

[Alb04] L. C. P. Albini. Diagnosis of symmetric graphs under the BGM model. The Computer Journal, 47(1):85– ??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_01/470085.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_01/pdf/
470085.pdf.

Alhir:2003:BR

[Alh03] Sinan Si Alhir. Book review. The Computer Journal, 46(4): 461–462, July 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_04/460461a.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_04/pdf/
460461a.pdf.

Alonso:2004:BRB

[Alo04] Gustavo Alonso. Book review: Perspectives on Web Services — Applying SOAP, WSDL and UDDI to Real-World Projects, by O. Zimmermann and M. Tomlinson and S. Peuser. The Computer Journal, 47(4):505, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470505.pdf;   http://
www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
04/470505.sgm.abs.html.

Amatya:2002:BRa

[Ama02a] Ananda Amatya. Book review. The Computer Journal, 45(2):252, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_02/450252.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_02/pdf/
450249.pdf.

Amatya:2002:BRb

[Ama02b] Ananda Amatya. Book review. The Computer Journal, 45(4):477, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_04/450477.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_04/pdf/
450477.pdf.

Ambler:2002:BR

[Amb02] Scott W. Ambler. Book review. The Computer Journal, 45(3):377, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450377.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450377.pdf.

Ahrabian:2007:PGA

[AND07] H. Ahrabian and A. Nowzari-Dalini. Parallel generation of t-ary trees in A-order. The Computer Journal, 50 (5):581–588, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/5/581;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/5/581;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/5/581.

Angelides:2005:EWM

[Ang05] Marios C. Angelides. Editorial: Working with multimedia standards: MPEG-7, MPEG-21 and H.264. The Computer Journal, 48 (5):502–503, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/reprint/48/5/502.

Anonymous:2002:CRa

[Ano02a] Anonymous. Capsule reviews. The Computer Journal, 45(4): 379–380, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_04/450379.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_04/pdf/
450379.pdf.

Anonymous:2002:CRb

[Ano02b] Anonymous. Capsule reviews.

The Computer Journal, 45(5): 479–480, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_05/450479.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_05/pdf/
450479.pdf.

Anonymous:2002:CRc

[Ano02c] Anonymous. Capsule reviews.

The Computer Journal, 45(6): 581–582, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_06/450581.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_06/pdf/
450581.pdf.

Anonymous:2003:CRa

[Ano03a] Anonymous. Capsule reviews.

The Computer Journal, 46(1): 1–2, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460001.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460001.pdf.

Anonymous:2003:CRb

[Ano03b]  Anonymous.  Capsule re-

views. The Computer Journal, 46(2):121–122, February 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_02/460121.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_02/pdf/
460121.pdf.

Anonymous:2003:CRc

[Ano03c] Anonymous. Capsule reviews. The Computer Journal, 46(3): 227–228, May 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_03/460227a.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_03/pdf/
460227a.pdf.

Anonymous:2003:CRd

[Ano03d] Anonymous. Capsule reviews. The Computer Journal, 46(4): 337–339, July 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_04/460337.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_04/pdf/
460337.pdf.

Anonymous:2003:CRe

[Ano03e]  Anonymous.  Capsule re-

views. The Computer Journal, 46(5):465–466, September 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_05/460465.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_05/pdf/
460465.pdf.

Anonymous:2003:CRf

[Ano03f]  Anonymous.  Capsule re-

views. The Computer Journal, 46(6):591–592, November 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_06/460591.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_06/pdf/
460591.pdf.

Anonymous:2004:CRa

[Ano04a] Anonymous. Capsule reviews.

The Computer Journal, 47(3): 275–277, May 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470275.pdf;   http://
www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
03/470275.sgm.abs.html.

Anonymous:2004:CRb

[Ano04b]  Anonymous.  Capsule re-

views. The Computer Journal, 47(5):509–510, September 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/free_
                                        pdf/470509.pdf;   http://
                                        www3.oup.co.uk/computer_
                                        journal/hdb/Volume_47/Issue_
                                        05/470509.sgm.abs.html.

Anonymous:2004:CRc

[Ano04c]  Anonymous.  Capsule re-

views. The Computer Journal, 47(6):625–626, November 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/free_
                                        pdf/470625.pdf;   http://
                                        www3.oup.co.uk/computer_
                                        journal/hdb/Volume_47/Issue_
                                        06/470625.sgm.abs.html.

Anonymous:2005:AR

[Ano05a]  Anonymous.  Acknowledge-

ment of referees. The Computer Journal, 48(2):254–256, March 2005. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_48/Issue_02/pdf/bxh094.
                                        pdf.

Anonymous:2005:CR

[Ano05b] Anonymous. Capsule reviews. The Computer Journal, 48(1):

1–3, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_01/bxh069.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_01/pdf/
bxh069.pdf.

Anonymous:2006:ARR

[Ano06a]  Anonymous.  Acknowledge-

ment of referees: Reviewers who have agreed to review between 1 January 2005 and 31 December 2005 for The Computer Journal. The Computer Journal, 49(3):378–380, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
3/378;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/3/378.

Anonymous:2006:CR

[Ano06b]  Anonymous.  Capsule re-

views. The Computer Journal, 49(2):133, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
2/133;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/2/133.

Anonymous:2006:DRM

[Ano06c] Anonymous. Discussion on Robin Milner’s First Computer Journal Lecture: Ubiquitous computing: Shall we understand it? The Computer Journal, 49(4):390–399, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
4/390;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/4/390.

Anonymous:2007:DHP

[Ano07a] Anonymous. Discussion on “Hedging Predictions in Machine Learning” by A. Gammerman and V. Vovk. The Computer Journal, 50(2):164– 172, March 2007. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/cgi/content/
full/50/2/164;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/2/164. See

[GV07a, GV07b].

Anonymous:2007:DIPa

[Ano07b] Anonymous. Discussion on “The Ideal of Program Correctness” by Tony Hoare. The Computer Journal, 50(3): 261–268, May 2007. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/cgi/content/
full/50/3/261;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/3/261. See

[Hoa07, Ano07c].

Anonymous:2007:DIPb

[Ano07c] Anonymous. Discussion on “The Ideal of Program Correctness”: Responses from Tony Hoare. The Computer Journal, 50(3):269–273, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
3/269;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/3/269.

See [Hoa07, Ano07b].

Anonymous:2007:EOV

[Ano07d] Anonymous. Editorial on the occasion of volume 50 of The Computer Journal. The Computer Journal, 50(1):4–6, January 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
1/4;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/1/4.

Anonymous:2007:KSJ

[Ano07e] Anonymous. Karen Spärck Jones (1935–2007): Professor (emeritus) of Computers and Information, University of Cambridge. The Computer Journal, 50(4):375–376, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
4/375;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/4/375.

Anonymous:2008:WBP

[Ano08]  Anonymous.  The Wilkes

Best Paper Award in volume 49 (2006). The Computer Journal, 51(1), January 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/reprint/51/1/NP.

Amati:2000:CGF

[AO00] Gianni Amati and Iadh Ounis. Conceptual graphs and first order logic. The Computer Journal, 43(1):1–12, 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_43/Issue_01/430001.
                                        sgm.abs.html.

Amer:2007:NFS

[AO07] Abdelrahman Amer and B. J. Oommen. A novel framework for self-organizing lists in environments with locality of reference: Lists-on-lists. The Computer Journal, 50(2):186–196, March 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/2/186;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/2/186;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/2/186.

Abdurazik:2009:UCB

[AO09] Aynur Abdurazik and Jeff Offutt. Using coupling-based weights for the class integration and test order problem. The Computer Journal, 52(5):557–570, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/5/557;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/5/557.

Al-Rabadi:2006:BRL

[AR06] Anas N. Al-Rabadi. Book review: Lee Spector, Automatic Quantum Computer Programming: A Genetic Programming Approach, Kluwer Academic Publishers (2004). ISBN 1-4020-7894-3. 153 pp. The Computer Journal, 49 (1):129–130, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
1/129;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/1/129.

Archenhold:2002:BR

[Arc02] Geoffrey Archenhold. Book review. The Computer Journal, 45(3):376–377, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_

                                        03/450376.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_45/Issue_
                                        03/pdf/450376.pdf.

Arnavut:2004:IC

[Arn04] Ziya Arnavut. Inversion coding. The Computer Journal, 47(1): 46–??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_01/470046.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_01/pdf/
                                        470046.pdf.

Al-Sadi:2001:PBF

[ASDOK01] J. Al-Sadi, K. Day, and M. Ould-Khaoua. Probabilitybased fault-tolerant routing in hypercubes. The Computer Journal, 44(5):368–373, ???? 2001. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_44/Issue_05/440368.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_05/pdf/
                                        440368.pdf.

Antoniou:2004:SWP

[AV04] G. (Grigoris) Antoniou and Frank Van Harmelen. A Semantic Web primer. Cooperative information systems. MIT Press, Cambridge, MA, USA, 2004. ISBN 0-262-01210-3. xx + 238 pp. LCCN TK5105.88815 . A58 2004.

Ayesh:2003:BR

[Aye03]  Aladdin Ayesh.  Book re-

view. The Computer Journal, 46(4):462, July 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_04/460462.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_04/pdf/
460462.pdf.

Azizi:2006:BRJ

[Azi06] Mostafa Azizi. Book review:

John Hunt and Chris Loftus, Guide to J2EE: Enterprise Java. Series Springer Professional Computing. Springer (2003). ISBN 1-85233-704-

4. $79.95/£45.00/£69.95. 672 pp. The Computer Journal, 49(2):250–251, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
2/250;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/2/250.

Bjorn-Andersen:2004:BRR

[BA04] Niels Bjørn-Andersen. Book review: Redesigning Human Systems — InfoSCI 2003 by Enid Mumford. The Computer Journal, 47(2):271, March 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/free_
                                        pdf/470271.pdf;   http://
                                        www3.oup.co.uk/computer_
                                        journal/hdb/Volume_47/Issue_
                                        02/470271.sgm.abs.html.

Balakirsky:2002:BCA

[Bal02a] Vladimir B. Balakirsky. Block codes for asynchronous data transmission designed from binary trees. The Computer Journal, 45(2):243–248, ???? 2002. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_02/450243.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_02/pdf/
                                        450243.pdf.

Balakirsky:2002:NCA

[Bal02b] Vladimir B. Balakirsky. A new coding algorithm for trees. The Computer Journal, 45(2):237– 242, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_02/450237.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_02/pdf/
                                        450237.pdf.

Balakirsky:2005:HDU

[Bal05] Vladimir B. Balakirsky. Hashing of databases with the use of metric properties of the Hamming space. The Computer Journal, 48(1):4– 16, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_01/bxh059.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_01/pdf/
bxh059.pdf.

Barbuti:2004:AIF

[Bar04]  Roberto Barbuti.  Analyz-

ing information flow properties in assembly code by abstract interpretation. The Computer Journal, 47(1):25– ??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_01/470025.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_01/pdf/
470025.pdf.

Barron:2005:DWP

[Bar05]  David W. Barron.  David

Wheeler: a personal memoir. The Computer Journal, 48(6):650–651, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/48/
6/650;   http://comjnl.oxfordjournals.
org/cgi/reprint/48/6/650.

Baskerville:2003:BR

[Bas03] Richard Baskerville. Book review. The Computer Journal, 46(6):660, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_46/Issue_
                                        06/460660a.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_46/Issue_
                                        06/pdf/460660a.pdf.

Brimkov:2005:AHG

[BB05] Valentin E. Brimkov and Reneta P. Barneva. Analytical honeycomb geometry for raster and volume graphics. The Computer Journal, 48(2): 180–199, March 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_48/Issue_02/bxh075.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_48/Issue_02/pdf/
                                        bxh075.pdf.

Bowman:2001:AMS

[BBD01] H. Bowman, J. W. Bryans, and J. Derrick. Analysis of a multimedia stream using stochastic process algebra. The Computer Journal, 44(4):230– 245, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_44/Issue_04/440230.

sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_04/pdf/
440230.pdf.

Baldoni:2007:EPS

[BBQV07] Roberto Baldoni, Roberto Beraldi, Leonardo Querzoni, and Antonino Virgillito. Efficient publish/subscribe through a self-organizing broker overlay and its application to SIENA. The Computer Journal, 50(4):444–459, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/4/444;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/4/444;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/4/444.

Baggio:2001:ETM

[BBvST01] Aline Baggio, Gerco Ballintijn, Maarten van Steen, and Andrew S. Tanenbaum. Efficient tracking of mobile objects in Globe. The Computer Journal, 44(5):340–353, ???? 2001. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_05/440340.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_05/pdf/
440340.pdf.

Bertot:2004:ITP

[BC04] Yves Bertot and P. (Pierre) Castéran. Interactive theorem proving and program development: Coq’Art: the calculus of inductive constructions. Texts in theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-20854-2 (hardcover). xxv + 469 pp. LCCN QA76.9.A96 B47 2004.

Benoit:2005:SSB

[BCGH05] A. Benoit, M. Cole, S. Gilmore, and J. Hillston. Scheduling skeleton-based grid applications using PEPA and NWS. The Computer Journal, 48(3):369–378, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/3/369;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        48/3/369.

Bressan:2005:DXT

[BCH^{+}05] Stéphane Bressan, Stefano Ceri, Ela Hunt, Zachary G. Ives, Zohra Bellahsène, Michael Rys, and Rainer Unland, editors. Database and XML technologies: Third International XML Database Symposium, XSym 2005, Trondheim, Norway, August 28– 29, 2005: proceedings, volume 3671 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CO-DEN LNCSD9. ISBN 3-540-28583-0 (softcover). ISSN 0302-9743. LCCN QA76.9.D3.I5978

2005.  URL http://www.
springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=3671.

Bate:2001:UMP

[BCKM01] Iain Bate, Philippa Conmy, Tim Kelly, and John McDermid. Use of modern processors in safety-critical applications. The Computer Journal, 44(6): 531–543, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_06/440531.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_06/pdf/
440531.pdf.

Bellavista:2002:JLD

[BCS02] Paolo Bellavista, Antonio Corradi, and Cesare Stefanelli. Java for on-line distributed monitoring of heterogeneous systems and services. The Computer Journal, 45(6):595– 607, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_06/450595.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/

                                        hdb/Volume_45/Issue_06/pdf/
                                        450595.pdf.

Bein:2007:OSS

[BDK07] Doina Bein, Ajoy K. Datta, and Mehmet Hakan Karaata. An optimal snap-stabilizing multiwave algorithm. The Computer Journal, 50(3):332–340, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/3/332;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/3/332;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/3/332.

Bein:2007:SSL

[BDV07] Doina Bein, Ajoy K. Datta, and Vincent Villain. Selfstabilizing local routing in ad hoc networks. The Computer Journal, 50(2):197–203, March 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/2/197;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/2/197;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/2/197.

Berkovich:2000:REC

[BEQ00] Simon Y. Berkovich and Eyas El-Qawasmeh. Reversing the error-correction scheme for a fault-tolerant indexing. The Computer Journal, 43(1):54– 64, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_01/430054.
sgm.abs.html.

Berleant:2002:BR

[Ber02a] Daniel Berleant. Book review.

The Computer Journal, 45(5): 578–579, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_05/450578.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_05/pdf/
450578.pdf.

Berry:2002:BR

[Ber02b] Simon Berry. Book review.

The Computer Journal, 45(3): 374, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450374.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450364.pdf.

Bettstetter:2004:CAH

[Bet04] Christian Bettstetter. On the connectivity of ad hoc networks. The Computer Journal, 47(4):432–447, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
04/470432.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
04/pdf/470432.pdf.

Borghuis:2000:CLS

[BF00] Tijn Borghuis and Loe Feijs. A constructive logic for services and information flow in computer networks. The Computer Journal, 43(4):274– 289, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_04/430274.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_04/pdf/
430274.pdf.

Bondavalli:2004:MAS

[BF04] Andrea Bondavalli and Roberto Filippini. Modeling and analysis of a scheduled maintenance system: a DSPN approach. The Computer Journal, 47(6):634– 650, November 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_06/470634.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_06/pdf/
470634.pdf.

Barringer:2002:RDS

[BFGW02] Howard Barringer, Donal Fellows, Graham Gough, and Alan Williams. Rainbow: Development, simulation and analysis tools for asynchronous micropipeline hardware design. The Computer Journal, 45(1): 2–11, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_01/450002.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_01/pdf/
450002.pdf.

Bernardeschi:2000:FVF

[BFS00] Cinzia Bernardeschi, Alessandro Fantechi, and Luca Simoncini. Formally verifying fault tolerant system designs. The Computer Journal, 43(3): 191–205, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_03/430191.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_03/pdf/
430191.pdf.

Blundo:2003:HII

[BG03] Carlo Blundo and Clemente Galdi. Hiding information in image mosaics. The Computer Journal, 46(2):202–212, February 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_02/460202.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_02/pdf/
460202.pdf.

Bousias:2006:ILP

[BHJ06] Kostas Bousias, Nabil Hasasneh, and Chris Jesshope. Instruction level parallelism through microthreading — a scalable approach to chip multiprocessors. The Computer Journal, 49(2):211–233, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/2/211;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/2/211;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/2/211.

Bar-Ilan:2006:BRM

[BI06]  Judit Bar-Ilan.  Book re-

view: Mark Levene, An Introduction to Search Engines and Web Navigation. Addison Wesley, Pearson Education (2006). ISBN 0-321-30677-5. £39.99. 365 pp. Softbound. The Computer Journal, 49(4):500, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
4/500-a;   http://comjnl.

oxfordjournals.org/cgi/reprint/
49/4/500-a.

Bakalis:2007:LMM

[BIM^{+}07] A. Bakalis, C. S. Iliopoulos, C. Makris, S. Sioutas, E. Theodoridis, A. Tsakalidis, and K. Tsichlas. Locating maximal multirepeats in multiple strings under various constraints. The Computer Journal, 50(2):178–185, March 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/2/178;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/2/178;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/2/178.

Bodlaender:2008:COG

[BK08] Hans L. Bodlaender and Arie M. C. A. Koster. Combinatorial optimization on graphs of bounded treewidth. The Computer Journal, 51(3):255–269, May 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/3/255;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/3/255;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/3/255.

Bodlaender:2004:ACG

[BKTV04] Hans L. Bodlaender, Ton Kloks, Richard B. Tan, and Jan Van Leeuwen. Approximations for λ-colorings of graphs. The Computer Journal, 47(2): 193–204, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_02/470193.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_02/pdf/
470193.pdf.

Bernardeschi:2006:UCD

[BLMM06] C. Bernardeschi, G. Lettieri, L. Martini, and P. Masci. Using control dependencies for space-aware bytecode verification. The Computer Journal, 49(2):234–248, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/2/234;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/2/234;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/2/234.

Bhowmick:2003:RWD

[BMN03] Sourav S. Bhowmick, Sanjay Madria, and Wee Keong Ng. Representation of Web data in a Web warehouse. The Computer Journal, 46(3):229– 262, May 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/

Volume_46/Issue_03/460229.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_03/pdf/
460229.pdf.

Bhowmick:2000:DVO

[BMNL00] Sourav S. Bhowmick, Sanjay Madria, Wee Keong Ng, and Ee-Peng Lim. Data visualization operators for WHOWEDA. The Computer Journal, 43(5):364–385, ???? 2000. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_05/430364.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_05/pdf/
430364.pdf.

Bozanis:2003:LTL

[BNM03] Panayiotis Bozanis, Alexandros Nanopoulos, and Yannis Manolopoulos. LR-tree: a logarithmic decomposable spatial index method. The Computer Journal, 46(3):319– 331, May 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_03/460319.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_03/pdf/
460319.pdf.

Badr:2005:SAT

[BO05] Ghada Hany Badr and B. John Oommen. Self-adjusting of ternary search tries using conditional rotations and randomized heuristics. The Computer Journal, 48(2):200– 219, March 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_48/Issue_02/bxh073.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_48/Issue_02/pdf/
                                        bxh073.pdf.

Bobbio:2002:BR

[Bob02] Andrea Bobbio. Book review. The Computer Journal, 45(6): 688–??, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_06/450688.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_06/pdf/
                                        450688.pdf.

Bonifati:2002:BR

[Bon02] Angela Bonifati. Book review. The Computer Journal, 45(4): 476, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_04/450476a.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/

hdb/Volume_45/Issue_04/pdf/
450476a.pdf.

Bono:2003:BR

[Bon03] Viviana Bono. Book review.

The Computer Journal, 46(4): 463–464, July 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_04/460463.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_04/pdf/
460463.pdf.

Botta:2002:BR

[Bot02] Marco Botta. Book review.

The Computer Journal, 45(6): 687–688, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_06/450687a.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_06/pdf/
450687a.pdf.

Botta:2003:BR

[Bot03] Marco Botta. Book review.

The Computer Journal, 46(1): 117, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460117a.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/

hdb/Volume_46/Issue_01/pdf/
460117a.pdf.

Bataineh:2003:ROI

[BQ03] Sameer Bataineh and Ghassan E. Qanzu’a. Reliable omega interconnected network for large-scale multiprocessor systems. The Computer Journal, 46(5):467–475, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_05/460467.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_05/pdf/
460467.pdf.

Baldoni:2001:CCT

[BQR01] Roberto Baldoni, Francesco Quaglia, and Michel Raynal. Consistent checkpointing for transaction systems. The Computer Journal, 44(2):92– 100, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_02/440092.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_02/pdf/
440092.pdf.

Brennan:2008:DPE

[Bre08a] M. H. Brennan. Data processing in the early cosmic ray experiments in Sydney. The Computer Journal, 51 (5):561–565, September 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/5/561;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/5/561;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/5/561.

Brent:2008:SCC

[Bre08b]  Richard P. Brent.  Some

comments on C. S. Wallace’s random number generators. The Computer Journal, 51(5):579–584, September 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/5/579;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/5/579;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/5/579.

Borger:2003:ASM

[BS03] E. (Egon) Börger and Robert F.

Stärk. Abstract state machines: a method for high-level system design and analysis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-00702-4. x + 438 pp. LCCN QA76.9.S88.B66; QA76.9.S88.B66 2003.

Bhatia:2003:HTC

[BSC03] Randeep Bhatia, Rakesh K.

Sinha, and Chung-Min Chen. A hierarchical technique for constructing efficient declustering schemes for range queries. The Computer Journal, 46(4): 358–377, July 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_04/460358.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_04/pdf/
460358.pdf.

Bae:2006:IAM

[BT06] Sung Eun Bae and Tadao Takaoka. Improved algorithms for the K-maximum subarray problem. The Computer Journal, 49(3):358–374, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/3/358;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/3/358;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/3/358.

Bundy:2006:VMD

[Bun06] Alan Bundy. A very mathematical dilemma. The Computer Journal, 49(4):480–486, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/4/480;   http://comjnl.

oxfordjournals.org/cgi/content/
full/49/4/480;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/4/480.

Burgin:2005:SRA

[Bur05] M. S. (Mark Semenovich) Burgin. Super-recursive algorithms. Monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-95569-0. xii + 304 pp. LCCN QA9.615 .B87

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0662/2004041748-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0817/2004041748-
b.html;   http://www.loc.
gov/catdir/enhancements/
fy0817/2004041748-t.html.

Burgin:2006:BRJ

[Bur06a] Mark Burgin. Book review:

Juraj Hromkovič, Design and Analysis of Randomized Algorithms: Introduction to Design Paradigms. Series: Texts in Theoretical Computer Science. An EATCS Series. Springer (2005). ISBN 3-540-23949-9. £30.50/£39.95/$49.95. 274 pp. Hardbound. The Computer Journal, 49(2):249–250, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
2/249;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/2/249.

Burgin:2006:BRK

[Bur06b] Mark Burgin. Book review: Kimiz Dalkir, Knowledge Management in Theory and Practice. Butterworth Heinemann (2005). ISBN 0-7506-7864-X. EUR 44.95/$49.95/£29.99. 368 pp. Hardbound. The Computer Journal, 49(4):501–503, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/49/
                                        4/501-a;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        49/4/501-a.

checkedabstract:2009:APC

[ca09] Get checked abstract. Appreciation of Professor Colin Tully (1936–2007): Professor of Software Practice, Middlesex University. The Computer Journal, 52(3):388–391, May 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/3/388;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/52/3/388;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/52/3/388.

Cai:2008:PCC

[Cai08] Leizhen Cai. Parameterized complexity of cardinality constrained optimization problems. The Computer Journal, 51(1):102–121, January 2008. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/1/102;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/1/102;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/1/102.

Calamoneri:2006:LPS

[Cal06]  Tiziana Calamoneri.  The

L(h, k)-labelling problem: a survey and annotated bibliography. The Computer Journal, 49(5):585–608, September 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/5/585;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/5/585;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/5/585.

Caldeira:2007:MDW

[Cal07] Carlos Caldeira. The Microsoft Data Warehouse Toolkit. The Computer Journal, 50(2): 249, March 2007. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/cgi/content/
full/50/2/249;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/2/249.

Carrasco:2003:TAR

[Car03] Juan A. Carrasco. Transient analysis of rewarded continuous time Markov models by regenerative randomization with Laplace transform inversion. The Computer Journal, 46(1): 84–99, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460084.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460084.pdf.

Cardone:2006:BRB

[Car06] Felice Cardone. Book review: Benjamin C. Pierce (ed.), Advanced Topics in Types and Programming Languages, The MIT Press (2005) ISBN 0-262-16228-8. £39.95. 574 pp. Hardbound. The Computer Journal, 49(1):130–131, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
1/130;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/1/130.

Carrasco:2008:CTA

[Car08]  Juan A. Carrasco.  Corri-

gendum: “Transient Analysis of Rewarded Continuous Time Markov Models by Regenerative Randomization with Laplace Transform Inversion”. The Computer Journal, 51(6):745, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/51/
6/745;   http://comjnl.oxfordjournals.
org/cgi/reprint/51/6/745.

Casteran:2006:BRI

[Cas06] Pierre Castéran. Book review:

Iman Hafiz Poernomo, John Newsome Crossley and Martin Wirsing, Adapting Proofsas-Programs — The Curry– Howard Protocol. Springer (2005). ISBN 0-387-23759-3. $79.95/£50.00/ 64.95. 420 pp. Hardbound. The Computer Journal, 49(2):251–252, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
2/251;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/2/251.

Colon-Bonet:2008:MEF

[CBW08] Glenn Colón-Bonet and Paul Winterrowd, Jr. Multiplier evolution: a family of multiplier VLSI implementations. The Computer Journal, 51 (5):585–594, September 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/5/585;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/5/585;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/5/585.

Cheng:2000:DRB

[CC00a] Sheng-Tzong Cheng and Chia-Mei Chen. Dynamic roundbased resource management for on-demand multimedia servers. The Computer Journal, 43(5): 386–395, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_43/Issue_05/430386.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_43/Issue_05/pdf/
                                        430386.pdf.

Chung:2000:NFA

[CC00b] Kuo-Liang Chung and Hsiu-Niang Chen. A neighborfinding algorithm for bincodebased images on reconfigurable meshes. The Computer Journal, 43(4):315–324, ???? 2000. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_43/Issue_04/430315.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_43/Issue_04/pdf/
                                        430315.pdf.

Chen:2006:HAS

[CC06] Ing-Ray Chen and Bojan Cukic. High assurance software systems. The Computer Journal, 49(5):507–508, September 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
5/507;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/5/507.

Chan:2004:EDM

[CCCL04] Alvin T. S. Chan, Siu-Nam Chuang, Jiannong Cao, and Hong-Va Leong. An eventdriven middleware for mobile context awareness. The Computer Journal, 47(3):278–288, May 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Chen:2007:IDP

[CCD07] Zhongqiang Chen, Zhongrong Chen, and Alex Delis. An inline detection and prevention framework for distributed denial of service attacks. The Computer Journal, 50(1):7–40, January 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/1/7;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/1/7;   http://comjnl.
oxfordjournals.org/cgi/reprint/
50/1/7.

Choi:2005:JMA

[CCM05] Hyung-Kyu Choi, Yoo C.

Chung, and Soo-Mook Moon. Java memory allocation with lazy worst fit for small objects. The Computer Journal, 48(4):437–442, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/4/437;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/4/437.

Cimato:2006:PVC

[CDD06] S. Cimato, R. De Prisco, and A. De Santis. Probabilistic visual cryptography schemes. The Computer Journal, 49(1):97–107, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/1/97;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/1/97;   http://comjnl.
oxfordjournals.org/cgi/reprint/
49/1/97.

Chen:2004:EMP

[CDM04] Ing-Ray Chen, Luiz DaSilva, and Scott Midkiff. Editorial: Mobile and pervasive computing. The Computer Journal, 47 (4):404–??, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_04/470404.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_04/pdf/
470404.pdf.

Cantarell:2005:GME

[CDPV05] Sébastien Cantarell, Ajoy K. Datta, Franck Petit, and Vincent Villain. Group mutual exclusion in token rings. The Computer Journal, 48(2):239– 252, March 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_02/bxh077.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_02/pdf/
bxh077.pdf.

Cournier:2006:SSD

[CDPV06] Alain Cournier, Stéphane Devismes, Franck Petit, and Vincent Villain. Snap-stabilizing depth-first search on arbitrary networks. The Computer Journal, 49(3):268–280, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/3/268;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/3/268;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/3/268.

Chen:2009:PMT

[CDW09] Zhongqiang Chen, Alex Delis, and Peter Wei. A pragmatic methodology for testing intrusion prevention systems. The Computer Journal, 52(4):429–460, July 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.

                                        org/cgi/content/abstract/
                                        52/4/429;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        52/4/429.

Clark:2003:AOM

[CEK03] Tony Clark, Andy Evans, and Stuart Kent. Aspect-oriented metamodelling. The Computer Journal, 46(5):566–577, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_05/460566.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_05/pdf/
                                        460566.pdf.

Ceri:2003:DDI

[CFB^{+}03] Stefano Ceri, Piero Fraternali, Aldo Bongio, Marco Brambilla, Sara Comai, and Maristella Matera, editors. Designing data-intensive Web applications. Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, San Francisco, CA, USA and Amsterdam, The Netherlands, 2003. ISBN 1-55860-843-5. xxix + 561 pp. LCCN TK5105.888 .D465 2003. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        description/els031/2002114096.
                                        html;   http://www.loc.gov/
                                        catdir/toc/els031/2002114096.
                                        html.

Canal:2001:ECI

[CFP^{+}01] C. Canal, L. Fuentes, E. Pimentel, J. M. Troya, and A. Vallecillo. Extending CORBA interfaces with protocols. The Computer Journal, 44(5):448–462, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_44/Issue_
05/440448.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_44/Issue_
05/pdf/440448.pdf.

Chen:2002:CCA

[CG02] Ing-Ray Chen and Baoshan Gu. A comparative cost analysis of degradable location management algorithms in wireless networks. The Computer Journal, 45(3):304– 319, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450304.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450304.pdf.

Chao:2003:ESC

[Cha03] Daniel Yuh Chao. Extended synchronized choice nets. The Computer Journal, 46(5):505– 523, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_05/460505.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_05/pdf/
                                        460505.pdf.

Cha:2005:SPI

[Cha05a] Guang-Ho Cha. The segmentpage indexing method for large multidimensional queries. The Computer Journal, 48(1):101– 114, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_48/Issue_01/bxh057.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_48/Issue_01/pdf/
                                        bxh057.pdf.

Chattopadhyay:2005:ACS

[Cha05b] Santanu Chattopadhyay. Area conscious state assignment with flip-flop and output polarity selection for finite state machine synthesis — a genetic algorithm approach. The Computer Journal, 48(4):443–450, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/4/443;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        48/4/443.

Chao:2006:CES

[Cha06] Daniel Yuh Chao. Computation of elementary siphons in Petri nets for deadlock control. The Computer Journal, 49(4):470–479, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/4/470;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/4/470;   http://
comjnl.oxfordjournals.org/
cgi/content/full/bxl019/
DC1;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/4/470.

Chen:2001:SIE

[Che01] Ing-Ray Chen. Special issue editorial. high assurance systems. The Computer Journal, 44(6): 471–472, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_06/440471.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_06/pdf/
440471.pdf.

Chen:2004:BFS

[Che04] Zhongqiang Chen. Building footprint simplification techniques and their effects on radio propagation predictions. The Computer Journal, 47(1):103– ??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/

                                        Volume_47/Issue_01/470103.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_01/pdf/
                                        470103.pdf.

Che:2005:RBP

[Che05] Haoyang Che. Review: From P2P to Web Services and Grids: Peers in a Client/ Server World. The Computer Journal, 48(3):379, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/reprint/48/3/379-
                                        a.

Chen:2007:PBR

[CHG07] Ing-Ray Chen, Weiping He, and Baoshan Gu. Proxybased regional registration for integrated mobility and service management in mobile IP systems. The Computer Journal, 50(3):281–293, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/3/281;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/3/281;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/3/281.

Chien:2008:SCA

[Chi08] Hung-Yu Chien. Selectively convertible authenticated encryption in the random oracle model. The Computer Journal, 51(4):419–434, July 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/4/419;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/4/419;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/4/419.

Conrad:2002:REW

[CHJ^{+}02] S. Conrad, W. Hasselbring, A. James, D. Kambur, R.-D. Kutsche, and P. Thiran. Report on the EFIS 2001 Workshop. The Computer Journal, 45(2): 249–251, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_02/450249.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_02/pdf/
450249.pdf.

Cai:2008:PCB

[CHL^{+}08] Liming Cai, Xiuzhen Huang, Chunmei Liu, Frances Rosamond, and Yinglei Song. Parameterized complexity and biopolymer sequence comparison. The Computer Journal, 51(3):270–291, May 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/3/270;   http://comjnl.

                                        oxfordjournals.org/cgi/content/
                                        full/51/3/270;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/51/3/270.

Choo:2007:PRY

[Cho07] Kim-Kwang Raymond Choo. A proof of revised Yahalom protocol in the Bellare and Rogaway (1993) model. The Computer Journal, 50(5): 591–601, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/5/591;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/5/591;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/5/591.

Choo:2008:PLP

[Cho08] Kim-Kwang Raymond Choo. Privacy on the line: The politics of wiretapping and encryption, updated and expanded edition. The Computer Journal, 51(6):744, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/51/
                                        6/744;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/51/6/744.

Churchill:2002:BR

[Chu02] Elizabeth F. Churchill. Book review. The Computer Journal, 45(5):575–577, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_
05/450575.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_
05/pdf/450575.pdf.

Chien:2003:RSA

[CJ03] Hung-Yu Chien and Jinn-Ke Jan. Robust and simple authentication protocol. The Computer Journal, 46(2):193– 201, February 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_02/460193.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_02/pdf/
460193.pdf.

Cho:2000:IIS

[CK00] Eun-Sun Cho and Hyoung-Joo Kim. Interface/implementation separation mechanism for integrating object-oriented management systems and generalpurpose programming languages. The Computer Journal, 43(6):512–523, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_43/Issue_
06/430512.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_43/Issue_
06/pdf/430512.pdf.

Campbell-Kelly:2005:CDW

[CK05] Martin Campbell-Kelly. Celebration of David Wheeler (1927–2004). The Computer Journal, 48(6):646– 648, November 2005. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

                                        tronic). URL http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/48/6/646;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/48/6/646.

Chan:2003:SCP

[CKM03] Tak-Ming Chan, Sam Kwong, and K. F. Man. Solving the converter placement problem in WDM ring networks using genetic algorithms. The Computer Journal, 46(4):427– 448, July 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_04/460427.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_04/pdf/
                                        460427.pdf.

Chen:2001:LCC

[CL01] Guihai Chen and Francis C. M. Lau. Layout of the cube-connected cycles without long wires. The Computer Journal, 44(5):374–383, ???? 2001. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/

Volume_44/Issue_05/440374.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_05/pdf/
440374.pdf.

Chen:2004:OAM

[CL04] Shan-Tai Chen and Shun-Shii Lin. Optimal algorithms for 2 × n mastermind games — a graph-partition approach. The Computer Journal, 47(5):602– 611, September 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_05/470602.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_05/pdf/
470602.pdf.

Chang:2005:PHS

[CL05] Chin-Chen Chang and Chih-Yang Lin. Perfect hashing schemes for mining association rules. The Computer Journal, 48(2):168–179, March 2005. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_02/bxh074.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_02/pdf/
bxh074.pdf.

Clark:2003:DNI

[Cla03] Martin P. Clark. Data networks, IP, and the Internet: protocols, design, and operation. Wiley, New York, NY, USA, 2003. ISBN 0-470-84856-1. xvi + 848 pp. LCCN TK5105.5 .C545 2003.

URL http://www.loc.gov/
catdir/description/wiley039/
2002191041.html;   http:/
/www.loc.gov/catdir/toc/
wiley031/2002191041.html.

Chen:2007:TPO

[CLH07] Shan-Tai Chen, Shun-Shii Lin, and Li-Te Huang. A twophase optimization algorithm for mastermind. The Computer Journal, 50(4):435–443, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/4/435;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/4/435;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/4/435.

Crampton:2001:LACa

[CLO01a] Jason Crampton, George Loizou, and Greg O’Shea. A logic of access control. The Computer Journal, 44(1):54– 66, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_01/440054.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_01/pdf/
440054.pdf.

Crampton:2001:LACb

[CLO01b] Jason Crampton, George Loizou, and Greg O’Shea. A logic of access control. The Computer Journal, 44(2):137– 149, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_02/440137.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_02/pdf/
440137.pdf.

Cockshott:2007:TNM

[CM07] Paul Cockshott and Greg Michaelson. Are there new models of computation? reply to Wegner and Eberbach. The Computer Journal, 50(2):232–247, March 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/2/232;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/2/232;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/2/232.

Chen:2008:PIH

[CM08] Jianer Chen and Jie Meng.

On parameterized intractability: Hardness and completeness. The Computer Journal, 51(1):39–59, January 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/1/39;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/1/39;   http://comjnl.
oxfordjournals.org/cgi/reprint/
51/1/39.

Chatterjee:2009:APG

[CMCS09] Priyam Chatterjee, Sujata Mukherjee, Subhasis Chaudhuri, and Guna Seetharaman. Application of Papoulis– Gerchberg method in image super-resolution and inpainting. The Computer Journal, 52(1):80–89, January 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/1/80;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/1/80;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/1/80.

Calder:2002:MLF

[CMS02] M. Calder, S. Maharaj, and C. Shankland. A modal logic for full LOTOS based on symbolic transition systems. The Computer Journal, 45(1):55– 61, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_01/450055.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/

hdb/Volume_45/Issue_01/pdf/
450055.pdf.

Chan:2005:JGA

[CMTK05] T. M. Chan, K. F. Man, K. S.

Tang, and S. Kwong. A jumping gene algorithm for multiobjective resource management in wideband CDMA systems. The Computer Journal, 48(6):749–768, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/6/749;   http://comjnl.
oxfordjournals.org/cgi/content/
full/48/6/749;   http://
comjnl.oxfordjournals.org/
cgi/reprint/48/6/749.

Croce:2008:RPC

[CMV08] Silvio Croce, Francesco Marcelloni, and Massimo Vecchio. Reducing power consumption in wireless sensor networks using a novel approach to data aggregation. The Computer Journal, 51(2):227–239, March 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/2/227;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/2/227;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/2/227.

Chao:2001:LSC

[CN01] Daniel Y. Chao and Jose A.

Nicdao. Liveness for synchronized choice Petri nets. The Computer Journal, 44(2):124– 136, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_02/440124.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_02/pdf/
440124.pdf.

Correia:2006:CAB

[CNV06] Miguel Correia, Nuno Ferreira Neves, and Paulo Veríssimo. From consensus to atomic broadcast: Time-free Byzantineresistant protocols without signatures. The Computer Journal, 49(1):82–96, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/1/82;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/1/82;   http://comjnl.
oxfordjournals.org/cgi/reprint/
49/1/82.

Cohen:2001:RCF

[Coh01]  Fred Cohen.  Reply to

“Comment on ‘A Framework for Modelling Trojans and Computer Virus Infection’ by E. Mäkinen”. The Computer Journal, 44(4):326– 327, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_04/440326.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_04/pdf/

440326.pdf. See [TAC98, Mäk01, TAC01].

Colyer:2004:RLA

[Col04]  Adrian Colyer.  Ramnivas

Laddad: AspectJ in actionpractical aspect-oriented programming. The Computer Journal, 47(1):134–135, January 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470134a.pdf;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
01/470134a.sgm.abs.html.

Colomb:2006:FVM

[Col06]  Robert M. Colomb.  For-

mal versus material ontologies for information systems interoperation in the Semantic Web. The Computer Journal, 49(1):4–19, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/1/4;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/1/4;   http://comjnl.
oxfordjournals.org/cgi/reprint/
49/1/4.

Contreras:2004:BRB

[Con04] Luis M. Contreras. Book review: Data Networks, IP and the Internet: Protocols, Design and Operation, by M. P. Clark. The Computer Journal, 47(4): 507–??, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/free_
                                        pdf/470507.pdf;   http://
                                        www3.oup.co.uk/computer_
                                        journal/hdb/Volume_47/Issue_
                                        04/470507.sgm.abs.html.

Cook:2002:REJ

[Coo02] Jonathan J. Cook. Reverse execution of Java bytecode. The Computer Journal, 45(6):608– 619, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_06/450608.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_06/pdf/
                                        450608.pdf.

Christensen:2000:MAP

[CP00] Søren Christensen and Laure Petrucci. Modular analysis of Petri nets. The Computer Journal, 43(3):224– 242, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_43/Issue_03/430224.
                                        sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/
hdb/Volume_43/Issue_03/pdf/
430224.pdf.

Chen:2004:EPA

[CPChX04] Ling Chen, Yi Pan, Yixin Chen, and Xiao hua Xu. Efficient parallel algorithms for Euclidean distance transform. The Computer Journal, 47(6):694– 700, November 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_06/470694.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_06/pdf/
470694.pdf.

Castro:2008:PCW

[CPK08] Maurice D. Castro, Ronald D.

Pose, and Carlo Kopp. Passwordcapabilities and the Walnut Kernel. The Computer Journal, 51(5):595–607, September 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/5/595;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/5/595;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/5/595.

Choi:2009:HDR

[CPK09] Jongseong Choi, Min Kyu Park, and Moon Gi Kang. High dynamic range image reconstruction with spatial resolution enhancement. The Computer Journal, 52 (1):114–125, January 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/1/114;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/1/114;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/1/114.

Chen:2009:UAD

[CQX^{+}09] Mingsong Chen, Xiaokang Qiu, Wei Xu, Linzhang Wang, Jianhua Zhao, and Xuandong Li. UML activity diagram-based automatic test case generation for Java programs. The Computer Journal, 52(5):545–556, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/5/545;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/5/545.

Crabtree:2003:DCS

[Cra03] Andy Crabtree. Designing collaborative systems: a practical guide to ethnography. Computer supported cooperative work. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 1-85233-718-4. xi + 178 pp. LCCN QA76.9.S88 C725 2003.

Chen:2004:OOB

[CSS04] Chia-Mei Chen, Huey-Min Sun, and LihChyun Shu. Optimizing object-based multimedia delivery across bandwidth constrained networks. The Computer Journal, 47(5):579– 590, September 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_05/470579.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_05/pdf/
470579.pdf.

Cheng:2007:SBM

[CT07] Sheng-Tzong Cheng and Ming-Hung Tao. A schedulingbased medium access control protocol for supporting multi-class services in wireless networks. The Computer Journal, 50(2):204–216, March 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/2/204;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/2/204;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/2/204.

Cairns:2008:ASU

[CT08] Paul Cairns and Harold Thimbleby. Affordance and symmetry in user interfaces. The Computer Journal, 51 (6):650–661, November 2008.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/6/650;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/6/650;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/6/650.

Chen:2002:CDF

[CU02] Xiao Jun Chen and Hasan Ural. Construction of deadlock-free designs of communication protocols from observations. The Computer Journal, 45(2):162– 173, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_02/450162.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_02/pdf/
450162.pdf.

Chen:2003:CBA

[CV03] Ing-Ray Chen and Naresh Verma. A cost-based admission control algorithm for digital library multimedia systems storing heterogeneous objects. The Computer Journal, 46(6):645– 659, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_06/460645.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/

hdb/Volume_46/Issue_06/pdf/
460645.pdf.

Catuogno:2004:AKL

[CV04] Luigi Catuogno and Ivan Visconti. An architecture for kernel-level verification of executables at run time. The Computer Journal, 47(5):511– 526, September 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_05/470511.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_05/pdf/
470511.pdf.

Corral:2005:AAD

[CV05] Antonio Corral and Michael Vassilakopoulos. On approximate algorithms for distancebased queries using R-trees. The Computer Journal, 48(2): 220–238, March 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_02/bxh060.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_02/pdf/
bxh060.pdf.

Chomicki:2004:LEA

[CVS04] Jan Chomicki, Ron Van der Meyden, and Gunter Saake, editors. Logics for emerging applications of databases. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-00705-9. xv + 368 pp. LCCN QA76.9.D3 L6425 2004.

Cavagnino:2008:EAI

[CW08] D. Cavagnino and A. E. Werbrouck. Efficient algorithms for integer division by constants using multiplication. The Computer Journal, 51(4):470–480, July 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        51/4/470;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/51/4/470;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/51/4/470.

Chien:2009:PSP

[CW09] Hung-Yu Chien and Tzong-Chen Wu. Provably secure password-based three-party key exchange with optimal message steps. The Computer Journal, 52(6):646–655, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/6/646;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        52/6/646.

Chien:2005:NRS

[CWY05] Hung-Yu Chien, Ren-Chiun Wang, and Chou-Chen Yang. Note on robust and simple authentication protocol. The Computer Journal, 48(1):27– 29, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_01/bxh061.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_01/pdf/
bxh061.pdf.

Cheung:2008:CBL

[CY08] Peter Y. K. Cheung and Alex Yakovlev. Comments on the BCS Lecture “The Future of Computer Technology and its Implications for the Computer Industry” by Professor Steve Furber. The Computer Journal, 51(6):741–742, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/51/
6/741;   http://comjnl.oxfordjournals.
org/cgi/reprint/51/6/741.

Chen:2009:DPM

[CZCD09] Zhongqiang Chen, Yuan Zhang, Zhongrong Chen, and Alex Delis. A digest and pattern matching-based intrusion detection engine. The Computer Journal, 52(6):699–723, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/6/699;   http://comjnl.

                                        oxfordjournals.org/cgi/reprint/
                                        52/6/699.

Dalkir:2005:KMT

[Dal05] Kimiz Dalkir. Knowledge management in theory and practice. Elsevier/Butterworth Heinemann, Amsterdam, The Netherlands, 2005. ISBN 0-7506-7864-X. xv + 356 pp. LCCN HD30.2 .D354

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0624/2005045074-d.html;
                                         http://www.loc.gov/catdir/
                                        toc/fy0601/2005045074.html.

Damiani:2003:BR

[Dam03] Ferruccio Damiani. Book review. The Computer Journal, 46(5):525–526, September 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_05/460525a.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_05/pdf/
                                        460525a.pdf.

Dandamudi:2005:GRP

[Dan05] Sivarama P. Dandamudi. Guide to RISC processors: for programmers and engineers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-21017-2. xv + 387 pp. LCCN QA76.5 .D2515 2005.

Davies:2001:HSB

[Dav01] D. W. Davies. An historical study of the beginnings of packet switching. The Computer Journal, 44(3):152– 162, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_03/pdf/440152.
pdf.

Doyle:2006:SCK

[DBS^{+}06] Barry Doyle, Stuart Bell, Alan F. Smeaton, Kealan Mc-Cusker, and Noel E. O’Connor. Security considerations and key negotiation techniques for power constrained sensor networks. The Computer Journal, 49(4):443–453, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/4/443;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/4/443;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/4/443.

DeMartin:2003:BR

[De 03] Juan Carlos De Martin. Book review. The Computer Journal, 46(2):224–??, February 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_02/460224.

                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_02/pdf/
                                        460224.pdf.

DeLaMorena:2004:BRB

[De 04a] Javier Farreres De La Morena. Book review: The Functional Approach to Data Management-Modeling, Analyzing and Integrating Heterogeneous Data, edited by P. M. D. Gray and L. Kerschberg and P. J. H. King and A. Poulovassilis. The Computer Journal, 47(4):506– 507, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/free_
                                        pdf/470506.pdf;   http://
                                        www3.oup.co.uk/computer_
                                        journal/hdb/Volume_47/Issue_
                                        04/470506.sgm.abs.html.

DeOliveiraGuimar:2004:GLE

[De 04b] José De Oliveira Guimar. The Green language exception system. The Computer Journal, 47(6):651–661, November 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_06/470651.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_06/pdf/
                                        470651.pdf.

Decandia:2002:BR

[Dec02] Giuseppe Decandia. Book review. The Computer Journal, 45(4):476, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_04/450476b.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_04/pdf/
450476b.pdf.

DeFrancesco:2001:FAM

[DFGI01] N. De Francesco, A. Fantechi, S. Gnesi, and P. Inverardi. Finite approximations for model checking nonfinite-state processes. The Computer Journal, 44(2):109– 123, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_02/440109.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_02/pdf/
440109.pdf.

Danicic:2005:SPS

[DFH^{+}05] Sebastian Danicic, Chris Fox, Mark Harman, Rob Hierons, John Howroyd, and Michael R. Laurence. Static program slicing algorithms are minimal for free liberal program schemas. The Computer Journal, 48(6):737–748, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.

                                        org/cgi/content/abstract/
                                        48/6/737;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/48/6/737;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/48/6/737.

Downey:2008:CJS

[DFL08] Rodney G. Downey, Michael R. Fellows, and Michael A. Langston. The Computer Journal special issue on parameterized complexity: Foreword by the Guest Editors. The Computer Journal, 51(1):1–6, January 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/51/
                                        1/1;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/51/1/1.

Dohi:2001:ESR

[DGPT01] Tadashi Dohi, Katerina Goševa-Popstojanova, and Kishor Trivedi. Estimating software rejuvenation schedules in high-assurance systems. The Computer Journal, 44(6):473– 485, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_44/Issue_06/440473.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_06/pdf/
                                        440473.pdf.

Datta:2004:SSM

[DGT04] Ajoy K. Datta, Maria Gradinariu, and Sébastien Tixeuil. Self-stabilizing mutual exclusion under arbitrary scheduler*. The Computer Journal, 47(3):289–298, May 2004. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Daly:2005:TDP

[DH05] Charlie Daly and Jane Horgan. A technique for detecting plagiarism in computer code. The Computer Journal, 48(6):662–666, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/6/662;   http://comjnl.
oxfordjournals.org/cgi/content/
full/48/6/662;   http://
comjnl.oxfordjournals.org/
cgi/reprint/48/6/662.

Demaine:2008:BTA

[DH08] Erik D. Demaine and MohammadTaghi Hajiaghayi. The bidimensionality theory and its algorithmic applications. The Computer Journal, 51(3): 292–302, May 2008. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/cgi/content/
abstract/51/3/292;   http:
//comjnl.oxfordjournals.
org/cgi/content/full/51/
3/292;   http://comjnl.oxfordjournals.
org/cgi/reprint/51/3/292.

Derderian:2006:AUI

[DHHG06] Karnig Derderian, Robert M. Hierons, Mark Harman, and Qiang Guo. Automated unique input output sequence generation for conformance testing of FSMs. The Computer Journal, 49(3):331–344, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        49/3/331;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/49/3/331;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/49/3/331.

DInverno:2004:UAS

[DL04] Mark D’Inverno and Michael (Michael M.) Luck. Understanding agent systems. Springer series on agent technology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second revised and expanded edition, 2004. ISBN 3-540-40700-6. xviii + 240 pp. LCCN QA76.76.I58.D56 2004; QA76.76.I58.D56.

Dini:2006:CPA

[DLL06] Gianluca Dini, Giuseppe Lettieri, and Lanfranco Lopriore. Caching and prefetching algorithms for programs with looping reference patterns. The Computer Journal, 49(1):42–61, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/1/42;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/1/42;   http://comjnl.
oxfordjournals.org/cgi/reprint/
49/1/42.

DAuriol:2005:PLA

[DM05] Brian J. D’Auriol and Rajesh Molakaseema. A parameterized linear array with a reconfigurable pipelined bus system: LARPBS(p). The Computer Journal, 48(1):115– 125, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_01/bxh067.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_01/pdf/
bxh067.pdf.

Djordjevic:2005:MSE

[DNM05] J. Djordjevic, B. Nikolic, and M. Mitrovic. A memory system for education. The Computer Journal, 48 (6):630–641, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/6/630;   http://comjnl.
oxfordjournals.org/cgi/content/
full/48/6/630;   http://
comjnl.oxfordjournals.org/
cgi/reprint/48/6/630.

Dogan:2005:BSA

[DÖ05] Atakan Dogan and Füsun Özgüner. Biobjective scheduling algorithms for execution time–reliability trade-off in heterogeneous computing systems. The Computer Journal, 48(3):300–314, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/3/300;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        48/3/300.

Donofrio:2001:TML

[Don01] Nick Donofrio. Turing Memorial Lecture. The Computer Journal, 44(2):67–74, ???? 2001. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_44/Issue_02/440067.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_02/freepdf/
                                        440067.pdf.

Dongarra:2004:BLT

[Don04] Jack Dongarra. The Boole Lecture: Trends in high performance computing. The Computer Journal, 47(4):399– 403, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_04/470399.

sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_04/pdf/
470399.pdf.

Dowe:2008:FRC

[Dow08] David L. Dowe. Foreword re C. S. Wallace. The Computer Journal, 51(5): 523–560, September 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/51/
5/523;   http://comjnl.oxfordjournals.
org/cgi/reprint/51/5/523;
http://comjnl.oxfordjournals.
org/content/51/5/523.full.
pdf+html.

Demiroz:2006:STS

[DT06] Betul Demiroz and Haluk Rahmi Topcuoglu. Static task scheduling with a unified objective on time and resource domains. The Computer Journal, 49(6):731–743, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/6/731;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/6/731;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/6/731.

Duddy:2006:BRK

[Dud06] Keith Duddy. Book review:

Kevin Lano, Advanced Systems Design with Java, UML and MDA. Elsevier (2005). ISBN 0-7506-6496-7. £27.99/US$44.95. 378 pp. Softbound. The Computer Journal, 49(4):500–501, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
4/500;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/4/500.

Edwards:2002:BAH

[EB02] Doug Edwards and Andrew Bardsley. Balsa: An asynchronous hardware synthesis language. The Computer Journal, 45(1):12–18, ???? 2002. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_01/450012.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_01/pdf/
450012.pdf.

Elad:2009:EBR

[ED09] Michael Elad and Dmitry Datsenko. Example-based regularization deployed to superresolution reconstruction of a single image. The Computer Journal, 52(1):15–30, January 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/1/15;   http://comjnl.
oxfordjournals.org/cgi/content/

full/52/1/15;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/1/15.

Elad:2010:CEB

[ED10] Michael Elad and Dmitry Datsenko. Corrigendum: “Example-Based Regularization Deployed to Super-Resolution Reconstruction of a Single Image”. The Computer Journal, 53(7):1131, September 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/53/
7/1131;   http://comjnl.
oxfordjournals.org/cgi/reprint/

53/7/1131. See [ED09].

Ebert:2005:DMH

[EHKL05] Jean-Pierre Ebert, Daniel Hollos, Holger Karl, and Marc Löbbers. Does multi-hop communication reduce electromagnetic exposure? The Computer Journal, 48(1):72– 83, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_01/bxh062.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_01/pdf/
bxh062.pdf.

El-Khatib:2002:UDA

[EKWMM02] Hazem T. El-Khatib, M. Howard Williams, David H. Marwick, and Lachlan M. Mackinnon.

Using a distributed approach to retrieve and integrate information from heterogeneous distributed databases. The Computer Journal, 45(4):381– 394, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_04/450381.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_04/pdf/
450381.pdf.

Eeckhout:2005:BAE

[ELDJ05] Lieven Eeckhout, Yue Luo, Koen De Bosschere, and Lizy K. John. BLRL: Accurate and efficient warmup for sampled processor simulation. The Computer Journal, 48(4):451–459, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/4/451;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/4/451.

Ellis:2000:SSM

[EM00] John Ellis and Minko Markov. In situ, stable merging by way of the perfect shuffle. The Computer Journal, 43(1):40– 53, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/

Volume_43/Issue_01/430040.
sgm.abs.html.

Eom:2008:IDC

[Eom08] Hyeonsang Eom. Informationdynamics-conscious development of routing software: a case of routing software that improves link-state routing based on future linkdelay-information estimation. The Computer Journal, 51 (2):144–161, March 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/2/144;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/2/144;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/2/144.

Elliott:2004:MCW

[EP04] Geoffrey Elliott and Nigel Phillips. Mobile commerce and wireless computing systems. Pearson/Addison Wesley, Harlow, England, 2004. ISBN 0-201-75240-9. xviii + 532 pp. LCCN QA76.59 .E45 2004.

Elliman:2004:NIA

[EY04] David G. Elliman and Sherin M.

Youssef. A new intelligent agent-based strategy for constrained multiple destination routing problems. The Computer Journal, 47(6):708–727, November 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_06/470708.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_06/pdf/
470708.pdf.

Fernandez:2006:AIS

[FFPS06] Juan Fernández, Eitan Frachtenberg, Fabrizio Petrini, and José-Carlos Sancho. An abstract interface for system software on large-scale clusters. The Computer Journal, 49(4):454–469, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/4/454;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/4/454;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/4/454.

Fitzpatrick:2004:BRB

[Fit04] Geraldine Fitzpatrick. Book review: Designing Collaborative Systems: A Practical Guide to Ethnography, by Andy Crabtree. The Computer Journal, 47(3):395– 396, May 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470395a.pdf;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
03/470395a.sgm.abs.html.

Feijs:2005:FSD

[FJ05] L. M. G. (Loe M. G.) Feijs and H. B. M. Jonkers. Formal specification and design, volume 35 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-43592-7 (paperback). 335 pp. LCCN QA76.76.D47 F44

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0632/2006273427-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0632/2006273427-
t.html.

Formica:2002:CSS

[FM02a] Anna Formica and Michele Missikoff. Concept similarity in SymOntos: An enterprise ontology management tool. The Computer Journal, 45(6):583– 594, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_06/450583.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_06/pdf/
450583.pdf.

Franceschet:2002:BR

[FM02b] Massimo Franceschet and Angelo Montanari. Book review. The Computer Journal, 45(6): 683–685, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/

                                        Volume_45/Issue_06/450683a.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_06/pdf/
                                        450683a.pdf.

Fornaro:2002:BR

[For02] Claudio Fornaro. Book review. The Computer Journal, 45(4):473, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_04/450473a.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_04/pdf/
                                        450473a.pdf.

Ford:2004:AJW

[For04a] Neal Ford. Art of Java Web development: Struts, Tapestry, Commons, Velocity, JUnit, Axis, Cocoon, InternetBeans, WebWorks. Manning Publications, Greenwich, CT, USA, 2004. ISBN 1-932394-06-0. xxx + 593 pp. LCCN QA76.73.J38 F66 2004. URL http:/

                                        /www.loc.gov/catdir/bios/
                                        ipg051/2004271136.html;
                                        http://www.loc.gov/catdir/
                                        description/ipg051/2004271136.
                                        html.

Formica:2004:LXS

[For04b]  Anna Formica.  Legality

of XML-schema type hierarchies. The Computer Journal, 47(5):591–601, September 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_05/470591.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_05/pdf/
470591.pdf;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_05/pdf/
470591.pdfhttp://www3.oup.
co.uk/computer_journal/hdb/
Volume_47/Issue_05/470591.
sgm.abs.html.

Formica:2008:SXS

[For08]  Anna Formica.  Similarity

of XML-schema elements: a structural and information content approach. The Computer Journal, 51(2):240–254, March 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/2/240;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/2/240;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/2/240.

Fallah:2007:SACa

[FS07a] Mehran S. Fallah and Ahmad R. Sharafat. Service availability in concurrent systems — Part I: a theory of hierarchical services of interacting processes. The Computer Journal, 50(5):522–534, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/5/522;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/5/522;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/5/522.

Fallah:2007:SACb

[FS07b] Mehran S. Fallah and Ahmad R. Sharafat. Service availability in concurrent systems — Part II: Analysis and case studies using HSIP. The Computer Journal, 50 (5):535–554, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/5/535;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/5/535;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/5/535.

Fadili:2009:IZU

[FSM09] M. J. Fadili, J.-L. Starck, and F. Murtagh. Inpainting and zooming using sparse representations. The Computer Journal, 52(1):64–79, January 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/1/64;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/52/1/64;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        52/1/64.

Furber:2002:VAM

[Fur02]  Steve Furber.  Validating

the AMULET microprocessors. The Computer Journal, 45(1): 19–26, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_01/450019.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_01/pdf/
450019.pdf.

Furber:2008:FCT

[Fur08] S. Furber. The future of computer technology and its implications for the computer industry. The Computer Journal, 51(6):735–740, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/6/735;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/6/735;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/6/735.

Farr:2002:CSM

[FW02] G. E. Farr and C. S. Wallace.

The complexity of strict minimum message length inference. The Computer Journal, 45(3): 285–292, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450285.

                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_03/pdf/
                                        450285.pdf.

Feng:2004:WWB

[FWLC04] Vincent W.-S. Feng, Lin-Yi Wu, Yi-Bing Lin, and Whai-En Chen. WGSN: WLANbased GPRS support node with push mechanism. The Computer Journal, 47(4):405– 417, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_04/470405.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_04/pdf/
                                        470405.pdf.

Gomez-Albarran:2005:TLP

[GA05]  Mercedes  Gómez-Albarrán.

The teaching and learning of programming: a survey of supporting software tools. The Computer Journal, 48(2):130– 144, March 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_48/Issue_02/bxh080.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_48/Issue_02/pdf/
                                        bxh080.pdf.

Galin:2004:SQA

[Gal04]  Daniel Galin.  Software

quality assurance. Pearson Education Limited, Harlow, England, 2004. ISBN 0-201-70945-7. xxvi + 590 pp. LCCN QA76.76.Q35.G35; QA76.76.Q35.G35 2004.

Gavoille:2000:DIR

[Gav00] Cyril Gavoille. On the dilation of interval routing. The Computer Journal, 43(3):243– 249, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_03/430243.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_03/pdf/
430243.pdf.

Gazi:2006:BRW

[Gaz06] Boran Gazi. Book review:

W. Van Wezel, R. J. Jorna and A. M. Meystel, Planning in Intelligent Systems: Aspects, Motivations, and Methods. John Wiley and Sons, Inc. (2006). ISBN 0-471-73427-6. £64.95/ 91.70/US $110.00. 592 pp. Hardbound. The Computer Journal, 49 (5):629–630, September 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
5/629-a;   http://comjnl.
oxfordjournals.org/cgi/reprint/

49/5/629-a. See [vWJM06].

Gencata:2003:DAT

[GB03] Ayşegül Gençata and Feza Buzluca. Determining the asynchronous threshold values of the timed token medium access protocol. The Computer Journal, 46(4):421–426, July 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_04/460421.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_04/pdf/
                                        460421.pdf.

Garcia:2001:UML

[GBR01] Alessandro F. Garcia, Delano M. Beder, and Cecília M. F. Rubira. A unified meta-level software architecture for sequential and concurrent exception handling. The Computer Journal, 44(6):569– 587, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_44/Issue_06/440569.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_06/pdf/
                                        440569.pdf.

Grandoni:2001:EFT

[GCDB01] F. Grandoni, S. Chiaradonna, F. Di Giandomenico, and A. Bondavalli. Evaluation of fault-tolerant multiprocessor systems for high assurance applications. The Computer Journal, 44(6):544– 556, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_06/440544.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_06/pdf/
440544.pdf.

Geil:2009:BRT

[Gei09]  Olav Geil.  Book review:

Tracey Ho and Desmond S. Lun, Network Coding: An Introduction. Cambridge University Press (2008). ISBN-13 978-0-521-87310-9. 184 pp. £30/ $60. Hardcover. The Computer Journal, 52(4):510, July 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/reprint/52/4/510.

Gelenbe:2006:EBA

[Gel06] Erol Gelenbe. Editorial: Building adaptivity into computer networks. The Computer Journal, 49(2):136, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
2/136;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/2/136.

Georgakopoulos:2004:GTS

[Geo04] George F. Georgakopoulos.

Generalized template splay: a basic theory and calculus. The Computer Journal, 47(1):10– ??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_01/470010.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_01/pdf/
470010.pdf.

Geroimenko:2005:BRB

[Ger05] Vladimir Geroimenko. Book review: A Semantic Web Primer, by G. Antoniou and F. van Harmelen. The Computer Journal, 48(1):126, January 2005. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_01/bxh070.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_01/pdf/
bxh070.pdf.

Gevrekci:2009:RBS

[GGA09] Murat Gevrekci, Bahadir K. Gunturk, and Yucel Altunbasak. Restoration of Bayersampled image sequences. The Computer Journal, 52(1):1– 14, January 2009. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/cgi/content/
abstract/52/1/1;   http:

//comjnl.oxfordjournals.
org/cgi/content/full/52/
1/1;   http://comjnl.oxfordjournals.
org/cgi/content/full/bxm041/
DC1;   http://comjnl.oxfordjournals.
org/cgi/reprint/52/1/1.

Glaser:2000:PNP

[GHG00] Hugh Glaser, Pieter H. Hartel, and Paul W. Garratt. Programming by numbers: a programming method for novices. The Computer Journal, 43(4): 252–265, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_04/430252.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_04/pdf/
430252.pdf.

Ghanea-Hercock:2007:HNG

[GHGJ^{+}07] Robert Ghanea-Hercock, E. Gelenbe, Nicholas R. Jennings, Oliver Smith, David N. Allsopp, Alex Healing, Hakan Duman, Simon Sparks, Nishan C. Karunatillake, and Perukrishnen Vytelingum. Hyperion — next-generation battlespace information services. The Computer Journal, 50 (6):632–645, November 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/6/632;   http://comjnl.
oxfordjournals.org/cgi/content/

                                        full/50/6/632;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/6/632.

Glinos:2004:VDB

[GHNS04] N. Glinos, D. B. Hoang, C. Nguyen, and A. Symvonis. Video-on-demand based on delayed-multicast: Algorithmic support. The Computer Journal, 47(5):545–559, September 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_05/470545.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_05/pdf/
                                        470545.pdf.

Goddard:2001:MLB

[GJ01] Steve Goddard and Kevin Jeffay. Managing latency and buffer requirements in processing graph chains. The Computer Journal, 44(6):486– 503, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_44/Issue_06/440486.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_06/pdf/
                                        440486.pdf.

Gasarch:2008:IFP

[GK08] William Gasarch and Keung Ma Kin. Invitation to fixed-parameter algorithms — parameterized complexity theory — parameterized algorithmics: Theory, practice and prospects. The Computer Journal, 51(1):137–140, January 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/51/
1/137;   http://comjnl.oxfordjournals.
org/cgi/reprint/51/1/137.

Gray:2004:FAD

[GKKP04] Peter M. D. Gray, Larry Kerschberg, Peter J. H. King, and Alexandra Poulovassilis, editors. The functional approach to data management: modeling, analyzing, and integrating heterogeneous data. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-00375-4. xiii + 483 pp. LCCN QA76.9.D3.F845; QA76.9.D3.F845 2004.

Giannopoulos:2008:PCG

[GKW08] Panos Giannopoulos, Christian Knauer, and Sue Whitesides. Parameterized complexity of geometric problems. The Computer Journal, 51(3):372–384, May 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/3/372;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/3/372;   http://

                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/51/3/372.

Glaeser:2003:BR

[Gla03] Georg Glaeser. Book review. The Computer Journal, 46(3): 332–??, May 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_03/460332a.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_03/pdf/
                                        460332a.pdf.

Glushko:2005:DEA

[GM05] Robert J. Glushko and Tim McGrath. Document engineering: analyzing and designing documents for business informatics and Web services. MIT Press, Cambridge, MA, USA, 2005. ISBN 0-262-07261-0. xxiii + 703 pp. LCCN QA76.9.T48 G54 2005.

Greenfield:2009:IET

[GM09] Daniel Greenfield and Simon Moore. Implications of electronics technology trends for algorithm design. The Computer Journal, 52(6):690–698, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/6/690;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        52/6/690.

Gramm:2008:FPA

[GNT08] Jens Gramm, Arfst Nickelsen, and Till Tantau. Fixedparameter algorithms in phylogenetics. The Computer Journal, 51(1):79–101, January 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/1/79;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/1/79;   http://comjnl.
oxfordjournals.org/cgi/reprint/
51/1/79.

Gyorgy:2006:ARU

[GO06] András György and György Ottucsák. Adaptive routing using expert advice. The Computer Journal, 49(2):180–189, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/2/180;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/2/180;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/2/180.

Gallagher:2009:TSG

[GO09] Leonard Gallagher and Jeff Offutt. Test sequence generation for integration testing of component software. The Computer Journal, 52(5):514–529, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.

                                        org/cgi/content/abstract/
                                        52/5/514;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        52/5/514.

Gomez:2006:BRS

[Gom06]  Jaime Gomez.  Book re-

view: Stefano Ceri, Piero Fraternali, Aldo Bongio, Marco Brambilla, Sara Comai and Maristella Matera, Designing Data-Intensive Web Applications, Morgan Kaufmann Publishers (2003). ISBN 1-55860-843-5. $59.95/ 58.90. 562 pp. Softbound. The Computer Journal, 49(1):127, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/49/
                                        1/127;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/49/1/127.

Gordon:2002:RET

[Gor02] Michael J. C. Gordon. Relating event and trace semantics of hardware description languages. The Computer Journal, 45(1):27–36, ???? 2002. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_01/450027.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_01/pdf/
                                        450027.pdf.

Gitzel:2007:OEM

[GOS07] Ralf Gitzel, Ingo Ott, and Martin Schader. Ontological extension to the MOF metamodel as a basis for code generation. The Computer Journal, 50(1):93–115, January 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/1/93;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/1/93;   http://comjnl.
oxfordjournals.org/cgi/reprint/
50/1/93.

Gouge:2003:SIO

[Gou03] Ian Gouge. Shaping the IT organization: the impact of outsourcing and the new business model. Springer professional computing; Practitioner series, 1439-9245. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 1-85233-727-3. x + 210 pp. LCCN HD30.2.G68 2003; HD30.2.G68.

Georgakopoulos:2007:MDR

[GP07] George F. Georgakopoulos and Kostas Politopoulos. MAX-DENSITY revisited: a generalization and a more efficient algorithm. The Computer Journal, 50(3):348–356, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/

                                        50/3/348;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/3/348;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/3/348.

Grosky:2005:DES

[GPLF05] William Grosky, Nilesh Patel, Xin Li, and Farshad Fotouhi. Dynamically emerging semantics in an MPEG-7 image database. The Computer Journal, 48(5):536–544, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/5/536;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        48/5/536.

Gibbons:2002:EAS

[GPM02] Alan Gibbons, Ida Pu, and Muthu Muthukrishnan. Exact analyses of a simple heuristic employed in array compression. The Computer Journal, 45(2): 213–220, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_02/450213.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_02/pdf/
                                        450213.pdf.

Gibbons:2003:CGP

[GPPR03] Alan Gibbons, Aris Pagourtzis, Igor Potapov, and Wojciech Rytter. Coarse-grained parallel transitive closure algorithm: Path decomposition technique. The Computer Journal, 46(4): 391–400, July 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_04/460391.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_04/pdf/
460391.pdf.

Guerraoui:2007:AIC

[GR07] Rachid Guerraoui and Michel Raynal. The alpha of indulgent consensus. The Computer Journal, 50(1):53–67, January 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/1/53;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/1/53;   http://comjnl.
oxfordjournals.org/cgi/reprint/
50/1/53.

Greenspan:2009:SRM

[Gre09]  Hayit Greenspan.  Super-

resolution in medical imaging. The Computer Journal, 52(1):43–63, January 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/1/43;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/1/43;   http://comjnl.

                                        oxfordjournals.org/cgi/reprint/
                                        52/1/43.

Grosse:2002:BR

[Gro02] André Grosse. Book review. The Computer Journal, 45(4): 475, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_04/450475a.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_04/pdf/
                                        450475a.pdf.

Grumbach:2003:HID

[GRS03] Stéphane Grumbach, Philippe Rigaux, and Luc Segoufin. Handling interpolated data. The Computer Journal, 46 (6):664–679, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_46/Issue_
                                        06/460664.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_46/Issue_
                                        06/pdf/460664.pdf.

Grunwald:2007:BRC

[Grü07] Peter Grünwald. Book review: Christopher S. Wallace, Statistical and Inductive Inference by Minimum Message Length, Springer (2005). ISBN 038723795X. £46.00. 432 pp. Hardbound. The Computer Journal, 50(3):369–370, May 2007. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
3/369;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/3/369.

Gerbessiotis:2003:IEE

[GS03] Alexandros V. Gerbessiotis and Constantinos J. Siniolakis. Increasing the efficiency of existing sorting algorithms by using randomized wrappers. The Computer Journal, 46(5):498– 504, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_05/460498.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_05/pdf/
460498.pdf.

Gottlob:2008:FPA

[GS08] Georg Gottlob and Stefan Szeider. Fixed-parameter algorithms for artificial intelligence, constraint satisfaction and database problems. The Computer Journal, 51(3):303–325, May 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/3/303;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/3/303;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/3/303.

Gibson:2007:HDS

[GSG07] John E. Gibson, William T. Scherer, and William F. Gibson. How to do systems analysis. Wiley series in systems engineering and management. Wiley-Interscience, New York, NY, USA, 2007. ISBN 0-470-00765-6 (cloth). xxiii + 335 pp. LCCN T57.6 .G543

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0741/2006033568-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0741/2006033568-
                                        d.html;   http://www.loc.
                                        gov/catdir/toc/ecip072/2006033568.
                                        html.

Gelenbe:2008:SIS

[GT08] Erol Gelenbe and Stelios Timotheou. Synchronized interactions in spiked neuronal networks. The Computer Journal, 51(6):723–730, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        51/6/723;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/51/6/723;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/51/6/723.

Gelenbe:2007:EIS

[GTV07] Erol Gelenbe, Andrew Tilbrook, and Richard Vinter. Editorial: Introduction to the special issue. The Computer Journal, 50(6):630–631, November 2007. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
6/630;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/6/630.

Gutmann:2004:CSA

[Gut04] Peter Gutmann. Cryptographic Security Architecture: Design and Verification. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-95387-

6. xviii + 320 pp. LCCN QA76.9.A25 G88 2002.

Gammerman:2007:HPM

[GV07a] Alexander Gammerman and Vladimir Vovk. Hedging predictions in machine learning: The Second Computer Journal Lecture. The Computer Journal, 50(2):151–163, March 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/2/151;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/2/151;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/2/151. See

discussion [Ano07a, GV07b].

Gammerman:2007:RHP

[GV07b] Alexander Gammerman and Vladimir Vovk. Rejoinder: “Hedging Predictions in Machine Learning”. The Computer Journal, 50(2):173–177, March 2007. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
2/173;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/2/173.

See [GV07a, Ano07a].

Gammerman:2009:SPA

[GVB^{+}09] Alex Gammerman, Volodya Vovk, Brian Burford, Ilia Nouretdinov, Zhiyuan Luo, Alexey Chervonenkis, Mike Waterfield, Rainer Cramer, Paul Tempst, Josep Villanueva, Musarat Kabir, Stephane Camuzeaux, John Timms, Usha Menon, and Ian Jacobs. Serum proteomic abnormality predating screen detection of ovarian cancer. The Computer Journal, 52(3):326–333, May 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/3/326;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/3/326;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/3/326.

Gutin:2008:SPP

[GY08] Gregory Gutin and Anders Yeo. Some parameterized problems on digraphs. The Computer Journal, 51(3):363–371, May 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/

51/3/363;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/3/363;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/3/363.

Hadjicostis:2004:CTF

[Had04] Christoforos N. Hadjicostis.

Coding techniques for faulttolerant parallel prefix computations in Abelian groups. The Computer Journal, 47(3):329– 341, May 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Harrin:2007:BRS

[Har07] Elizabeth Harrin. Book review: Susan Snedaker, IT Security Project Management Handbook. Syngress (2006). ISBN 1-59749-076-8. 47.95/ £32.99/$59.95. 612 pp. Softbound. The Computer Journal, 50(1):129, January 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
1/129;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/1/129.

Harrin:2008:UUD

[Har08]  Elizabeth Harrin.  Univer-

sal usability: Designing computer interfaces for diverse users. The Computer Journal, 51(6):743–744, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.

                                        org/cgi/content/full/51/
                                        6/743-a;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        51/6/743-a.

Harrin:2009:BRJb

[Har09a] Elizabeth Harrin. Book review: Jean Tabaka, Collaboration Explained: Facilitation Skills for Software Project Leaders. Addison Wesley (2006). ISBN-13 978-0-321-26877-8. £31.99. 412 pp. Softbound. The Computer Journal, 52(6):725, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/reprint/52/6/725.

Harrin:2009:BRJa

[Har09b] Elizabeth Harrin. Book review: John E. Gibson, William T. Scherer and William F. Gibson, How to do Systems Analysis, Wiley (2007). ISBN-13 978-0-470-00765-5. £47.50. 360 pp. Hardcover. The Computer Journal, 52(4):510–511, July 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/reprint/52/4/510-
                                        a.

Huang:2002:DIP

[HC02a] Chung-Ming Huang and Yu-Chang Chao. Differentiated information presentations for heterogeneous I/O devices using a single markup language approach. The Computer Journal, 45(1):72– 87, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_01/450072.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_01/pdf/
450072.pdf.

Chan:2002:BR

[hC02b] Wai hung Chan. Book review.

The Computer Journal, 45(2): 254–255, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_02/450254.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_02/pdf/
450254.pdf.

He:2002:FDA

[HD02] Xudong He and Yi Deng. A framework for developing and analyzing software architecture specifications in SAM. The Computer Journal, 45(1):111– 128, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_01/450111.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_01/pdf/
450111.pdf.

Herranz:2006:DIB

[Her06] Javier Herranz. Deterministic identity-based signatures for partial aggregation. The Computer Journal, 49(3):322–330, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        49/3/322;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/49/3/322;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/49/3/322.

Haji:2005:SBC

[HGDD05] Mohammed H. Haji, Iain Gourlay, Karim Djemame, and Peter M. Dew. A SNAPbased community resource broker using a three-phase commit protocol: a performance study. The Computer Journal, 48(3):333–346, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/3/333;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        48/3/333.

He:2005:PER

[HGS^{+}05] Fei He, Ming Gu, Xiaoyu Song, Zhiwei Tang, Guowu Yang, and Lerong Cheng. Probabilistic estimation for routing space. The Computer Journal, 48(6):667–676, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/6/667;   http://comjnl.
oxfordjournals.org/cgi/content/
full/48/6/667;   http://
comjnl.oxfordjournals.org/
cgi/reprint/48/6/667.

Huang:2005:REA

[HH05] Chung-Ming Huang and Tz-Heng Hsu. A resource exchange architecture for peer-to-peer file sharing applications. The Computer Journal, 48(1): 53–71, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_01/bxh055.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_01/pdf/
bxh055.pdf.

Hertog:2006:SRP

[HH06] D. Den Hertog and P. B. Hulshof. Solving Rummikub problems by integer linear programming. The Computer Journal, 49(6):665–669, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/6/665;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/6/665;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/6/665.

Hierons:2005:BCT

[HHF05] R. M. Hierons, M. Harman, and C. J. Fox. Branchcoverage testability transformation for unstructured programs. The Computer Journal, 48(4):421–436, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/4/421;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        48/4/421.

Hiasat:2004:SFR

[Hia04] Ahmad A. Hiasat. A suggestion for a fast residue multiplier for a family of moduli of the form (2n − (2p ± 1)). The Computer Journal, 47(1): 93–??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_01/470093.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_01/pdf/
                                        470093.pdf.

Hibbert:2003:BR

[Hib03] Adrian Hibbert. Book review. The Computer Journal, 46(1): 115, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_01/460115a.
                                        sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460115a.pdf.

Hierons:2003:GCW

[Hie03]  R. M. Hierons.  Generat-

ing candidates when testing a deterministic implementation against a non-deterministic finite-state machine. The Computer Journal, 46(3):307– 318, May 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_03/460307.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_03/pdf/
460307.pdf.

Hillston:2005:TSC

[Hil05] Jane Hillston. Tuning systems: From composition to performance. The Computer Journal, 48(4):385–400, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/4/385;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/4/385.

Hlineny:2008:WPB

[HiOSG08] Petr Hlineny, Sang il Oum, Detlef Seese, and Georg Gottlob. Width parameters beyond tree-width and their applications. The Computer Journal, 51(3):326–362, May 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/3/326;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/3/326;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/3/326.

Hirvensalo:2004:BRQ

[Hir04] Mika Hirvensalo. Book review: Quantum Computation: A Grand Mathematical Challenge for the Twenty-First Century and the Millennium, by Samuel J. Lomonaco, Jr. (ed.). The Computer Journal, 47(2): 270, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470270.pdf;   http://
www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
02/470270.sgm.abs.html.

Harrison:2008:DCC

[HKG08] Peter G. Harrison, Taskin Kocak, and Erol Gelenbe. Discussant contributions for the Computer Journal Lecture by Erol Gelenbe. The Computer Journal, 51(6):731–734, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/51/
6/731;   http://comjnl.oxfordjournals.
org/cgi/reprint/51/6/731.

Hunt:2003:GJE

[HL03] John Hunt and Chris Loftus. Guide to J2EE: Enterprise Java. Springer professional computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 1-85233-704-4. xxv + 672 pp. LCCN QA76.73.J38 H855 2003.

Huang:2007:EPS

[HL07] Chung-Ming Huang and Jian-Wei Li. Efficient and provably secure IP multimedia subsystem authentication for UMTS. The Computer Journal, 50(6):739–757, November 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/6/739;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/6/739;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/6/739.

Ho:2008:NCI

[HL08] Tracey Ho and Desmond S.

(Desmond Sui) Lun. Network coding: an introduction. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-87310-X (hardcover). xi + 170 pp. LCCN TK5105 .H585

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0810/2008007583-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0810/2008007583-
d.html;   http://www.loc.

                                        gov/catdir/enhancements/
                                        fy0810/2008007583-t.html.

Huang:2009:PVD

[HLL09] Chung-Ming Huang, Chung-Wei Lin, and Xin-Ying Lin. A predictive video-on-demand bandwidth management using the Kalman filter over heterogeneous networks. The Computer Journal, 52(2):171–185, March 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/2/171;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/52/2/171;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/52/2/171.

Heun:2003:EEH

[HM03] Volker Heun and Ernst W. Mayr. Efficient embeddings into hypercube-like topologies. The Computer Journal, 46(6):632–644, November 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_06/460632.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_06/pdf/
                                        460632.pdf.

Hoare:2005:GCC

[HM05] Tony Hoare and Robin Milner. Grand challenges for computing research. The Computer Journal, 48(1):49– 52, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_01/bxh065.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_01/pdf/
bxh065.pdf.

Harrin:2006:BRR

[HM06] Elizabeth Harrin and MBCS.

Book review: Robert S Coble and Shel Israel, Naked Conversations: How Blogs are Changing the Way Businesses Talk with Customers, John Wiley & Sons (2006). ISBN 0-471-74719-X. £16.99. 251 pp. Hardcover. The Computer Journal, 49(6):764, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
6/764;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/6/764.

Huszerl:2002:QAU

[HMP^{+}02] Gábor Huszerl, István Majzik, András Pataricza, Konstantinos Kosmidis, and Mario Dal Cin. Quantitative analysis of UML statechart models of dependable systems. The Computer Journal, 45(3):260– 277, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.

uk/computer_journal/hdb/
Volume_45/Issue_03/450260.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450260.pdf.

Huffner:2008:TPF

[HNW08] Falk Hüffner, Rolf Niedermeier, and Sebastian Wernicke. Techniques for practical fixed-parameter algorithms. The Computer Journal, 51(1):7–25, January 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/1/7;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/1/7;   http://comjnl.
oxfordjournals.org/cgi/reprint/
51/1/7.

Hanaoka:2002:HNI

[HNZI02] Goichiro Hanaoka, Tsuyoshi Nishioka, Yuliang Zheng, and Hideki Imai. A hierarchical non-interactive key-sharing scheme with low memory size and high resistance against collusion attacks. The Computer Journal, 45(3):293–303, ???? 2002. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450293.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450293.pdf.

Hoare:2007:IPC

[Hoa07] Tony Hoare. The ideal of program correctness: Third Computer Journal Lecture. The Computer Journal, 50(3): 254–260, May 2007. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/cgi/content/
full/50/3/254;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/3/254. See

[Ano07b, Ano07c].

Holt:2005:BRB

[Hol05] Alison Holt. Book review:

MOF, A Pocket Guide-IT Service Operations Management, by D. Pultorak, P. Quagliarello, R. Akker and J. van Bon. The Computer Journal, 48(1): 126, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_01/bxh072.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_01/pdf/
bxh072.pdf.

Hassaballah:2008:RSM

[HOM08] M. Hassaballah, Saleh Omran, and Youssef B. Mahdy. A review of SIMD multimedia extensions and their usage in scientific and engineering applications. The Computer Journal, 51(6):630–649, November 2008. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/6/630;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/6/630;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/6/630.

Horng:2001:ESP

[Hor01] Gwoboa Horng. An efficient and secure protocol for multiparty key establishment. The Computer Journal, 44(5):463– 470, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_05/440463.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_05/pdf/
440463.pdf.

Holyer:2000:RMS

[HP00] Ian Holyer and Hüseyin Pehlivan. A recovery mechanism for shells. The Computer Journal, 43(3):168–176, ???? 2000. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_03/430168.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_03/pdf/
430168.pdf.

Hsung-Pin:2004:CAR

[HPCC^{+}04] Hsung-Pin, Chang, Ray-I Chang, Wei-Kuan Shih, and Ruei-Chuan Chang. Cacheaware real-time disk scheduling. The Computer Journal, 47(5):560–578, September 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_05/470560.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_05/pdf/
470560.pdf.

Hromkovic:2003:BR

[Hro03] Juraj Hromkovic. Book review.

The Computer Journal, 46(3): 334–335, May 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_03/460334a.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_03/pdf/
460334a.pdf.

Hromkovic:2005:DAR

[Hro05] Juraj Hromkovič. Design and analysis of randomized algorithms: introduction to design paradigms. Texts in theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-23949-9 (hardcover). xii + 274 pp. LCCN QA274 .H76 2005.

Hwang:2007:PEA

[HS07] Ren-Junn Hwang and Sheng-Hua Shiau. Provably efficient authenticated key agreement protocol for multi-servers. The Computer Journal, 50 (5):602–615, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/5/602;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/5/602;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/5/602.

He:2009:HGA

[HSGS09] Fei He, Xiaoyu Song, Ming Gu, and Jiaguang Sun. Heuristicguided abstraction refinement. The Computer Journal, 52(3):280–287, May 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/3/280;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/52/3/280;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/52/3/280.

Hanaoka:2006:USA

[HSHI06] Goichiro Hanaoka, Junji Shikata, Yumiko Hanaoka, and Hideki Imai. Unconditionally secure anonymous encryption and group authentication. The Computer Journal, 49(3):

310–321, May 2006. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/cgi/content/
abstract/49/3/310;   http:
//comjnl.oxfordjournals.
org/cgi/content/full/49/
3/310;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/3/310.

Huang:2009:JTC

[HSM09] Tzu-Chi Huang, Ce-Kuen Shieh, and Yu-Ben Miao. Justin-time cache for capsule-based active networks. The Computer Journal, 52(6):671–689, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/6/671;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/6/671.

Huang:2006:DIV

[HTT06] Chung-Ming Huang, Ching-Hsien Tsai, and Ming-Chi Tsai. Design and implementation of video streaming hot-plug between wired and wireless networks using SCTP. The Computer Journal, 49(4):400–417, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/4/400;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/4/400;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/4/400.

Hierons:2008:EDT

[HU08] R. M. Hierons and H. Ural. The effect of the distributed test architecture on the power of testing. The Computer Journal, 51(4):497–510, July 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        51/4/497;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/51/4/497;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/51/4/497.

Huang:2006:DTL

[Hua06]  Yi-Sheng Huang.  Design

of traffic light control systems using statecharts. The Computer Journal, 49(6): 634–649, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        49/6/634;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/49/6/634;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/49/6/634.

Hulse:2005:RBC

[Hul05] Paul Hulse. Review: Beowulf Cluster Computing with Linux, second edition. The Computer Journal, 48(3):379–380, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/reprint/48/3/379.

Hunt:2003:BRa

[Hun03a] James J. Hunt. Book review.

The Computer Journal, 46(1): 115, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460115b.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460115b.pdf.

Hunt:2003:BRb

[Hun03b] John Hunt. Book review.

The Computer Journal, 46(4): 462–463, July 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_04/460462a.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_04/pdf/
460462a.pdf.

Ipate:2003:MSX

[Ipa03] Florentin Ipate. On the minimality of stream X-machines. The Computer Journal, 46(3): 295–306, May 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_03/460295.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/

                                        hdb/Volume_46/Issue_03/pdf/
                                        460295.pdf.

Ipate:2005:MFA

[Ipa05]  Florentin Ipate.  On the

minimality of finite automata and stream X-machines for finite languages. The Computer Journal, 48(2):157–167, March 2005. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_48/Issue_02/bxh068.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_48/Issue_02/pdf/
                                        bxh068.pdf.

Ipate:2009:TSH

[Ipa09] Florentin Ipate. Test selection for hierarchical and communicating finite state machines. The Computer Journal, 52(3):334–347, May 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/3/334;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/52/3/334;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/52/3/334.

Iqbal:2008:EEZ

[Iqb08] Azlan Iqbal. Evaluation of economy in a zero-sum perfect information game. The Computer Journal, 51(4):408–418, July 2008. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/4/408;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/4/408;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/4/408.

Iribarne:2004:TSC

[ITV04] Luis Iribarne, José M. Troya, and Antonio Vallecillo. A trading service for COTS components. The Computer Journal, 47(3):342–357, May 2004. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Ibarguengoytia:2006:PMI

[IVS06] Pablo H. Ibargüengoytia, Sunil Vadera, and L. Enrique Sucar. A probabilistic model for information and sensor validation. The Computer Journal, 49(1):113–126, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/1/113;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/1/113;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/1/113.

Iwata:2002:BR

[Iwa02] Hiroo Iwata. Book review.

The Computer Journal, 45(3): 373, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450373.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450364.pdf.

Ibrahim:2009:SOS

[IZ09] Abdelgadir Ibrahim and Liping Zhao. Supporting the OSGi service platform with mobility and service distribution in ubiquitous home environments. The Computer Journal, 52(2):210–239, March 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/2/210;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/2/210;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/2/210.

Jasinski:2006:BRI

[Jas06] Ricardo Jasinski. Book review: S. Ivarama and P. D. Andamudi, Guide to RISC Processors for Programmers and Engineers, Springer-Verlag (2005). ISBN 0-387-21017-2. £36.00, US$46.95, US$59.95. 387 pp. Hardbound. The Computer Journal, 49(3):375, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/

3/375;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/3/375.

Jasinski:2007:FTT

[Jas07]  Ricardo Jasinski.  Fault-

tolerance techniques for SRAMbased FPGAs. The Computer Journal, 50(2):248, March 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
2/248;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/2/248.

Jiao:2005:HSP

[JCL05] Li Jiao, To-Yat Cheung, and Weiming Lu. Handling synchronization problem in Petri net-based system design by property-preserving transitionreduction. The Computer Journal, 48(6):692–701, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/6/692;   http://comjnl.
oxfordjournals.org/cgi/content/
full/48/6/692;   http://
comjnl.oxfordjournals.org/
cgi/reprint/48/6/692.

Jie:2002:PPO

[JCT02] Wei Jie, Wentong Cai, and Stephen J. Turner. POEMS: a parallel object-oriented environment for multi-computer systems. The Computer Journal, 45(5):540–560, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_
05/450540.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_
05/pdf/450540.pdf.

Joo:2007:AMM

[JL07] Kil Hong Joo and Won Suk Lee. An adaptive materialization method for globally optimized query evaluation in a hierarchical mediator system. The Computer Journal, 50(1):68–80, January 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/1/68;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/1/68;   http://comjnl.
oxfordjournals.org/cgi/reprint/
50/1/68.

Jia:2003:PWS

[JLH^{+}03] Xiaohua Jia, Deying Li, Xiaodong Hu, Weili Wu, and Dingzhu Du. Placement of Web-server proxies with consideration of read and update operations on the Internet. The Computer Journal, 46(4): 378–390, July 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_04/460378.
sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/
hdb/Volume_46/Issue_04/pdf/
460378.pdf.

Jia:2001:OPW

[JLHD01] Xiaohua Jia, Deying Li, Xiaodong Hu, and DingZhu Du. Optimal placement of Web proxies for replicated Web servers in the Internet. The Computer Journal, 44(5):329– 339, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_05/440329.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_05/pdf/
440329.pdf.

Juang:2002:VMA

[JLL02] Wen-Shenq Juang, Chin-Laung Lei, and Horng-Twu Liaw. A verifiable multi-authority secret election allowing abstention from voting. The Computer Journal, 45(6):672– 682, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_06/450672.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_06/pdf/
450672.pdf.

Jorgensen:2008:WAU

[JM08] Murray A. Jorgensen and Geoffrey J. McLachlan. Wallace’s approach to unsupervised learning: The snob program. The Computer Journal, 51(5):571–578, September 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/5/571;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/5/571;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/5/571.

Jun:2002:EPT

[JMT02] Yang Jun, Greg Michaelson, and Phil Trinder. Explaining polymorphic types. The Computer Journal, 45(4):436– 452, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_04/450436.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_04/pdf/
450436.pdf.

John:2000:DPS

[Joh00] Lizy Kurian John. Data placement schemes to reduce conflicts in interleaved memories. The Computer Journal, 43(2): 138–151, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_02/430138.
sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/
hdb/Volume_43/Issue_02/pdf/
430138.pdf.

Jorge:2007:TWB

[Jor07] Martinez-Gil Jorge. Thinking on the Web: Berners-Lee, Gödel and Turing. The Computer Journal, 50(3):371–372, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
3/371-a;   http://comjnl.
oxfordjournals.org/cgi/reprint/
50/3/371-a.

Jouannaud:2004:TTR

[Jou04] Jean-Pierre Jouannaud. Terese:

Term rewriting systems. The Computer Journal, 47(1):135– ??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470135.pdf;   http://
www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
01/470135.sgm.abs.html.

Juarna:2007:SGS

[JV07] Asep Juarna and Vincent Vajnovszki. Some generalizations of a Simion–Schmidt bijection. The Computer Journal, 50(5):574–580, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/

                                        50/5/574;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/5/574;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/5/574.

Ji:2009:AWC

[JWLQ09] Meng-Luo Ji, Ji Wang, Shuhao Li, and Zhi-Chang Qi. Automated worst-case execution time analysis based on program modes. The Computer Journal, 52(5):530–544, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/5/530;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        52/5/530.

Kamareddine:2004:CRa

[Kam04a] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 47(1):1–3, January 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/free_
                                        pdf/470001.pdf;   http://
                                        www3.oup.co.uk/computer_
                                        journal/hdb/Volume_47/Issue_
                                        01/470001.sgm.abs.html.

Kamareddine:2004:CRb

[Kam04b] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 47(2):137–139, March 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470137.pdf;   http://
www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
02/470137.sgm.abs.html.

Kamareddine:2005:CRa

[Kam05a] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 48(3):257–258, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/3/257;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/3/257.

Kamareddine:2005:CRb

[Kam05b] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 48(4):382–384, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/4/382;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/4/382.

Kamareddine:2005:CRc

[Kam05c] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 48(5):501, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/5/501;   http://comjnl.

                                        oxfordjournals.org/cgi/reprint/
                                        48/5/501.

Kamareddine:2005:CRd

[Kam05d] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 48(6):627–629, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/6/627;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/48/6/627;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/48/6/627.

Kamareddine:2006:CRa

[Kam06a] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 49(1):1–3, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        49/1/1;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/49/1/1;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        49/1/1.

Kamareddine:2006:CRb

[Kam06b] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 49(3):255–257, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        49/3/255;   http://comjnl.

oxfordjournals.org/cgi/content/
full/49/3/255;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/3/255.

Kamareddine:2006:CRc

[Kam06c] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 49(4):381–382, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/4/381;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/4/381;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/4/381.

Kamareddine:2006:CRd

[Kam06d] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 49(5):505–506, September 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/5/505;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/5/505;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/5/505.

Kamareddine:2006:CRe

[Kam06e] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 49(6):631–633, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.

                                        org/cgi/content/abstract/
                                        49/6/631;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/49/6/631;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/49/6/631.

Kamareddine:2007:CRa

[Kam07a] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 50(1):1–3, January 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/1/1;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/1/1;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        50/1/1.

Kamareddine:2007:CRb

[Kam07b] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 50(2):131–133, March 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/2/131;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/2/131;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/2/131.

Kamareddine:2007:CRc

[Kam07c] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 50(3):251–253, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/3/251;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/3/251;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/3/251.

Kamareddine:2007:CRd

[Kam07d] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 50(4):373–374, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/4/373;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/4/373;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/4/373.

Kamareddine:2007:CRe

[Kam07e] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 50(5):501–504, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/5/501;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/5/501;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/5/501.

Kamareddine:2007:CRf

[Kam07f] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 50(6):629, November 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/6/629;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/6/629;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/6/629.

Kamareddine:2008:CRa

[Kam08a] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 51(2):141–143, March 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        51/2/141;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/51/2/141;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/51/2/141.

Kamareddine:2008:CRb

[Kam08b] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 51(4):405–407, July 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        51/4/405;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/51/4/405;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/51/4/405.

Kamareddine:2008:CRc

[Kam08c] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 51(6):609–611, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/6/609;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/6/609;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/6/609.

Kamareddine:2009:CRa

[Kam09a] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 52(2):169–170, March 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/2/169;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/2/169;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/2/169.

Kamareddine:2009:CRb

[Kam09b] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 52(3):277–279, May 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/3/277;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/3/277;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/3/277.

Kamareddine:2009:CRc

[Kam09c] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 52(4):393–394, July 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/4/393;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        52/4/393.

Kang:2004:SMD

[KBX^{+}04] Porlin Kang, Cristian Borcea, Gang Xu, Akhilesh Saxena, Ulrich Kremer, and Liviu Iftode. Smart messages: a distributed computing platform for networks of embedded systems. The Computer Journal, 47(4): 475–494, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_04/470475.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_04/pdf/
                                        470475.pdf.

Kelleher:2004:BRB

[Kel04] B. Kelleher. Book review: Shaping the IT Organization: The Impact of Outsourcing and the New Business Model, by I. Gouge. The Computer Journal, 47(6):728–729, November 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.

uk/computer_journal/free_
pdf/470728a.pdf;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
06/470728a.sgm.abs.html.

Kerren:2004:BRB

[Ker04] Andreas Kerren. Book review:

Software Visualization — From Theory to Practice, by Kang Zhang. The Computer Journal, 47(3):395, May 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470395.pdf;   http://
www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
03/470395.sgm.abs.html.

Kalva:2005:CEH

[KF05] Hari Kalva and Borko Furht.

Complexity estimation of the H.264 coded video bitstreams. The Computer Journal, 48 (5):504–513, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/5/504;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/5/504.

Kuo:2003:IWT

[KHL03] Tei-Wei Kuo, Yen-Hsi Hou, and Kam-Yiu Lam. The impacts of write-through procedures and checkpointing on realtime concurrency control. The Computer Journal, 46(2):174– 192, February 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_02/460174.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_02/pdf/
460174.pdf.

Kirchner:2002:BR

[Kir02] Claude Kirchner. Book review. The Computer Journal, 45(2): 255–256, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_02/450255.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_02/pdf/
450255.pdf.

Kirda:2006:PUA

[KK06] Engin Kirda and Christopher Kruegel. Protecting users against phishing attacks. The Computer Journal, 49(5):554–561, September 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/5/554;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/5/554;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/5/554.

Kapus-Kolar:2007:TCE

[KK07] Monika Kapus-Kolar. Testing as collecting of evidence: An integrated approach to test generation for finite state machines. The Computer Journal, 50(3):315–331, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/3/315;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/3/315;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/3/315.

Kapus-Kolar:2009:TGG

[KK09] Monika Kapus-Kolar. Three generalizations to a generic integrated test generation method for finite state machines. The Computer Journal, 52(6):599–625, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/6/599;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/6/599.

Kaporis:2001:LIU

[KKK^{+}01] Alexis C. Kaporis, Lefteris M. Kirousis, Evangelos Kranakis, Danny Krizanc, Yannis C. Stamatiou, and Elias C. Stavropoulos. Locating information with uncertainty in fully interconnected networks with applications to World Wide Web information retrieval. The Computer Journal, 44(4):221– 229, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_04/440221.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_04/pdf/
440221.pdf.

Kim:2002:PAT

[KL02a] Daeyoung Kim and Yann-Hang Lee. Periodic and aperiodic task scheduling in strongly partitioned integrated realtime systems. The Computer Journal, 45(4):395–409, ???? 2002. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_04/450395.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_04/pdf/
450395.pdf.

Korsh:2002:LGT

[KL02b] James F. Korsh and Paul LaFollette. Loopless generation of trees with specified degrees. The Computer Journal, 45(3): 364–372, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450364.

sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450364.pdf.

Korsh:2003:LGS

[KL03] James F. Korsh and Paul S.

LaFollette. Loopless generation of Schröder trees. The Computer Journal, 46(1):106– 113, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460106.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460106.pdf.

Korsh:2004:LAG

[KL04] James F. Korsh and Paul S.

LaFollette. Loopless array generation of multiset permutations. The Computer Journal, 47(5):612–621, September 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_05/470612.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_05/pdf/
470612.pdf.

Klonowska:2004:COP

[KLLB04] Kamilla Klonowska, Lars Lundberg, Håkan Lennerstad, and Magnus Broberg. Comparing the optimal performance of parallel architectures. The Computer Journal, 47(5):527– 544, September 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_05/470527.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_05/pdf/
                                        470527.pdf.

Kratochvil:2002:UXL

[KM02] Milan Kratochvíl and Barry McGibbon. UML xtra-light: how to specify your software requirements. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-89242-2. LCCN QA76.76.A65 K72 2002b; **See. URL http:

                                        //library.books24x7.com/
                                        library.asp?B&isbn=0521892422.

Katsaggelos:2009:GE

[KM09] Aggelos K. Katsaggelos and Rafael Molina. Guest editorial. The Computer Journal, 52(4):395–396, July 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/reprint/52/4/395.

Kihlstrom:2003:BFD

[KMMS03] Kim Potter Kihlstrom, Louise E. Moser, and P. M. Melliar-Smith. Byzantine fault detectors for solving consensus. The Computer Journal, 46(1):16– 35, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460016.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460016.pdf.

Kocak:2006:HNP

[Koc06]  Taskin Kocak.  A hybrid

network processor with support for high-speed execution of CPN and upper layer IP protocols. The Computer Journal, 49(2):190–200, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/2/190;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/2/190;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/2/190.

Kone:2001:LAT

[Kon01]  Ousmane Koné.  A lo-

cal approach to the testing of real-time systems. The Computer Journal, 44(5):435– 447, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_05/440435.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_05/pdf/
440435.pdf.

Koriem:2000:FPN

[Kor00] Samir M. Koriem. A fuzzy Petri net tool for modeling and verification of knowledgebased systems. The Computer Journal, 43(3):206–223, ???? 2000. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_43/Issue_03/430206.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_43/Issue_03/pdf/
                                        430206.pdf.

Korsh:2005:GAT

[Kor05] James F. Korsh. Generating T -ary trees in linked representation. The Computer Journal, 48(4):488–497, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/4/488;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        48/4/488.

Koriem:2006:DAE

[Kor06a] Samir M. Koriem. Development, analysis and evaluation of performance models for mobile multi-agent networks. The Computer Journal, 49(6):685–709, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/

49/6/685;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/6/685;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/6/685.

Korsh:2006:SAA

[Kor06b]  James F. Korsh.  Skar-

bek’s algorithm for t-ary trees. The Computer Journal, 49(3):351–357, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/3/351;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/3/351;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/3/351.

Kovac:2003:CAG

[Kov03] Caroline Kovac. Computing in the age of the genome. The Computer Journal, 46(6):593– 597, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_06/460593.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_06/pdf/
460593.pdf.

Kozatchok:2002:UTM

[KP02] Irina Kozatchok and Samuel Pierre. User tracking and mobility management algorithm for wireless networks. The Computer Journal, 45(5):525– 539, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_05/450525.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_05/pdf/
450525.pdf.

Koloniari:2004:FXB

[KP04] Georgia Koloniari and Evaggelia Pitoura. Filters for XMLbased service discovery in pervasive computing. The Computer Journal, 47(4):461– 474, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_04/470461.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_04/pdf/
470461.pdf.

Kamareddine:2002:PTS

[KR02] Fairouz Kamareddine and Alejandro Ríos. Pure type systems with de Bruijn indices. The Computer Journal, 45(2): 187–201, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_02/450187.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_02/pdf/
450187.pdf.

Kurki-Suonio:2005:PTR

[KS05] Reino Kurki-Suonio. A practical theory of reactive systems: incremental modeling of dynamic behaviors. Texts in theoretical computer science: an EATCS series. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-23342-3 (hardcover). xxi + 418 pp. LCCN QA76.76.D47 K87

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0663/2005920137-d.html;
http://www.loc.gov/catdir/
enhancements/fy0818/2005920137-
b.html;   http://www.loc.
gov/catdir/enhancements/
fy0818/2005920137-t.html.

Kuznetsov:2005:SI

[Kuz05] Nikolaj A. Kuznetsov. The science of infocommunications. The Computer Journal, 48(4):460–465, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/4/460;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/4/460.

Klein:2003:PHD

[KW03] S. T. Klein and Y. Wiseman.

Parallel Huffman decoding with applications to JPEG files. The Computer Journal, 46(5):487– 497, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_05/460487.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_05/pdf/

460487.pdf. This is an extended version of a paper that appeared earlier in Klein, S. T. and Wiseman, Y. (2000) Proc. Data Compression Conf. DCC’00, Snowbird, Utah, pp. 383–392. IEEE Computer Society Press, Los Alamitos, CA.

Kuo:2008:ACC

[KW08] Yu-Chen Kuo and Po-Yao Wu. The availability of complemental k-coteries. The Computer Journal, 51(6): 677–687, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/6/677;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/6/677;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/6/677.

Lamb:2004:BRB

[Lam04] Luís C. Lamb. Book review: Logics for Emerging Applications of Databases, edited by J. Chomicki and R. van der Meyden and G. Saake. The Computer Journal, 47(4):505– 506, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.

uk/computer_journal/free_
pdf/470505a.pdf;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
04/470505a.sgm.abs.html.

Land:2003:BR

[Lan03] Frank Land. Book review.

The Computer Journal, 46(5): 526, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_05/460526.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_05/pdf/
460526.pdf.

Lano:2005:ASD

[Lan05] K. Lano. Advanced systems design with Java, UML, and MDA. Elsevier Butterworth-Heinemann, Amsterdam, The Netherlands, 2005. ISBN 0-7506-6496-7. viii + 378 pp. LCCN QA76.76.D47 L37 2005; QA76.76.D47 L375 2005.

Land:2006:LIM

[Lan06]  Frank Land.  Leo II and

the Model T Ford. The Computer Journal, 49(6): 650–656, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/6/650;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/6/650;   http://

                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/49/6/650.

Lavington:2009:ADS

[Lav09] Simon Lavington. An appreciation of Dina St Johnston (1930– 2007) founder of the UK’s first software house. The Computer Journal, 52(3):378–387, May 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/3/378;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/52/3/378;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/52/3/378.

Losada:2001:LMI

[LB01] David E. Losada and Alvaro Barreiro. A logical model for information retrieval based on propositional logic and belief revision. The Computer Journal, 44(5):410– 424, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_44/Issue_05/440410.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_05/pdf/
                                        440410.pdf.

Levene:2007:CTO

[LBI07] Mark Levene and Judit Bar-Ilan. Comparing typical opening move choices made by humans and chess engines.

The Computer Journal, 50 (5):567–573, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/5/567;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/5/567;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/5/567.

Laufer:2000:SSC

[LBR00] Konstantin Läufer, Gerald Baumgartner, and Vincent F. Russo. Safe structural conformance for Java. The Computer Journal, 43(6):469– 481, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_06/430469.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_06/pdf/
430469.pdf.

Li:2007:CDW

[LC07] Der-Chiang Li and Fengming M. Chang. An in–out combined dynamic weighted round-robin method for network load balancing. The Computer Journal, 50(5): 555–566, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/

                                        50/5/555;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/5/555;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/5/555.

Lee:2004:RBS

[LCBK04] Byung Suk Lee, Li Chen, Jeff Buzas, and Vinod Kannoth. Regression-based selftuning modeling of smooth user-defined function costs for an object-relational database management system query optimizer. The Computer Journal, 47(6):673–693, November 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_06/470673.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_06/pdf/
                                        470673.pdf.

Lacks:2009:DED

[LCK09] Daniel Lacks, Mainak Chatterjee, and Taskin Kocak. Design and evaluation of a distributed clustering algorithm for mobile ad hoc networks. The Computer Journal, 52(6):656–670, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/6/656;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        52/6/656.

Lau:2001:AMP

[LCT01] F. C. M. Lau, P. K. W. Cheng, and S. S. H. Tse. An algorithm for the 2-median problem on two-dimensional meshes. The Computer Journal, 44(2):101– 108, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_02/440101.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_02/pdf/
440101.pdf.

Luck:2001:CFA

[Ld01] Michael Luck and Mark d’Inverno.

A conceptual framework for agent definition and development. The Computer Journal, 44(1):1–20, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_01/440001.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_01/pdf/
440001.pdf.

Lent:2006:DMT

[Len06]  Ricardo Lent.  Design of

a MANET testbed management system. The Computer Journal, 49(2):171–179, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/

                                        49/2/171;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/49/2/171;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/49/2/171.

Leonardi:2003:BR

[Leo03] Emilio Leonardi. Book review. The Computer Journal, 46(3):334, May 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_03/460334.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_03/pdf/
                                        460334.pdf.

Levin:2005:ATB

[Lev05]  Leonid A. Levin.  Aperi-

odic tilings: Breaking translational symmetry. The Computer Journal, 48(6): 642–645, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/6/642;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/48/6/642;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/content/full/bxh124/
                                        DC1;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/48/6/642.

Levene:2006:ISE

[Lev06a] M. (Mark) Levene. An introduction to search engines and Web navigation. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-30677-5 (paperback). xxvi + 365 pp. LCCN TK5105.884 .L48 2006.

Levene:2006:BRS

[Lev06b] Mark Levene. Book review:

A. Spink and B. J. Jansen, Web Search: Public Searching of the Web. Information Science & Knowledge Management, Vol. 6. Kluwer Academic Publishers (2004). ISBN 1-4020-2268-9. 100. 199 pp. Hardcover. The Computer Journal, 49(1):128–129, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
1/128;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/1/128.

Levene:2007:AIG

[Lev07] Mark Levene. Artificial intelligence for games. series in interactive 3D technology. The Computer Journal, 50(3):371, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
3/371;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/3/371.

Lano:2000:SCF

[LFB00] K. Lano, R. France, and J.- M. Bruel. A semantic comparison of fusion and syntropy. The Computer Journal, 43(6): 451–468, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_06/430451.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_06/pdf/
430451.pdf.

Lema:2003:AMO

[LFG^{+}03] José Antonio Cotelo Lema, Luca Forlizzi, Ralf Hartmut Güting, Enrico Nardelli, and Markus Schneider. Algorithms for moving objects databases. The Computer Journal, 46(6):680–712, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_06/460680.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_06/pdf/
460680.pdf.

Lee:2001:UPC

[LH01] SangKeun Lee and Chong-Sun Hwang. Unified protocols of concurrency control and recovery in distributed object-based databases. The Computer Journal, 44(3):201– 213, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_03/440201.
sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/
hdb/Volume_44/Issue_03/pdf/
440201.pdf.

Lhoussaine:2004:BRB

[Lho04] Cdric Lhoussaine. Book review: The Pi-Calculus — A Theory of Mobile Processes, by D. Sangiorgi and D. Walker. The Computer Journal, 47(3): 396, May 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470396.pdf;   http://
www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
03/470396.sgm.abs.html.

Li:2003:SPP

[Li03]  Keqin Li.  Speed-up of

parallel processing of divisible loads on k-dimensional meshes and tori. The Computer Journal, 46(6):625–631, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_06/460625.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_06/pdf/
460625.pdf.

Li:2007:RCI

[Li07] Wei Li. R-calculus: An inference system for belief revision. The Computer Journal, 50(4):378–390, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/4/378;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/4/378;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/4/378.

Li:2008:ACP

[Li08] Keqin Li. Average-case performance analysis of online non-clairvoyant scheduling of parallel tasks with precedence constraints. The Computer Journal, 51(2):216–226, March 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/2/216;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/2/216;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/2/216.

Limb:2002:BR

[Lim02] Rowan Limb. Book review. The Computer Journal, 45(3): 373–374, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450373b.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450364.pdf.

Lioy:2003:BRa

[Lio03a] Antonio Lioy. Book review. The Computer Journal, 46(2):223– 224, February 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_02/460223a.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_02/pdf/
460223a.pdf.

Lioy:2003:BRb

[Lio03b] Antonio Lioy. Book review.

The Computer Journal, 46(5): 524, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_05/460524.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_05/pdf/
460524.pdf.

Lai:2001:PED

[LIPP01] C.-H. Lai, C. S. Ierotheou, C. J. Palansuriya, and K. A. Pericleous. Performance evaluation of a distributed algorithm for an inverse heat conduction problem. The Computer Journal, 44(3):214– 220, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_03/440214.
sgm.abs.html;   http://www3.

                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_03/pdf/
                                        440214.pdf.

Lopez:2009:IWS

[LKCK09] Tomás Sánchez López, Daeyoung Kim, Gonzalo Huerta Canepa, and Koudjo Koumadi. Integrating wireless sensors and RFID tags into energyefficient and dynamic context networks. The Computer Journal, 52(2):240–267, March 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/2/240;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/52/2/240;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/52/2/240.

Liu:2007:SOA

[LL07] Tzong-Jye Liu and Chia-Lin Lee. State-optimal alternator for uniform synchronous rings. The Computer Journal, 50(3):341–347, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/3/341;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/3/341;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/3/341.

Liu:2000:CVD

[LLH00] Duen-Ren Liu, Chen-Hsien Lin, and Jing-Jang Hwang. Classifying video documents by hierarchical structure of video contents. The Computer Journal, 43(5):396–410, ???? 2000. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_05/430396.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_05/pdf/
430396.pdf.

Lee:2007:HSL

[LLL07] Jaesung Lee, Hyuk-Jae Lee, and Chanho Lee. A high-speed link layer architecture for low latency and memory cost reduction. The Computer Journal, 50(5):616–628, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/5/616;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/5/616;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/5/616.

Lee:2009:PBA

[LLL09] Jaesung Lee, Hyuk-Jae Lee, and Chanho Lee. A phasebased approach for on-chip bus architecture optimization. The Computer Journal, 52(6):626–645, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/6/626;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        52/6/626.

Lee:2002:CCU

[LLS02] Victor C. S. Lee, Kwok-Wa Lam, and Sang H. Son. Concurrency control using timestamp ordering in broadcast environments. The Computer Journal, 45(4):410–422, ???? 2002. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_04/450410.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_04/pdf/
                                        450410.pdf.

Lam:2005:UHS

[LLZ05] Kam-Yiu Lam, BiYu Liang, and ChuanLin Zhang. On using handoff statistics and velocity for location management in cellular wireless networks. The Computer Journal, 48(1):84– 100, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_48/Issue_01/bxh063.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_48/Issue_01/pdf/
                                        bxh063.pdf.

Luccio:2000:MIR

[LMOP00] F. Luccio, M. Mahafzah, M. Omari, and L. Pagli. Masked interval routing: a new routing scheme. The Computer Journal, 43(2):121– 129, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_02/430121.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_02/pdf/
430121.pdf.

Lamma:2004:SMF

[LMR04] Evelina Lamma, Paola Mello, and Fabrizio Riguzzi. A system for measuring function points from an ER-DFD specification. The Computer Journal, 47(3): 358–372, May 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Li:2007:QSR

[LN07] Sanjiang Li and Bernhard Nebel. Qualitative spatial representation and reasoning: a hierarchical approach. The Computer Journal, 50(4):391–402, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/4/391;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/4/391;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/4/391.

Lieberherr:2003:ACC

[LO03] David H. Lorenz Karl Lieberherr and Johan Ovlinger. Aspectual collaborations: Combining modules and aspects. The Computer Journal, 46 (5):542–565, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_46/Issue_
                                        05/460542.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_46/Issue_
                                        05/pdf/460542.pdf.

Locatelli:2002:BR

[Loc02] Marco Locatelli. Book review. The Computer Journal, 45(2): 253, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_02/450253.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_02/pdf/
                                        450249.pdf.

Lomonaco:2002:QCG

[Lom02] Samuel J. Lomonaco, Jr., editor. Quantum computation: a grand mathematical challenge for the Twenty-First Century and the Millennium: American Mathematical Society, Short Course, January 17–18, 2000, Washington, DC, volume 58 of Proceedings of symposia in applied mathematics; AMS short course lecture notes. American Mathematical Society, Providence, RI, USA, 2002. ISBN 0-8218-2084-2. ISSN 0160-7634. xvi + 358 pp. LCCN QA76.889 .A44 2000. URL http://www.

loc.gov/catdir/toc/fy034/
2001055366.html.

Lombardo:2003:BR

[Lom03] Vincenzo Lombardo. Book review. The Computer Journal, 46(1):117–119, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_46/Issue_
01/460117b.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_46/Issue_
01/pdf/460117b.pdf.

Li:2009:OAF

[LPC09] Yamin Li, Shietung Peng, and Wanming Chu. Optimal algorithms for finding a trunk on a tree network and its applications. The Computer Journal, 52(2):268–275, March 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/2/268;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/2/268;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/2/268.

Langston:2008:ICM

[LPS^{+}08] Michael A. Langston, Andy D.

Perkins, Arnold M. Saxton, Jon A. Scharff, and Brynn H. Voy. Innovative computational methods for transcriptomic data analysis: a case study in the use of FPT for practical algorithm design and implementation. The Computer Journal, 51(1):26–38, January 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/1/26;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/1/26;   http://comjnl.
oxfordjournals.org/cgi/reprint/
51/1/26.

Law:2004:PCA

[LS04] K. L. Eddie Law and Sunny So. Pervasive computing on active networks. The Computer Journal, 47(4):418– 431, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_04/470418.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_04/pdf/
470418.pdf.

Li:2006:DEL

[LS06] Guohui Li and LihChyun Shu. Design and evaluation of a lowlatency checkpointing scheme for mobile computing systems. The Computer Journal, 49(5):527–540, September 2006. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/5/527;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/5/527;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/5/527.

Latif-Shabgahi:2004:MEF

[LSBB04] G. Latif-Shabgahi, J. M. Bass, and S. Bennett. Multiple error filtering in cyclic systems. The Computer Journal, 47(3):299– 308, May 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Little:2002:UBF

[LSS02] M. C. Little, S. K. Shrivastava, and N. A. Speirs. Using Bloom filters to speed-up name lookup in distributed systems. The Computer Journal, 45(6): 645–652, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_06/450645.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_06/pdf/
450645.pdf.

Liu:2005:TFD

[LSY05] Wei Yi Liu, Ning Song, and Hong Yao. Temporal functional dependencies and temporal nodes Bayesian networks. The Computer Journal, 48(1): 30–41, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_01/bxh066.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_01/pdf/
bxh066.pdf.

Lin:2004:ELP

[LTL04] Ting-Yu Lin, Yu-Chee Tseng, and Yuan-Ting Lu. An efficient link polling policy by pattern matching for Bluetooth Piconets. The Computer Journal, 47(2):169–178, March 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_02/470169.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_02/pdf/
470169.pdf.

Lucas:2004:UBT

[Luc04] Joan M. Lucas. Untangling binary trees via rotations. The Computer Journal, 47(2):259– 269, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_02/470259.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_02/pdf/
470259.pdf.

Luigi:2003:BR

[Lui03] Liquori Luigi. Book review.

The Computer Journal, 46(6): 661, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_06/460661.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_06/pdf/
460661.pdf.

Luk:2001:CMD

[Luk01] Robert W. P. Luk. Compact morphic directed acyclic word graphs. The Computer Journal, 44(5):425–434, ???? 2001. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_05/440425.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_05/pdf/
440425.pdf.

Lorentzos:2006:RFM

[LV06] Nikos A. Lorentzos and Jose R. R. Viqueira. Relational formalism for the management of spatial data. The Computer Journal, 49(1):62–81, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/1/62;   http://comjnl.
oxfordjournals.org/cgi/content/

                                        full/49/1/62;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        49/1/62.

Liang:2007:UOF

[LWY07] Hongli Liang, Jue Wang, and Yiyu Yao. User-oriented feature selection for machine learning. The Computer Journal, 50(4):421–434, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/4/421;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/4/421;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/4/421.

Laurentowski:2004:EIE

[LZ04a] Aleksander Laurentowski and Krzysztof Zieliński. Experiences from implementation and evaluation of event processing and distribution of notifications in an object monitoring service. The Computer Journal, 47(2): 205–220, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_02/470205.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_02/pdf/
                                        470205.pdf.

Li:2004:CAB

[LZ04b] Hua Li and Chang N. Zhang. A cellular automata based reconfigurable architecture for hybrid cryptosystems. The Computer Journal, 47(3):320–328, May 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Lu:2007:GES

[LZ07] Ruqian Lu and Zhi-Hua Zhou.

Guest editorial to the special issue on advances of artificial intelligence and knowledge engineering in China. The Computer Journal, 50(4):377, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
4/377;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/4/377.

Martin:2006:AH

[MA06] Duncan Martin and Helen Ashman. Atomic hypermedia. The Computer Journal, 49(1):20–31, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/1/20;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/1/20;   http://comjnl.
oxfordjournals.org/cgi/reprint/
49/1/20.

Macii:2003:BR

[Mac03] Enrico Macii. Book review. The Computer Journal, 46(1):115– 116, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460115c.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460115c.pdf.

McGough:2005:MGP

[MAD^{+}05] Andrew Stephen McGough, Ali Afzal, John Darlington, Nathalie Furmento, Anthony Mayer, and Laurie Young. Making the grid predictable through reservations and performance modelling. The Computer Journal, 48(3):358–368, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/3/358;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/3/358.

Mahdoum:2006:BRF

[Mah06] Ali Mahdoum. Book review: Franz Rothlauf, Representations for Genetic and Evolutionary Algorithms. Springer-Verlag (2006, second edition). ISBN 3-540-25059-X. $103.95/$129.00/£77.00. 325 pp. Hardbound. The Computer Journal, 49(5):629, September 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
5/629;   http://comjnl.oxfordjournals.

org/cgi/reprint/49/5/629.

See [Rot06].

Mairs:2007:IED

[Mai07] Chris Mairs. Inclusion and exclusion in the digital world: Turing Lecture 2006. The Computer Journal, 50(3):274–280, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/3/274;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/3/274;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/3/274.

Makinen:2001:CFM

[Mäk01]  Erkki Mäkinen.  Comment

on ‘A Framework for Modelling Trojans and Computer Virus Infection’. The Computer Journal, 44(4):321–323, ???? 2001. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_04/440321.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_04/pdf/

440321.pdf. See [TAC98, TAC01, Coh01].

Manolopoulos:2003:ISS

[Man03] Yannis Manolopoulos. Introduction to the special section on spatiotemporal databases. The Computer Journal, 46 (6):662–663, November 2003.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_46/Issue_
                                        06/460662.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_46/Issue_
                                        06/pdf/460662.pdf.

Manzanares:2007:PPT

[Man07] Antonio Izquierdo Manzanares. Professional pen testing for Web applications — programmer to programmer series (P2P). The Computer Journal, 50(4):500, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/50/
                                        4/500;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/50/4/500.

Martin:2000:TCC

[Mar00] Jeremy M. R. Martin. A tool for checking the CSP sat property. The Computer Journal, 43(1):13–23, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/content/43/1/13.full.
                                        pdf+html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_43/Issue_01/430013.
                                        sgm.abs.html.

Mares:2005:BRB

[Mar05] Peter Mares. Book review: Art of Java Web Development, by N. Ford. The Computer Journal, 48(2):253, March 2005. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/bxh071.pdf;   http://
www3.oup.co.uk/computer_
journal/hdb/Volume_48/Issue_
02/bxh071.sgm.abs.html.

Martins:2008:BEB

[Mar08a] Manuel A. Martins. On the behavioral equivalence between kdata structures. The Computer Journal, 51(2):181–191, March 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/2/181;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/2/181;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/2/181.

Marx:2008:PCA

[Mar08b] Dániel Marx. Parameterized complexity and approximation algorithms. The Computer Journal, 51(1):60–78, January 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/1/60;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/1/60;   http://comjnl.
oxfordjournals.org/cgi/reprint/
51/1/60.

Marecek:2009:BRY

[Mar09] Jakub Marecek. Book review: Yves Pochet and Laurence A. Wolsey, Production Planning Using Mixed Integer Programming, Springer (2006). ISBN-13 978-0-387-29959-4. £38.50/$59.95. 477 pp. Hardbound. The Computer Journal, 52(6):724–725, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/reprint/52/6/724.

McGettrick:2005:GCC

[MBI^{+}05] Andrew McGettrick, Roger Boyle, Roland Ibbett, John Lloyd, Gillian Lovegrove, and Keith Mander. Grand challenges in computing: Education — a summary. The Computer Journal, 48(1):42– 48, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_48/Issue_01/bxh064.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_48/Issue_01/pdf/
                                        bxh064.pdf.

Meier:2005:TDE

[MC05] René Meier and Vinny Cahill. Taxonomy of distributed eventbased programming systems. The Computer Journal, 48 (5):602–626, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/5/602;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/5/602.

McCue:2007:DMP

[McC07] Colleen McCue. Data mining and predictive analysis: intelligence gathering and crime analysis. Butterworth-Heinemann, Amsterdam, The Netherlands, 2007. ISBN 0-7506-7796-1 (paperback). xxxi + 332 pp. LCCN HV7936.C88

M37 2007. URL http://www.
loc.gov/catdir/enhancements/
fy0646/2006040568-d.html.

Mohammad-Djafari:2009:SRS

[MD09] Ali Mohammad-Djafari. Superresolution: a short review, A new method based on hidden Markov modeling of HR image and future challenges. The Computer Journal, 52(1):126–141, January 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/1/126;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/1/126;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/1/126.

Meghanathan:2009:LPB

[Meg09] Natarajan Meghanathan. A location prediction-based reactive routing protocol to minimize the number of route discoveries and hop count per path in mobile ad hoc networks. The Computer Journal, 52(4):461–482, July 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/4/461;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/4/461.

Menzies:2002:BR

[Men02] David Menzies. Book review. The Computer Journal, 45(3): 374–375, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450374a.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450364.pdf.

Meo:2003:BRa

[Meo03a]  Rosa Meo.  Book review.

The Computer Journal, 46(2): 223, February 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_02/460223.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_02/pdf/
460223.pdf.

Meo:2003:BRb

[Meo03b]  Rosa Meo.  Book review.

The Computer Journal, 46(6): 661, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_06/460661a.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_06/pdf/
460661a.pdf.

Ma:2006:PIV

[MH06] Jixin Ma and Pat Hayes. Primitive intervals versus pointbased intervals: Rivals or allies? The Computer Journal, 49(1):32–41, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/1/32;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/1/32;   http://comjnl.
oxfordjournals.org/cgi/reprint/
49/1/32.

Michaelson:2003:BR

[Mic03] Greg Michaelson. Book review.

The Computer Journal, 46(5): 525, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_05/460525.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/

                                        hdb/Volume_46/Issue_05/pdf/
                                        460525.pdf.

Milner:2006:UCS

[Mil06]  Robin Milner.  Ubiquitous

computing: Shall we understand it? The Computer Journal, 49(4):383–389, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/49/
                                        4/383;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/49/4/383.

Mishra:2003:BR

[Mis03] Alok Mishra. Book review. The Computer Journal, 46(6): 660, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_06/460660.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_06/pdf/
                                        460660.pdf.

Mishra:2004:BRBb

[Mis04a] Alok Mishra. Book review: Software Quality Assurance — From Theory to Implementation, by D. Galin. The Computer Journal, 47(6):728, November 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/free_
                                        pdf/470728.pdf;   http://
                                        www3.oup.co.uk/computer_

journal/hdb/Volume_47/Issue_
06/470728.sgm.abs.html.

Mishra:2004:BRBa

[Mis04b]  Alok Mishra.  Book re-

view: UML Xtra-Light — How to Specify Your Software Requirements, by M. Kratochvíl and B. McGibbon. The Computer Journal, 47 (4):507, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470507a.pdf;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
04/470507a.sgm.abs.html.

Ma:2003:RDI

[MK03] Jixin Ma and Brian Knight.

Representing the dividing instant. The Computer Journal, 46(2):213–222, February 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_02/460213.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_02/pdf/
460213.pdf.

Martin-Lof:2006:YZA

[ML06] Per Martin-Löf. 100 years of Zermelo’s Axiom of Choice: what was the problem with it? The Computer Journal, 49(3):345–350, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
3/345;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/3/345.

Manousaka:2000:FAT

[MM00] Antigoni Manousaka and Yannis Manolopoulos. Fringe analysis of 2-3 trees with lazy parent split. The Computer Journal, 43(5):420– 429, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_05/430420.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_05/pdf/
430420.pdf.

Marinescu:2007:BLQ

[MM07] Dan C. Marinescu and Gabriela M. Marinescu. The Boole Lecture: Quantum information: a glimpse at the strange and intriguing future of information. The Computer Journal, 50(5):505–521, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/5/505;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/5/505;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/5/505.

Moller:2002:EFM

[Mol02] Faron G. Moller. Editorial:

Formal methods in computation. The Computer Journal, 45(1):1, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_01/450001.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_01/pdf/
450001.pdf.

Montuschi:2002:BRa

[Mon02a] Paolo Montuschi. Book review. The Computer Journal, 45(4):475, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_04/450475b.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_04/pdf/
450475b.pdf.

Montuschi:2002:BRb

[Mon02b] Paolo Montuschi. Book review. The Computer Journal, 45(5):579, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_05/450579.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_05/pdf/
450579.pdf.

Morrison:2006:BRK

[Mor06] Ronald Morrison. Book review: K. C. Tan, E. F. Khor and T. H. Lee, Multiobjective Evolutionary Algorithms and Applications. Series: Advanced Information & Knowledge Processing. Springer-Verlag (2005). ISBN 1-85233-836-9. $89.95/ $99.95/£50.00. 295 pp. Hardbound. The Computer Journal, 49(2):252–253, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/49/
                                        2/252;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/49/2/252.

Moth:2002:BR

[Mot02] Daniel D. G. Moth. Book review. The Computer Journal, 45(4):473–474, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_45/Issue_
                                        04/450473b.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_45/Issue_
                                        04/pdf/450473b.pdf.

Martin:2006:ACM

[MPXT06] Patrick Martin, Wendy Powley, Xiaoyi Xu, and Wenhu Tian. Automated configuration of multiple buffer pools. The Computer Journal, 49(4):487–499, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/4/487;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/4/487;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/4/487.

Mitianoudis:2007:JFB

[MS07] Nikolaos Mitianoudis and Tania Stathaki. Joint fusion and blind restoration for multiple image scenarios with missing data. The Computer Journal, 50(6):660–673, November 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/6/660;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/6/660;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/6/660.

Murtagh:2000:OCD

[MSB00] Fionn Murtagh, Jean-Luc Starck, and Michael W. Berry. Overcoming the curse of dimensionality in clustering by means of the wavelet transform. The Computer Journal, 43(2): 107–120, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_02/430107.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/

hdb/Volume_43/Issue_02/pdf/
430107.pdf.

Mahmud:2002:FTH

[MSK02] Syed Masud Mahmud, L. Tissa Samaratunga, and Shilpa Kommidi. Fault-tolerant hierarchical networks for shared memory multiprocessors and their bandwidth analysis. The Computer Journal, 45(2):147– 161, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_02/450147.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_02/pdf/
450147.pdf.

Mahmood:2007:AMD

[MTO^{+}07] Asher Mahmood, Philip M. Tudor, William Oxford, Robert Hansford, James D. B. Nelson, Nicholas G. Kingsbury, Antonis Katartzis, M. Petrou, N. Mitianoudis, T. Stathaki, Alin Achim, David Bull, Nishan Canagarajah, Stavri Nikolov, Artur Loza, and Nedeljko Cvejic. Applied multi-dimensional fusion. The Computer Journal, 50(6):646–659, November 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/6/646;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/6/646;   http://

comjnl.oxfordjournals.org/
cgi/reprint/50/6/646.

Mumford:2003:RHS

[Mum03]  Enid Mumford.  Redesign-

ing human systems. Information Science Publishing, Hershey, PA, 2003. ISBN 1-59140-118-6 (hardcover), 1-59140-119-4 (ebook). ix + 292 pp. LCCN HD58.8 .M85 2003.

Munro:2002:BR

[Mun02] Barry Munro. Book review.

The Computer Journal, 45(3): 373, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450373a.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450364.pdf.

Murtagh:2002:EST

[Mur02]  Fionn Murtagh.  Edito-

rial: Statistics and trends. The Computer Journal, 45(3): 259, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/pdf/450259.
pdf.

Murtagh:2003:MRN

[Mur03] Fionn Murtagh. In memoriam: Roger Needham. The Computer Journal, 46(3):227, May 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_03/pdf/460227.
                                        pdf.

Murtagh:2005:E

[Mur05] Fionn Murtagh. Editorial. The Computer Journal, 48(4):381, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/reprint/48/4/381.

Murtagh:2006:E

[Mur06]  Fionn Murtagh.  Edito-

rial. The Computer Journal, 49(2):134–135, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/49/
                                        2/134;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/49/2/134.

Murtagh:2008:E

[Mur08]  Fionn Murtagh.  Editorial.

The Computer Journal, 51 (6):612–614, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/51/
                                        6/612;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/51/6/612.

Moussiades:2005:PCA

[MV05] Lefteris Moussiades and Athena Vakali. PDetect: a clustering approach for detecting plagiarism in source code datasets. The Computer Journal, 48(6):651–661, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/6/651;   http://comjnl.
oxfordjournals.org/cgi/content/
full/48/6/651;   http://
comjnl.oxfordjournals.org/
cgi/reprint/48/6/651.

Nixon:2002:FEI

[NA02] Mark S. Nixon and Alberto S.

Aguado. Feature extraction and image processing. Newnes, Oxford, UK, 2002. ISBN 0-7506-5078-8. xii + 350 pp. LCCN TA1637 .N58 2002.

Napoli:2002:BR

[Nap02] Margherita Napoli. Book review. The Computer Journal, 45(3):375–376, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_
03/450375a.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_
03/pdf/450364.pdf.

Nguyen:2001:CBT

[Ngu01] Ngoc Thanh Nguyen. Consensusbased timestamps in distributed temporal databases. The Computer Journal, 44(5): 398–409, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_05/440398.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_05/pdf/
440398.pdf.

Ng:2007:MIT

[NLGP07] William Ng, Jack Li, Simon Godsill, and Sze Kim Pang. Multitarget initiation, tracking and termination using Bayesian Monte Carlo methods. The Computer Journal, 50(6):674–693, November 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/6/674;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/6/674;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/6/674.

Noble:2000:AR

[Nob00]  James Noble.  Arguments

and results. The Computer Journal, 43(6):439–450, ???? 2000. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_06/430439.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_06/pdf/
430439.pdf.

Nikmehr:2007:FRF

[NPL07] Hooman Nikmehr, Braden Phillips, and Cheng-Chew Lim. A fast radix-4 floating-point divider with quotient digit selection by comparison multiples. The Computer Journal, 50(1):81–92, January 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/1/81;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/1/81;   http://comjnl.
oxfordjournals.org/cgi/reprint/
50/1/81.

Norrish:2002:THD

[NS02] Michael Norrish and Konrad Slind. A thread of HOL development. The Computer Journal, 45(1):37– 45, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_01/450037.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_01/pdf/
450037.pdf.

Older:2002:FMA

[OC02] Susan Older and Shiu-Kai Chin. Formal methods for assuring security of protocols. The Computer Journal, 45(1): 46–54, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_01/450046.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_01/pdf/
                                        450046.pdf.

ONeill:2005:PCL

[OFW^{+}05] Paul O’Neill, Karl Fraser, Zidong Wang, Paul Kellam, Joost N. Kok, and Xiaohui Liu. Pyramidic clustering of large-scale microarray images. The Computer Journal, 48(4):466–479, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/4/466;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        48/4/466.

OKeefe:2002:BR

[O’K02] R. A. O’Keefe. Book review. The Computer Journal, 45(2): 253–254, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_02/450253a.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_02/pdf/
                                        450249.pdf.

OLeary:2005:RBS

[O’L05] Daniel O’Leary. Review: Ontologies: A Silver Bullet for Knowledge Management and Electronic Commerce. The Computer Journal, 48(4):498, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/reprint/48/4/498.

Oke:2007:DSD

[ÖL07] Gülay Öke and Georgios Loukas. A denial of service detector based on maximum likelihood detection and the random neural network. The Computer Journal, 50 (6):717–727, November 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/6/717;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/6/717;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/6/717.

Oommen:2002:EHL

[OR02] B. John Oommen and Luis G.

Rueda. The efficiency of histogram-like techniques for database query optimization. The Computer Journal, 45(5): 494–510, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_05/450494.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_05/pdf/
450494.pdf.

Outram:2002:BR

[Out02] Matthew Outram. Book review. The Computer Journal, 45(4):474, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_04/450474a.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_04/pdf/
                                        450474a.pdf.

Owei:2003:DCQ

[Owe03] Vesper Owei. Development of a conceptual query language: Adopting the usercentered methodology. The Computer Journal, 46(6):602– 624, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_06/460602.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_06/pdf/
                                        460602.pdf.

OReilly:2005:GPT

[OYRW05] Una-May O’Reilly, Tina Yu, Rick Riolo, and Bill Worzel, editors. Genetic programming theory and practice II, volume GPEM 08 of Genetic programming series. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-23253-2. xiv + 320 pp. LCCN QA76.623 .G49 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/fy053/2004058923.html.

Pallis:2006:BRR

[Pal06a] George Pallis. Book review:

Robert J. Glushko and T. Im. McGrath, Document Engineering — Analyzing and Designing Document for Business Informatics and Web Services. The MIT Press (2005). ISBN 0-262-07261-0 . £21.95. 703 pp. Hardbound. The Computer Journal, 49(3):376, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
3/376;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/3/376.

Pallis:2006:BRX

[Pal06b] George Pallis. Book review:

Xiaofang Zhou, Stanley Su, Mike P. Papazoglou, Maria E. Orlowska and Keith G. Jeffery, Web Information Systems Engineering — WISE 2004. Springer-Proceedings Series: Lecture Notes in Computer Science, Vol. 3306 (2004). ISBN 3-540-23894-8. £34.55/ US$117.00. 745 pp. Softbound. The Computer Journal, 49(4):501, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/

                                        4/501;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/49/4/501.

Pan:2005:RBS

[Pan05] Yi Pan. Review: Active Services: Concepts, Architecture and Implementation. The Computer Journal, 48(4):498–499, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/reprint/48/4/498-
                                        a.

Papadopoulos:2002:ETL

[Pap02] Constantinos Papadopoulos. An extended temporal logic for CSCW. The Computer Journal, 45(4):453–472, ???? 2002. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_04/450453.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_04/pdf/
                                        450453.pdf.

Parimala:2002:EOS

[Par02] N. Parimala. Explicit operation specification for component databases. The Computer Journal, 45(2):202–212, ???? 2002. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_02/450202.
                                        sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/
hdb/Volume_45/Issue_02/pdf/
450202.pdf.

Parmar:2005:RBM

[Par05] Minaz Parmar. Review: Distributed Multimedia Database Technologies Supported by MPEG-7 and MPEG-21. The Computer Journal, 48(5): 563–564, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/reprint/48/5/563-
a.

Pickup:2009:BMI

[PCRZ09] Lyndsey C. Pickup, David P.

Capel, Stephen J. Roberts, and Andrew Zisserman. Bayesian methods for image superresolution. The Computer Journal, 52(1):101–113, January 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/1/101;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/1/101;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/1/101.

Poernomo:2005:APP

[PCW05] Iman Hafiz Poernomo, John N.

Crossley, and M. (Martin) Wirsing. Adapting proofs-as-programs: the Curry–Howard protocol. Monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-23759-3 (hardback). xi + 420 pp. LCCN QA9.54 .P64 2005.

Peters:2007:CIP

[Pet07] Jan Peters. Computational intelligence: Principles, techniques and applications. The Computer Journal, 50(6): 758, November 2007. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

                                        tronic). URL http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/6/758;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/6/758.

Pitoura:2001:DLD

[PF01] Evaggelia Pitoura and Ioannis Fudos. Distributed location databases for tracking highly mobile objects. The Computer Journal, 44(2):75– 91, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_44/Issue_02/440075.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_02/pdf/
                                        440075.pdf.

PereiraeSilva:2009:MEU

[PF09] Ricardo Pereira e Silva and Evandro César Freiberger. Metrics to evaluate the use of object oriented frameworks. The Computer Journal, 52(3):288–304, May 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/3/288;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/3/288;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/3/288.

Pinto:2005:DCA

[PFT05] Mónica Pinto, Lidia Fuentes, and José María Troya. A dynamic component and aspectoriented platform. The Computer Journal, 48(4):401–420, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/4/401;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/4/401.

Politopoulos:2001:PCS

[PGT01] Kostas Politopoulos, George F.

Georgakopoulos, and Panagiotis Tsanakas. Precedence constrained scheduling: a case in P. The Computer Journal, 44(3):163– 173, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_03/440163.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_03/pdf/
440163.pdf.

Phillips:2004:BRB

[Phi04]  Nigel Phillips.  Book re-

view: Mobile Communications, by Jochen Schiller. The Computer Journal, 47(5):622, September 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/free_
                                        pdf/470622.pdf;   http://
                                        www3.oup.co.uk/computer_
                                        journal/hdb/Volume_47/Issue_
                                        05/470622.sgm.abs.html.

Pengelly:2000:QMI

[PI00] A. D. Pengelly and D. C. Ince. Quotient machines, the interface equation and protocol conversion. The Computer Journal, 43(1):24– 39, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_43/Issue_01/430024.
                                        sgm.abs.html.

Pierce:2005:ATT

[Pie05] Benjamin C. Pierce, editor. Advanced topics in types and programming languages. MIT Press, Cambridge, MA, USA, 2005. ISBN 0-262-16228-8. xiii + 574 pp. LCCN QA76.7 .A36 2005.

Piper:2005:TLC

[Pip05] Fred Piper. Turing Lecture: Cyberworld security — the good, the bad and the ugly. The Computer Journal, 48(2):

145–156, March 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_48/Issue_02/bxh076.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_48/Issue_02/pdf/
bxh076.pdf.

Pitt:2006:VMA

[PKSA06] Jeremy Pitt, Lloyd Kamara, Marek Sergot, and Alexander Artikis. Voting in multiagent systems. The Computer Journal, 49(2):156–170, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/2/156;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/2/156;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/2/156.

Pan:2002:CTA

[PL02] Tien-Tai Pan and Shun-Shii Lin. Constant-time algorithms for minimum spanning tree and related problems on processor array with reconfigurable bus systems. The Computer Journal, 45(2):174– 186, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_02/450174.
sgm.abs.html;   http://www3.

                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_02/pdf/
                                        450174.pdf.

Papadopoulos:2006:PDJ

[PNM06] Apostolos N. Papadopoulos, Alexandros Nanopoulos, and Yannis Manolopoulos. Processing distance join queries with constraints. The Computer Journal, 49(3):281–296, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        49/3/281;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/49/3/281;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/49/3/281.

Poernomo:2006:BRY

[Poe06] Iman Hafiz Poernomo. Book review: Yves Bertot and Pierre Castéran, Interactive Theorem Proving and Program Development (Coq’Art: The Calculus of Inductive Constructions). Texts in Theoretical Computer Science. An EATCS Series. (2004). ISBN 3-540-20854-2. 469pp. Hardcover. The Computer Journal, 49(1):130–131, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/49/
                                        1/130-a;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        49/1/130-a.

Polkowski:2006:BRM

[Pol06] Lech Polkowski. Book review: Una-May O’Reilly, Tina Yu, Rick Riolo and Bill Worzel (eds), Genetic Programming Theory and Practice II, Springer Science+Business Media, Inc. (2005). ISBN 0-387-23253-2. $129.00. 320 pp. Hardbound. The Computer Journal, 49(1):127–128, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
1/127-a;   http://comjnl.
oxfordjournals.org/cgi/reprint/
49/1/127-a.

Pomerance:2002:BR

[Pom02] Carl Pomerance. Book review.

The Computer Journal, 45(2): 256, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_02/450256.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_02/pdf/
450255.pdf.

Pont:2002:BR

[Pon02] M. J. Pont. Book review. The Computer Journal, 45(6):685– 686, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_06/450685.

                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_06/pdf/
                                        450685.pdf.

Portinale:2003:BRa

[Por03a] Luigi Portinale. Book review. The Computer Journal, 46(3): 332–333, May 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_03/460332b.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_03/pdf/
                                        460332b.pdf.

Portinale:2003:BRb

[Por03b] Luigi Portinale. Book review. The Computer Journal, 46(3): 333–334, May 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_03/460333.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_03/pdf/
                                        460333.pdf.

Pastrana:2008:CSA

[PPK08] J. L. Pastrana, E. Pimentel, and M. Katrib. Composition of self-adapting components for customizable systems. The Computer Journal, 51(4):481–496, July 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/4/481;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/4/481;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/4/481.

Pultorak:2003:MPG

[PQA^{+}03] Dave Pultorak, Pete Quagliariello, Rolf Akker, Jan van Bon, Mike Pieper, and Jan Schaap. MOF pocket guide: IT service operations management. Van Haren Publishing, Zaltbommel, The Netherlands, 2003. ISBN 90-77212-14-0. 96 pp. LCCN ????

Prencipe:2002:BR

[Pre02] Giuseppe Prencipe. Book review. The Computer Journal, 45(5):574–575, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_
05/450574a.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_
05/pdf/450574a.pdf.

Paterno:2003:SRA

[PS03] Fabio Paternò and Carmen Santoro. Support for reasoning about interactive systems through human-computer interaction designers’ representations. The Computer Journal, 46(4):340–357, July 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_04/460340.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_04/pdf/
460340.pdf.

Perkins:2005:RDC

[PS05] S. Perkins and D. H. Smith. Robust data compression: Variable length codes and burst errors. The Computer Journal, 48(3):315–322, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/3/315;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/3/315.

Perkins:2004:RDC

[PSR04] S. Perkins, D. H. Smith, and A. Ryley. Robust data compression: Consistency checking in the synchronization of variable length codes. The Computer Journal, 47(3):309–319, May 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Power:2009:FNR

[PSS09] David Power, Mark Slaymaker, and Andrew Simpson. On formalizing and normalizing role-based access control systems. The Computer Journal, 52(3):305–325, May 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/3/305;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/3/305;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/3/305.

Papatriantafilou:2000:WFH

[PT00] Marina Papatriantafilou and Phillippas Tsigas. Waitfree handshaking using rainbow colouring. The Computer Journal, 43(2):130–137, ???? 2000. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_02/430130.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_02/pdf/
430130.pdf.

Puustjarvi:2001:WCC

[Puu01] Juha Puustjärvi. Workflow concurrency control. The Computer Journal, 44(1):42– 53, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_01/440042.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_01/pdf/
440042.pdf.

Pagnia:2003:FE

[PVG03] Henning Pagnia, Holger Vogt, and Felix C. Gärtner. Fair exchange. The Computer Journal, 46(1):55–75, January 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460055.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460055.pdf.

Pochet:2006:PPM

[PW06] Yves Pochet and Laurence A. Wolsey. Production planning by mixed integer programming. Springer series in operations research and financial engineering. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-29959-9 (hardcover). xxiii + 499 pp. LCCN TS176 .P57 2006; MLCM 2006/40914

(T). URL http://www.loc.
gov/catdir/enhancements/
fy0663/2005935294-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0814/2005935294-
t.html.

Randell:2000:TML

[Ran00] Brian Randell. Turing Memorial Lecture: Facing up to faults. The Computer Journal, 43(2):95–106, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_43/Issue_

02/430095.pdf;   http://
www3.oup.co.uk/computer_
journal/hdb/Volume_43/Issue_
02/430095.sgm.abs.html.

Rashid:2003:EAO

[RB03] Awais Rashid and Lynne Blair.

Editorial: Aspect-oriented programming and separation of crosscutting concerns. The Computer Journal, 46(5):527– 528, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_05/460527.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_05/pdf/
460527.pdf.

Refaat:2009:BRC

[Ref09] Mamdouh Refaat. Book review: Colleen Mccue, Data Mining and Predictive Analytics: Intelligence Gathering and Crime Analysis, Elsevier (2007). ISBN 0-7506-7796-1. 28.99 GBP. 332 pp. Paperback. The Computer Journal, 52(6):725–726, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/reprint/52/6/725-
a.

Reinhard:2004:TAM

[Rei04] Erik Reinhard. Tomas Akenine-Möller and Eric Haines: Real-Time Rendering, 2nd edition.

The Computer Journal, 47(1): 134–??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/free_
                                        pdf/470134.pdf;   http://
                                        www3.oup.co.uk/computer_
                                        journal/hdb/Volume_47/Issue_
                                        01/470134.sgm.abs.html.

reviews:2005:CR

[rev05] Capsule reviews. Capsule reviews. The Computer Journal, 48(2):127–129, March 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/free_pdf/bxh095.
                                        pdf.

Rae:2005:IFA

[RF05] Andrew Rae and Colin Fidge. Information flow analysis for fail-secure devices. The Computer Journal, 48(1):17– 26, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_48/Issue_01/bxh056.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_48/Issue_01/pdf/
                                        bxh056.pdf.

Robinson:2009:ORA

[RFM09] Dirk Robinson, Sina Farsiu, and Peyman Milanfar. Optimal registration of aliased images using variable projection with applications to superresolution. The Computer Journal, 52(1):31–42, January 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/1/31;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/1/31;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/1/31.

Raymond:2002:RCG

[RGW02] John W. Raymond, Eleanor J.

Gardiner, and Peter Willett. RASCAL: Calculation of graph similarity using maximum common edge subgraphs. The Computer Journal, 45(6):631– 644, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_06/450631.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_06/pdf/
450631.pdf.

Rice:2008:IRL

[Ric08] J. E. Rice. An introduction to reversible latches. The Computer Journal, 51 (6):700–709, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/6/700;   http://comjnl.
oxfordjournals.org/cgi/content/

                                        full/51/6/700;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/51/6/700.

Rinderknecht:2003:AVA

[Rin03] Christian Rinderknecht. An algorithm for validating ASN.1 (X.680) specifications using set constraints. The Computer Journal, 46(4):401–420, July 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_04/460401.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_04/pdf/
                                        460401.pdf.

Rissanen:2006:SFD

[Ris06]  J. Rissanen.  The struc-

ture function and distinguishable models of data. The Computer Journal, 49 (6):657–664, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/49/
                                        6/657;   http://comjnl.oxfordjournals.
                                        org/cgi/reprint/49/6/657.

Raskovic:2004:MMA

[RMJ04] Dejan Raskovic, Thomas Martin, and Emil Jovanov. Medical monitoring applications for wearable computing. The Computer Journal, 47(4):495– 504, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_04/470495.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_04/pdf/
470495.pdf.

Rocca:2004:BRA

[Roc04] Simona Ronchi Della Rocca.

Book review: Abstract State Machines — A Method for High-Level System Design and Analysis, by Egon Boerger and Robert Staerk. The Computer Journal, 47(2):270– 271, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470270a.pdf;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
02/470270a.sgm.abs.html.

Rothlauf:2006:RGE

[Rot06] Franz Rothlauf. Representations for genetic and evolutionary algorithms. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2006. ISBN 3-540-25059-X. xvii + 325 pp. LCCN QA76.623 .R68 2006.

Rai:2003:ESM

[RS03] Siddhartha Rai and Priti Shankar. Efficient statistical modeling for the compression of tree structured intermediate code. The Computer Journal, 46(5):476–486, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_05/460476.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_05/pdf/
460476.pdf.

Roberts:2002:LBO

[RSW02] Alan Roberts, Antonios Symvonis, and David R. Wood. Lower bounds for one-to-one packet routing on trees using hot-potato algorithms. The Computer Journal, 45(4):423– 435, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_04/450423.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_04/pdf/
450423.pdf.

Romero:2008:MOC

[RTV08] José Raúl Romero, José M. Troya, and Antonio Vallecillo. Modeling ODP computational specifications using UML. The Computer Journal, 51(4):435–450, July 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/

51/4/435;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/4/435;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/4/435.

Rudiger:2007:QPL

[Rüd07]  Roland Rüdiger.  Quan-

tum programming languages: An introductory overview. The Computer Journal, 50 (2):134–150, March 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/2/134;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/2/134;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/2/134.

Rusu:2006:VAP

[Rus06]  Vlad Rusu.  Verifying an

ATM protocol using a combination of formal techniques. The Computer Journal, 49 (6):710–730, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/6/710;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/6/710;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/6/710.

Rubino:2006:CMQ

[RVB06] Gerardo Rubino, Martín Varela, and Jean-Marie Bonnin. Controlling multimedia QoS in the future home network using the PSQA metric. The Computer Journal, 49(2):137–155, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/2/137;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/2/137;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/2/137.

Reuven:2006:MTS

[RW06] Moses Reuven and Yair Wiseman. Medium-term scheduler as a solution for the thrashing effect. The Computer Journal, 49(3):297–309, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/3/297;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/3/297;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/3/297.

Sofokleous:2009:DSV

[SA09] Anastasis A. Sofokleous and Marios C. Angelides. Dynamic selection of a video content adaptation strategy from a Pareto front. The Computer Journal, 52(4):413–428, July 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.

org/cgi/content/abstract/
52/4/413;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/4/413.

Savsar:2006:RDA

[SAA06] Mehmet Savsar and Fawaz S.

Al-Anzi. Reliability of data allocation on a centralized service configuration with distributed servers. The Computer Journal, 49(3):258–267, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/3/258;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/3/258;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/3/258.

Sanders:2002:RRI

[San02a]  Peter Sanders.  Random-

ized receiver initiated loadbalancing algorithms for treeshaped computations. The Computer Journal, 45(5):561– 573, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_05/450561.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_05/pdf/
450561.pdf.

Sanna:2002:BR

[San02b] Andrea Sanna. Book review.

The Computer Journal, 45(4):

474, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_04/450474b.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_04/pdf/
450474b.pdf.

Sarbazi-Azad:2000:MLH

[SAOKM00] H. Sarbazi-Azad, M. Ould-Khaoua, and L. M. Mackenzie. Message latency in hypercubes in the presence of matrix-transpose traffic. The Computer Journal, 43(5):411– 419, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_05/430411.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_05/pdf/
430411.pdf.

Sapino:2003:BR

[Sap03] M. L. Sapino. Book review. The Computer Journal, 46(1): 116, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460116.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460116.pdf.

Sapir:2004:THF

[Sap04] Amir Sapir. The Tower of Hanoi with forbidden moves. The Computer Journal, 47(1): 20–??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_01/470020.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_01/pdf/
470020.pdf.

Shatnawi:2002:OSD

[SAS02] Ali Shatnawi, M. O. Ahmad, and M. N. S. Swamy. Optimal scheduling of digital signal processing data-flow graphs using shortest-path algorithms. The Computer Journal, 45(1): 88–100, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_01/450088.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_01/pdf/
450088.pdf.

Stewart:2001:BSC

[SC01] A. Stewart and M. Clint.

BSP-style computation: a semantic investigation. The Computer Journal, 44(3):174– 185, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/

                                        Volume_44/Issue_03/440174.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_03/pdf/
                                        440174.pdf.

Szymanski:2008:CTN

[SC08] Boleslaw K. Szymanski and Gilbert G. Chen. Computing with time: From neural networks to sensor networks. The Computer Journal, 51(4):511–522, July 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        51/4/511;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/51/4/511;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/51/4/511.

Schiller:2003:MC

[Sch03a] Jochen H. Schiller. Mobile communications. Addison-Wesley, Reading, MA, USA, second edition, 2003. ISBN 0-321-12381-6 (paperback). xviii + 492 pp. LCCN TK6570.M6 S35 2003.

Schneier:2003:BFT

[Sch03b] Bruce Schneier. Beyond fear: thinking sensibly about security in an uncertain world. Copernicus Books, New York, NY, USA, 2003. ISBN 0-387-02620-

7. 295 pp. LCCN HV6432 .S36

2003.

Schiller:2004:BRM

[Sch04] Jochen Schiller. Book review: Mobile Commerce and Wireless Computing Systems by Nigel Phillips and Geoffrey Elliott. The Computer Journal, 47(2): 272–273, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470272.pdf;   http://
www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
02/470272.sgm.abs.html.

Schneider:2006:BRR

[Sch06] Klaus Schneider. Book review: R. Kurki-Suonio, A Practical Theory of Reactive Systems. Springer-Verlag (2005). ISBN 3-540-23342-3. 49.95/£38.50. 418 pp. Hardbound. The Computer Journal, 49(6):762, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
6/762;   http://comjnl.oxfordjournals.
org/cgi/reprint/49/6/762.

Schroder:2007:BRM

[Sch07] Lutz Schröder. Book review:

L. M. G. Feijs and H. B. M. Jonkers, Formal Specification and Design. Cambridge University Press (2005). ISBN-13 978-0-521-43592-5, ISBN-10 0-521-43592-7. £29.00/$50.00. Softbound. The Computer Journal, 50(1):129–130, January 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/50/
                                        1/129-a;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        50/1/129-a.

Spooner:2005:PAW

[SCJ^{+}05] D. P. Spooner, J. Cao, S. A. Jarvis, L. He, and G. R. Nudd. Performance-aware workflow management for grid computing. The Computer Journal, 48(3):347–357, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/3/347;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        48/3/347.

Seetharaman:2004:BRB

[See04]  Priya Seetharaman.  Book

review: Beyond Fear — Thinking Sensibly About Security in an Uncertain World, by Bruce Schneier. The Computer Journal, 47(3):397, May 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/free_
                                        pdf/470397.pdf;   http://
                                        www3.oup.co.uk/computer_
                                        journal/hdb/Volume_47/Issue_
                                        03/470397.sgm.abs.html.

Segnan:2007:WDM

[Seg07] Marino Segnan. Web data management practices — emerging techniques and technologies. The Computer Journal, 50(2):249–250, March 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/50/
2/249-a;   http://comjnl.
oxfordjournals.org/cgi/reprint/
50/2/249-a.

Sehgal:2004:BRB

[Seh04] Satish K. Sehgal. Book review: Understanding Agent Systems, by Mark d’Inverno and Michael Luck. The Computer Journal, 47(5):623, September 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470623.pdf;   http://
www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
05/470623.sgm.abs.html.

Sereno:2003:BR

[Ser03] Matteo Sereno. Book review.

The Computer Journal, 46(1): 114, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460114.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460114.pdf.

Sethuraman:2005:RBR

[Set05]  J. Sethuraman.  Review:

Network Recovery: Protection and Restoration of Optical, SONET-SDH, IP, and MPLS. The Computer Journal, 48(4):499–500, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/reprint/48/4/499.

Sharafat:2004:FAD

[SF04] Ahmad R. Sharafat and Mehran S. Fallah. A framework for the analysis of denial of service attacks. The Computer Journal, 47(2):179– 192, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_02/470179.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_02/pdf/
470179.pdf.

Sroubek:2009:SRB

[SFC09] Filip Sroubek, Jan Flusser, and Gabriel Cristóbal. Superresolution and blind deconvolution for rational factors with an application to color images. The Computer Journal, 52(1):142–152, January 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/1/142;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/1/142;   http://

comjnl.oxfordjournals.org/
cgi/content/full/bxm098/
DC1;   http://comjnl.oxfordjournals.
org/cgi/reprint/52/1/142.

Schoenberg:2003:IDT

[SFL03] Frederic Paik Schoenberg, Thomas Ferguson, and Cheng Li. Inverting Dirichlet tessellations. The Computer Journal, 46(1):76–83, January 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460076.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460076.pdf.

Sierra:2006:DOP

[SFVFM06] José Luis Sierra, Alfredo Fernández-Valmayor, and Baltasar Fernández-Manjón. A document-oriented paradigm for the construction of contentintensive applications. The Computer Journal, 49(5): 562–584, September 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/5/562;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/5/562;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/5/562.

Song:2000:GAV

[SG00] Suk-Kyu Song and Narasimhaiah Gorla. A genetic algorithm for vertical fragmentation and access path selection. The Computer Journal, 43(1): 81–93, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_43/Issue_01/430081.
                                        sgm.abs.html.

Scoble:2006:NCH

[SI06] Robert Scoble and Shel Israel. Naked conversations: how blogs are changing the way businesses talk with customers. Wiley, New York, NY, USA, 2006. ISBN 0-471-74719-X (cloth). x + 251 pp. LCCN TK5105.8884 .S26

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0623/2005027678-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0740/2005027678-
                                        b.html;   http://www.loc.
                                        gov/catdir/toc/ecip061/2005027678.
                                        html.

Skyt:2002:PVM

[SJ02] Janne Skyt and Christian S. Jensen. Persistent views

— A mechanism for managing ageing data. The Computer Journal, 45(5):481– 493, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/

Volume_45/Issue_05/450481.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_05/pdf/
450481.pdf.

Spink:2004:WSP

[SJ04] Amanda Spink and Bernard J.

Jansen. Web search: public searching on the Web, volume 6 of Information science and knowledge management. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2004. ISBN 1-4020-2268-9 (hardback), 1-4020-2269-7 (ebook). xiii + 198 pp. LCCN ZA4230 .S65 2004.

Scott:2005:GBP

[SJ05] Elizabeth Scott and Adrian Johnstone. Generalized bottom up parsers with reduced stack activity. The Computer Journal, 48(5):565–587, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/5/565;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/5/565.

Sihman:2003:SAO

[SK03] Marcelo Sihman and Shmuel Katz. Superimpositions and aspect-oriented programming. The Computer Journal, 46 (5):529–541, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_46/Issue_
                                        05/460529.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_46/Issue_
                                        05/pdf/460529.pdf.

Siltaneva:2002:CRB

[SM02] Jarmo Siltaneva and Erkki Mäkinen. A comparison of random binary tree generators. The Computer Journal, 45(6): 653–660, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_06/450653.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_06/pdf/
                                        450653.pdf.

Smith:2002:BR

[Smi02] Edward Smith. Book review. The Computer Journal, 45(2): 252–253, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_02/450252a.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_02/pdf/
                                        450249.pdf.

Smith:2006:BRM

[Smi06] Marc L. Smith. Book review: Mark Burgin, Super-Recursive Algorithms. Series: Monographs in Computer Science. Springer-Verlag (2005). ISBN 0-387-95569-0. $79.95/ £61.50. 304 pp. Hardbound. The Computer Journal, 49 (6):762–763, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
6/762-a;   http://comjnl.
oxfordjournals.org/cgi/reprint/
49/6/762-a.

Snedaker:2006:SIS

[Sne06] Susan Snedaker. Syngress IT Security Project Management Handbook. Syngress Publishing, Inc., Rockland, MA, USA, 2006. ISBN 1-59749-076-8. 612 (est.) pp. LCCN ????

Shen:2009:SRR

[SNLZ09] Huanfeng Shen, Michael K.

Ng, Pingxiang Li, and Liangpei Zhang. Super-resolution reconstruction algorithm to MODIS remote sensing images. The Computer Journal, 52(1):90–100, January 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/1/90;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/1/90;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/1/90.

Sofokleous:2005:RBM

[Sof05]  Anastasis Sofokleous.  Re-

view: H.264 and MPEG-4 Video Compression: Video Coding for Next-generation Multimedia. The Computer Journal, 48(5):563, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/reprint/48/5/563.

Solomonoff:2003:KLU

[Sol03] Ray J. Solomonoff. The Kolmogorov Lecture: The universal distribution and machine learning. The Computer Journal, 46(6):598–601, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_06/460598.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_06/pdf/
460598.pdf.

Solomonoff:2008:TKP

[Sol08] Ray J. Solomonoff. Three kinds of probabilistic induction: Universal distributions and convergence theorems. The Computer Journal, 51 (5):566–570, September 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/5/566;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/5/566;   http://

comjnl.oxfordjournals.org/
cgi/reprint/51/5/566.

Spector:2004:AQC

[Spe04] Lee C. Spector. Automatic quantum computer programming: a genetic programming approach, volume 7 of Genetic programming series. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2004. ISBN 1-4020-7894-3. x + 153 pp. LCCN QA76.889 .S64 2004.

Smart:2007:ATD

[SRS^{+}07] Paul R. Smart, Alistair Russell, Nigel R. Shadbolt, m. c. schraefel, and Leslie A. Carr. AKTiveSA: a technical demonstrator system for enhanced situation awareness. The Computer Journal, 50(6): 703–716, November 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/6/703;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/6/703;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/6/703.

Shapira:2005:PDF

[SS05] Dana Shapira and James A.

Storer. In place differential file compression. The Computer Journal, 48(6): 677–691, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/6/677;   http://comjnl.
oxfordjournals.org/cgi/content/
full/48/6/677;   http://
comjnl.oxfordjournals.org/
cgi/reprint/48/6/677.

Su:2007:MCT

[SSL07] Kaile Su, Abdul Sattar, and Xiangyu Luo. Model checking temporal logics of knowledge via OBDDs. The Computer Journal, 50(4):403–420, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/4/403;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/4/403;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/4/403.

Shutler:2008:ALT

[SSL08] Paul M. E. Shutler, Seok Woon Sim, and Wei Yin Selina Lim. Analysis of linear time sorting algorithms. The Computer Journal, 51(4):451–469, July 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/4/451;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/4/451;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/4/451.

Susilo:2000:NEF

[SSNGS00] Willy Susilo, Rei Safavi-Naini, Marc Gysin, and Jennifer Seberry. A new and efficient failstop signature scheme. The Computer Journal, 43(5):430– 437, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_05/430430.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_05/pdf/
430430.pdf.

Shu:2004:ABP

[SSS04] Lih Chyun Shu, John A.

Stankovic, and Sang H. Son. Achieving bounded and predictable recovery using realtime logging. The Computer Journal, 47(3):373–394, May 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Sloper:2008:OTD

[ST08] Christian Sloper and Jan Arne Telle. An overview of techniques for designing parameterized algorithms. The Computer Journal, 51(1):122–136, January 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/1/122;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/1/122;   http://

                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/51/1/122.

Stakhov:2002:BTP

[Sta02] Alexey Stakhov. Brousentsov’s ternary principle, Bergman’s number system and ternary mirror-symmetrical arithmetic. The Computer Journal, 45(2): 221–236, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_02/450221.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_02/pdf/
                                        450221.pdf.

Saffre:2009:ADO

[STH^{+}09] Fabrice Saffre, Richard Tateson, José Halloy, Mark Shackleton, and Jean Louis Deneubourg. Aggregation dynamics in overlay networks and their implications for self-organized distributed applications. The Computer Journal, 52(4): 397–412, July 2009. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

                                        tronic). URL http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        abstract/52/4/397;   http:
                                        //comjnl.oxfordjournals.
                                        org/cgi/reprint/52/4/397.

Subramani:2005:CFS

[Sub05] K. Subramani. A comprehensive framework for specifying clairvoyance, constraints and periodicity in real-time scheduling. The Computer Journal, 48(3):259–272, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/3/259;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/3/259.

Sangiorgi:2001:PCT

[SW01] Davide Sangiorgi and David Walker. The [pi]-calculus: a theory of mobile processes. Cambridge University Press, Cambridge, [England], 2001. ISBN 0-521-78177-9. xii + 580 pp. LCCN QA76.59.S36 2001; QA76.59.S36.

Syropoulos:2006:FS

[Syr06] Apostolos Syropoulos. Fuzzifying P systems. The Computer Journal, 49(5): 619–628, September 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/5/619;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/5/619;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/5/619.

Shan:2009:GSC

[SZ09] Lijun Shan and Hong Zhu.

Generating structurally complex test cases by data mutation: a case study of testing an automated modelling tool. The Computer Journal, 52(5):571–588, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        52/5/571;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        52/5/571.

Tabaka:2006:CEF

[Tab06] Jean Tabaka. Collaboration explained: facilitation skills for software project leaders. The Agile software development series. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-26877-6 (paperback). xxxvi + 412 pp. LCCN QA76.76.D47 T32 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0519/2005027650.html.

Thimbleby:1998:FMT

[TAC98] Harold Thimbleby, Stuart Anderson, and Paul Cairns. A framework for modelling Trojans and computer virus infection. The Computer Journal, 41(7):445–458, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/Volume_41/Issue_
                                        07/Vol41_07.body.html#AbstractThimbleby.

See comments [Mäk01, TAC01,

Coh01].

Thimbleby:2001:RCF

[TAC01] H. Thimbleby, S. O. Anderson, and P. A. Cairns. Reply to “Comment on ‘A Framework for Modelling Trojans and Computer Virus Infection’ by E. Mäkinen”. The Computer Journal, 44(4):324– 325, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_04/440324.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_04/pdf/

440324.pdf. See [TAC98, Mäk01, Coh01].

Tadjouddine:2008:VOA

[Tad08] E. M. Tadjouddine. Vertexordering algorithms for automatic differentiation of computer codes. The Computer Journal, 51(6):688– 699, November 2008. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/cgi/content/
abstract/51/6/688;   http:
//comjnl.oxfordjournals.
org/cgi/content/full/51/
6/688;   http://comjnl.oxfordjournals.
org/cgi/reprint/51/6/688.

Tambouratzis:2000:CCT

[Tam00] Tatiana Tambouratzis. Counterclustering for training pattern selection. The Computer Journal, 43(3):177–190, ???? 2000. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.

                                        uk/computer_journal/hdb/
                                        Volume_43/Issue_03/430177.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_43/Issue_03/pdf/
                                        430177.pdf.

Tang:2007:SGK

[Tan07] Qiang Tang. On the security of a group key agreement protocol. The Computer Journal, 50(5):589–590, September 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/5/589;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/5/589;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/5/589.

Tartaro:2002:BR

[Tar02] Ing. Felice Tartaro. Book review. The Computer Journal, 45(4):476–477, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_45/Issue_
                                        04/450476c.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_45/Issue_
                                        04/pdf/450476c.pdf.

Tseng:2003:CSR

[TC03] Hsien-Wen Tseng and Chin-Chen Chang. Construction of symmetrical reversible variable length codes using backtracking. The Computer Journal, 46(1):100–105, January 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_01/460100.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_01/pdf/
460100.pdf.

Tseng:2004:BBA

[TC04] Hsien-Wen Tseng and Chin-Chen Chang. A branch-and-bound algorithm for the construction of reversible variable length codes. The Computer Journal, 47(6):701–707, November 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_06/470701.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_06/pdf/
470701.pdf.

Teuhola:2009:TCI

[Teu09]  Jukka Teuhola.  Tourna-

ment coding of integer sequences. The Computer Journal, 52(3):368–377, May 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/3/368;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/3/368;   http://

                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/52/3/368.

Thomasian:2007:ARP

[TFN07] Alexander Thomasian, Gang Fu, and Spencer W. Ng. Analysis of rebuild processing in RAID5 disk arrays. The Computer Journal, 50 (2):217–231, March 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        50/2/217;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/50/2/217;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/50/2/217.

Tian:2005:PCT

[TGBS05] Xinmin Tian, Milind Girkar, Aart Bik, and Hideki Saito. Practical compiler techniques on efficient multithreaded code generation for OpenMP programs. The Computer Journal, 48(5):588–601, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/5/588;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        48/5/588.

Thawani:2007:CAT

[TGSR07] Amit Thawani, Srividya Gopalan, V. Sridhar, and Krithi Ramamritham. Context-aware timely information delivery in mobile environments. The Computer Journal, 50(4):460–472, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/4/460;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/4/460;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/4/460.

Thomasian:2005:ABR

[TH05] Alexander Thomasian and Chunqi Han. Affinity-based routing in zoned mirrored disks. The Computer Journal, 48(3):292–299, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/3/292;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/3/292.

Theodoridis:2003:TBD

[The03]  Yannis Theodoridis.  Ten

benchmark database queries for location-based services. The Computer Journal, 46(6):713– 725, November 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_06/460713.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_06/pdf/
460713.pdf.

Thomas:2005:EGP

[Tho05a] Nigel Thomas. Editorial: Grid performability. The Computer Journal, 48(3):323–324, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/reprint/48/3/323.

Thomasian:2005:CRA

[Tho05b] Alexander Thomasian. Clustered RAID arrays and their access costs. The Computer Journal, 48(6):702–713, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        48/6/702;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/48/6/702;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/48/6/702.

Timmis:2003:BR

[Tim03] Jon Timmis. Book review. The Computer Journal, 46(5):524– 525, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_05/460524a.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_05/pdf/
                                        460524a.pdf.

Tsai:2001:MVH

[TJ01] Jeffrey J. P. Tsai and Eric Y. T. Juan. Modeling and verification of high-assurance properties of safety-critical systems. The Computer Journal, 44(6): 504–530, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_06/440504.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_06/pdf/
440504.pdf.

Tan:2005:MEA

[TKL05] K. C. Tan, E. F. Khor, and T. H. (Tong H.) Lee. Multiobjective evolutionary algorithms and applications. Advanced information and knowledge processing. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-85233-836-9. x + 295 pp. LCCN QA76.9.A43 T36 2005.

Tseng:2004:LTW

[TKLH04] Yu-Chee Tseng, Sheng-Po Kuo, Hung-Wei Lee, and Chi-Fu Huang. Location tracking in a wireless sensor network by mobile agents and its data fusion strategies. The Computer Journal, 47(4):448– 460, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_04/470448.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/

                                        hdb/Volume_47/Issue_04/pdf/
                                        470448.pdf.

Tousidou:2000:IMS

[TNM00] Eleni Tousidou, Alex Nanopoulos, and Yannis Manolopoulos. Improved methods for signature-tree construction. The Computer Journal, 43(4): 301–314, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_43/Issue_04/430301.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_43/Issue_04/pdf/
                                        430301.pdf.

Tofts:2002:TMP

[Tof02] C. Tofts. Traffic modelling and probabilistic process abstraction. The Computer Journal, 45 (1):62–67, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_01/450062.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_01/pdf/
                                        450062.pdf.

Torres:2005:MDE

[TS05] Manuel Torres and José Samos. A methodology to define external schemas in ODMG databases. The Computer Journal, 48(6):714–736, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/6/714;   http://comjnl.
oxfordjournals.org/cgi/content/
full/48/6/714;   http://
comjnl.oxfordjournals.org/
cgi/reprint/48/6/714.

Tseng:2005:RMP

[Tse05] Yuh-Min Tseng. A robust multi-party key agreement protocol resistant to malicious participants. The Computer Journal, 48(4):480–487, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/4/480;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/4/480.

Tseng:2007:SAG

[Tse07] Yuh-Min Tseng. A secure authenticated group key agreement protocol for resourcelimited mobile devices. The Computer Journal, 50(1):41– 52, January 2007. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/cgi/content/
abstract/50/1/41;   http:
//comjnl.oxfordjournals.
org/cgi/content/full/50/
1/41;   http://comjnl.oxfordjournals.
org/cgi/reprint/50/1/41.

Tzouramanis:2000:OLQ

[TVM00] Theodoros Tzouramanis, Michael Vassilakopoulos, and Yannis Manolopoulos. Overlapping linear quadtrees and spatiotemporal query processing. The Computer Journal, 43(4):325– 324, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_04/430325.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_04/pdf/
430325.pdf.

Tucker:2002:BR

[TVT02] J. V. Tucker, Sergei Vorobyov, and Chris Tofts. Book reviews. The Computer Journal, 45(1): 68–71, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_01/450068.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_01/pdf/
450068.pdf.

Tian:2008:DRS

[TWN08] Ye Tian, Di Wu, and Kam-Wing Ng. On distributed rating systems for peer-to-peer networks. The Computer Journal, 51(2):162–180, March 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/2/162;   http://comjnl.

oxfordjournals.org/cgi/content/
full/51/2/162;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/2/162.

Thomasian:2008:RPM

[TX08] Alexander Thomasian and Jun Xu. Reliability and performance of mirrored disk organizations. The Computer Journal, 51(6):615–629, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/6/615;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/6/615;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/6/615.

Tong:2004:ERL

[TYH04] Sheau-Ru Tong, Yuan-Tse Yu, and Chung-Ming Huang. Efficient region layouts for regionbased video servers employing multi-zone disks. The Computer Journal, 47(2):153– 168, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_02/470153.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_02/pdf/
470153.pdf.

Thomasian:2006:PSD

[TZ06a] Alexander Thomasian and Lijuan Zhang. Persistent semidynamic ordered partition index. The Computer Journal, 49(6):670–684, November 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/6/670;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/6/670;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/6/670.

Thomasian:2006:SDI

[TZ06b] Alexander Thomasian and Lijuan Zhang. The Stepwise Dimensionality Increasing (SDI) index for high-dimensional data. The Computer Journal, 49(5):609–618, September 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/5/609;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/5/609;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/5/609.

Uren:2002:HWC

[UA02] Victoria S. Uren and Thomas R. Addis. How weak categorizers based upon different principles strengthen performance. The Computer Journal, 45(5): 511–524, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/

Volume_45/Issue_05/450511.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_05/pdf/
450511.pdf.

Ungerer:2002:MP

[URŠ02] Theo Ungerer, Borut Robič, and Jurij Šilc. Multithreaded processors. The Computer Journal, 45(3):320– 348, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_03/450320.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_03/pdf/
450320.pdf.

Uzun:2004:BRB

[Uzu04] Isa Servan Uzun. Book review:

Cryptographic Security Architecture — Design and Verification, by Peter Gutmann. The Computer Journal, 47(5):622– 623, September 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470622a.pdf;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
05/470622a.sgm.abs.html.

Vakali:2006:BRB

[Vak06]  Athena Vakali.  Book re-

view: S. Bressan, S. Ceri, E. Hunt, Z. G. Ives, Z. Bellahséne, M. Rys and R. Unland, (EDS) Database and XML Technologies. Third International XML Database Symposium, XSym 2005, Trondheim, Norway, August 28–29, 2005, Proceedings. Series: Lecture Notes in Computer Science, Vol. 3671. Springer-Verlag (2005). ISBN 3-540-28583-0. 44.94/£32.50/US$58.00. 236 pp. Softbound. The Computer Journal, 49(3):376–377, May 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
3/376-a;   http://comjnl.
oxfordjournals.org/cgi/reprint/
49/3/376-a.

Valente:2002:BR

[Val02] Andrea Valente. Book review. The Computer Journal, 45(6): 686–687, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_06/450686a.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_06/pdf/
450686a.pdf.

VandeSompel:2005:AMS

[VBL^{+}05] Herbert Van de Sompel, Jeroen Bekaert, Xiaoming Liu, Luda Balakireva, and Thorsten Schwander. aDORe: a modular, standards-based digital object repository. The Computer Journal, 48(5):

514–535, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/5/514;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/5/514.

Verbeek:2001:DWP

[VBvdA01] H. M. W. Verbeek, T. Basten, and W. M. P. van der Aalst. Diagnosing workflow processes using Woflan. The Computer Journal, 44(4):246– 279, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_04/440246.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_04/pdf/
440246.pdf.

Vandierendonck:2004:GSI

[VD04] Hans Vandierendonck and Koen De Bosschere. On generating set index functions for randomized caches. The Computer Journal, 47(2):245– 258, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_02/470245.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_02/pdf/
470245.pdf.

Vernitski:2006:CUM

[Ver06] Alexei Vernitski. Can unbreakable mean incomputable? The Computer Journal, 49 (1):108–112, January 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        49/1/108;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/49/1/108;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/49/1/108.

Ertvelde:2008:AEC

[VHE08] Luk Van Ertvelde, Filip Hellebaut, and Lieven Eeckhout. Accurate and efficient cache warmup for sampled processor simulation through NSL– BLRL. The Computer Journal, 51(2):192–206, March 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        51/2/192;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/51/2/192;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/51/2/192.

Viroli:2003:TPA

[Vir03] Mirko Viroli. A type-passing approach for the implementation of parametric methods in Java. The Computer Journal, 46(3):263–294, May 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_03/460263.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_03/pdf/
460263.pdf.

vanMoorsel:2005:GMS

[vM05] Aad P. A. van Moorsel.

Grid, management and selfmanagement. The Computer Journal, 48(3):325–332, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/3/325;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/3/325.

Vega:2009:SRM

[VMMK09] M. Vega, J. Mateos, R. Molina, and A. K. Katsaggelos. Superresolution of multispectral images. The Computer Journal, 52(1):153–167, January 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/1/153;   http://comjnl.
oxfordjournals.org/cgi/content/
full/52/1/153;   http://
comjnl.oxfordjournals.org/
cgi/reprint/52/1/153.

Vega-Rodriguez:2004:BRF

[VR04] Miguel A. Vega-Rodriguez.

Book review: Feature Extraction and Image Processing, by Mark S. Nixon and Alberto S. Aguado. The Computer Journal, 47(2):271– 272, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/free_
pdf/470271a.pdf;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_47/Issue_
02/470271a.sgm.abs.html.

vanRooij:2008:PCC

[vRW08] Iris van Rooij and Todd Wareham. Parameterized complexity in cognitive modeling: Foundations, applications and opportunities. The Computer Journal, 51(3):385–404, May 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/3/385;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/3/385;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/3/385.

Verbeek:2007:VWC

[VvdAtH07] H. M. W. Verbeek, W. M. P. van der Aalst, and A. H. M. ter Hofstede. Verifying workflows with cancellation regions and OR-joins: An approach based on relaxed soundness and invariants. The Computer Journal, 50(3):294–314, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/3/294;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/3/294;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/3/294.

vanWezel:2006:PIS

[vWJM06] Wout van Wezel, Rene Jorna, and A. (Alex) Meystel, editors. Planning in intelligent systems: aspects, motivations, and methods. Wiley-Interscience, New York, NY, USA, 2006. ISBN 0-471-73427-6 (hardback). xii + 574 pp. LCCN QA76.76.E95 P533 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0516/2005021353.html.

Ware:2001:IDP

[War01] Willis H. Ware. Introduction to Davies’ paper. The Computer Journal, 44(3):151, ???? 2001. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_03/440151.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_03/pdf/
440151.pdf.

Wang:2004:ENN

[WB04] Hui Wang and David Bell. Extended k-nearest neighbours based on evidence theory. The Computer Journal, 47(6):662– 672, November 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_06/470662.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_06/pdf/
                                        470662.pdf.

Wegener:2003:BR

[Weg03] Ingo Wegener. Book review. The Computer Journal, 46(3): 332, May 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_03/460332.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_03/pdf/
                                        460332.pdf.

Wegner:2004:NMC

[Weg04] Peter Wegner. New models of computation. The Computer Journal, 47(1):4– ??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_01/470004.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_01/pdf/
                                        470004.pdf.

Wang:2004:SSC

[WH04] Chian Wang and Chung-Ming Huang. Synchronization schemes for controlling VCR-like user interactions in interactive Multimedia-on-Demand (MOD) systems. The Computer Journal, 47(2):140– 152, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_02/470140.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_02/pdf/
470140.pdf.

Wu:2003:RCP

[WKwH03] Jun Wu, Tei-Wei Kuo, and Chih wen Hsueh. R.C.P.C.P.: a ceiling-based protocol for multiple-disk environments. The Computer Journal, 46(2): ??, February 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). This paper is an extended version of a paper that appeared in the Proceedings of the 5th International Conference on Real-Time Computing Systems and Applications, 1999.

Wong:2006:EDA

[WLLW06] T. Y. Wong, K. T. Law, John C. S. Lui, and M. H. Wong. An efficient distributed algorithm to identify and traceback DDoS traffic. The Computer Journal, 49(4):418–442, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/

                                        49/4/418;   http://comjnl.
                                        oxfordjournals.org/cgi/content/
                                        full/49/4/418;   http://
                                        comjnl.oxfordjournals.org/
                                        cgi/reprint/49/4/418.

Wang:2001:SFB

[WMS01] Yuke Wang, Carl McCrosky, and Xiaoyu Song. Singlefaced Boolean functions and their minimization. The Computer Journal, 44(4):280– 291, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_44/Issue_04/440280.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_04/pdf/
                                        440280.pdf.

Wheeler:1995:TTE

[WN95] D. J. Wheeler and R. M. Needham. TEA, a tiny encryption algorithm. Lecture Notes in Computer Science, 1008:363– ??, 1995. CODEN LNCSD9. ISSN 0302-9743.

Waldron:2007:ICS

[WPMK07] Samuel M. Waldron, John Patrick, Phillip L. Morgan, and Sophia King. Influencing cognitive strategy by manipulating information access. The Computer Journal, 50(6):694–702, November 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.

org/cgi/content/abstract/
50/6/694;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/6/694;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/6/694.

Welch:2003:RES

[WS03] Ian S. Welch and Robert J.

Stroud. Re-engineering security as a crosscutting concern. The Computer Journal, 46(5):578– 589, September 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_05/460578.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_05/pdf/
460578.pdf.

Wu:2002:MCM

[WTLS02] Shih-Lin Wu, Yu-Chee Tseng, Chih-Yu Lin, and Jang-Ping Sheu. A multi-channel MAC protocol with power control for multi-hop mobile ad hoc networks. The Computer Journal, 45(1):101–110, ???? 2002. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_45/Issue_01/450101.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_45/Issue_01/pdf/
450101.pdf.

Wang:2001:BTL

[WTSS01] San-Yuan Wang, Yu-Chee Tseng, Ching-Sung Shiu, and Jang-Ping Sheu. Balancing traffic load for multinode multicast in a wormhole 2-D torus/mesh. The Computer Journal, 44(5):354– 367, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_44/Issue_05/440354.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_05/pdf/
                                        440354.pdf.

Wu:2002:FTA

[Wu02]  Jie Wu.  A fault-tolerant

adaptive and minimal routing scheme in n-D meshes. The Computer Journal, 45(3):349– 363, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_03/450349.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_03/pdf/
                                        450349.pdf.

Wu:2004:CWW

[Wu04]  Lin-Yi Wu.  Corrigendum:

“WGSN: WLAN-based GPRS Support Node with Push Mechanism”. The Computer Journal, 47(5):624, September 2004. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_05/470624.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_05/pdf/
470624.pdf.

Xia:2006:EPS

[XCWG06] Jinchun Xia, Carl K. Chang, Jeff Wise, and Yujia Ge. An empirical performance study on PSIM. The Computer Journal, 49(5):509–526, September 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/5/509;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/5/509;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/5/509.

Xu:2001:SQP

[XK01] Zhiwei Xu and Taghi M.

Khoshgoftaar. Software quality prediction for high-assurance network telecommunications systems. The Computer Journal, 44(6):557–568, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_44/Issue_
06/440557.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_44/Issue_
06/pdf/440557.pdf.

Xiang:2001:TAC

[XU01] Limin Xiang and Kazuo Ushijima. On O(1) time algorithms for combinatorial generation. The Computer Journal, 44(4): 292–302, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_44/Issue_04/440292.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_44/Issue_04/pdf/
                                        440292.pdf.

Xu:2004:MCF

[Xu04]  Ying Xu.  Model check-

ing for a first-order temporal logic using Multiway Decision Graphs (MDGs). The Computer Journal, 47(1):71– ??, January 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_01/470071.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_01/pdf/
                                        470071.pdf.

Xiang:2000:GRA

[XUA00] Limin Xiang, Kazuo Ushijima, and Selim G. Akl. Generating regular k-ary trees efficiently. The Computer Journal, 43(4): 290–300, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/

Volume_43/Issue_04/430290.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_04/pdf/
430290.pdf.

Xiao:2001:CNU

[XZJ01] Jitian Xiao, Yanchun Zhang, and Xiaohua Jia. Clustering non-uniform-sized spatial objects to reduce I/O cost for spatial-join processing. The Computer Journal, 44(5):384– 397, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_05/440384.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_05/pdf/
440384.pdf.

Yilmaz:2005:DDC

[YE05] Güray Yilmaz and Nadia Erdogan. DCOBE: Distributed composite object-based environment. The Computer Journal, 48(3):273–291, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/3/273;   http://comjnl.
oxfordjournals.org/cgi/reprint/
48/3/273.

Yen:2007:EOS

[Yen07] William Chung-Kung Yen.

Edge-orienting on split, planar and treelike graphs. The Computer Journal, 50(3):357–368, May 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/3/357;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/3/357;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/3/357.

Ye:2006:BBA

[YLK06] Qizhi Ye, Yu Lei, and David Kung. A blocking-based approach to protocol validation. The Computer Journal, 49(5):541–553, September 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
49/5/541;   http://comjnl.
oxfordjournals.org/cgi/content/
full/49/5/541;   http://
comjnl.oxfordjournals.org/
cgi/reprint/49/5/541.

Lam:2000:RCP

[yLKTL00] Kam yiu Lam, Tei-Wei Kuo, Wai-Hung Tsang, and Gary C. K. Law. The reduced ceiling protocol for concurrency control in real-time databases with mixed transactions. The Computer Journal, 43(1):65– 80, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/

Volume_43/Issue_01/430065.
sgm.abs.html.

Yang:2009:SCB

[YLW09] Qian Yang, J. Jenny Li, and David M. Weiss. A survey of coverage-based testing tools. The Computer Journal, 52(5):589–597, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/5/589;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/5/589.

Yamauchi:2009:PFC

[YMB09] Yukiko Yamauchi, Toshimitsu Masuzawa, and Doina Bein. Preserving the faultcontainment of ring protocols executed on trees. The Computer Journal, 52(4):483–498, July 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/4/483;   http://comjnl.
oxfordjournals.org/cgi/reprint/
52/4/483.

Younessi:2000:AAC

[You00] Houman Younessi. An algorithm for automatic construction of class diagrams. The Computer Journal, 43(6):482– 490, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.

uk/computer_journal/hdb/
Volume_43/Issue_06/430482.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_06/pdf/
430482.pdf.

Yang:2008:BDS

[YSHP08] Guowu Yang, Xiaoyu Song, William N. N. Hung, and Marek A. Perkowski. Bidirectional synthesis of 4-bit reversible circuits. The Computer Journal, 51(2):207–215, March 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/2/207;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/2/207;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/2/207.

Yanovsky:2007:LTC

[YWB07] Vladimir M. Yanovsky, Israel A. Wagner, and Alfred M. Bruckstein. A lineartime constant-space algorithm for the boundary fill problem. The Computer Journal, 50(4):473–477, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/4/473;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/4/473;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/4/473.

Yang:2008:VCS

[YWC08] Ching-Nung Yang, Chung-Chun Wang, and Tse-Shih Chen. Visual cryptography schemes with reversing. The Computer Journal, 51 (6):710–722, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/6/710;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/6/710;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/6/710.

Yan:2009:APU

[YWW09] Kuo-Qin Yan, Shun-Sheng Wang, and Shu-Ching Wang. The agreement problem in unreliable scale-free networks. The Computer Journal, 52(4): 499–509, July 2009. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/cgi/content/
abstract/52/4/499;   http:
//comjnl.oxfordjournals.
org/cgi/reprint/52/4/499.

Zhenglin:2002:AMP

[ZC02] Liu Zhenglin and Xu Chongyang.

An analytical model for the performance of the DOC-SIS CATV network. The Computer Journal, 45(3):278– 284, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_03/450278.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_03/pdf/
                                        450278.pdf.

Zhu:2009:GES

[ZCHL09] Hong Zhu, S. C. Cheung, Joseph R. Horgan, and J. Jenny Li. Guest editorial to the special issue on automation of software testing. The Computer Journal, 52(5):513, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/reprint/52/5/513.

Zhu:2004:PSC

[ZDL^{+}04] Qiang Zhu, Brian Dunkel, Wing Lau, Suyun Chen, and Berni Schiefer. Piggyback statistics collection for query optimization: Towards a self-maintaining database management system. The Computer Journal, 47(2):221– 244, March 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_47/Issue_02/470221.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_47/Issue_02/pdf/
                                        470221.pdf.

Zeadally:2000:IPQ

[Zea00]  S. Zeadally.  Implemen-

tation and performance of QoS-aware Java applications over ATM networks. The Computer Journal, 43(4):266– 273, ???? 2000. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_43/Issue_04/430266.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_43/Issue_04/pdf/
430266.pdf.

Zhang:2003:SVT

[Zha03] Kang Zhang, editor. Software visualization: from theory to practice, volume SECS 734 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-7448-4. xxvi + 442 pp. LCCN QA76.65 .S565 2003.

Zimmer:2003:BR

[Zim03]  Pascal Zimmer.  Book re-

view. The Computer Journal, 46(4):461, July 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_46/Issue_04/460461.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_46/Issue_04/pdf/
460461.pdf.

Zito:2002:BR

[Zit02] Michele Zito. Book review. The Computer Journal, 45(5): 577–578, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_45/Issue_05/450577a.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_45/Issue_05/pdf/
                                        450577a.pdf.

Zhang:2003:DRV

[ZL03] Chang N. Zhang and Hua Li. Design of reconfigurable VLSI architecture for hybrid arithmetic in GF (2^{m}). The Computer Journal, 46(4):449– 460, July 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_04/460449.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_04/pdf/
                                        460449.pdf.

Zhang:2001:DWC

[ZLB01] Sijing Zhang, E. Stewart Lee, and Alan Burns. Determining the worst-case synchronous message response time in FDDI networks. The Computer Journal, 44(1):31– 41, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/

Volume_44/Issue_01/440031.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_01/pdf/
440031.pdf.

Zhu:2008:CIT

[ZMSH08] Qiang Zhu, Brahim Medjahed, Anshuman Sharma, and Henry Huang. The collective index: a technique for efficient processing of progressive queries. The Computer Journal, 51(6):662–676, November 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
51/6/662;   http://comjnl.
oxfordjournals.org/cgi/content/
full/51/6/662;   http://
comjnl.oxfordjournals.org/
cgi/reprint/51/6/662.

Zhou:2004:WIS

[ZSP^{+}04] Xiaofang Zhou, Stanley Su, Mike P. Papazoglou, Maria E. Orlowska, and Keith G. Jeffery, editors. Web Information Systems — WISE 2004: 5th International Conference on Web Information Systems Engineering, Brisbane, Australia, November 22– 24, 2004: Proceedings, volume 3306 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-23894-

8. ISSN 0302-9743. LCCN TA168 .I5265 2004. URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3306.htm;   http:
                                        //www.springerlink.com/openurl.
                                        asp?genre=issue&issn=0302-
                                        9743&volume=3306;   http:
                                        //www.springerlink.com/openurl.
                                        asp?genre=volume&id=doi:
                                        10.1007/b103344.

Zimmermann:2003:PWS

[ZTP03] Olaf Zimmermann, Mark Tomlinson, and Stefan Peuser. Perspectives on Web services: applying SOAP, WSDL and UDDI to real-world projects. Springer professional computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-00914-

0. xxxii + 648 pp. LCCN

TK5105.88813 .Z54 2003.

Ziavras:2003:VAH

[ZWP03] Sotirios G. Ziavras, Qian Wang, and Paraskevi Papathanasiou. Viable architectures for highperformance computing. The Computer Journal, 46(1):36– 54, January 2003. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.
                                        uk/computer_journal/hdb/
                                        Volume_46/Issue_01/460036.
                                        sgm.abs.html;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_46/Issue_01/pdf/
                                        460036.pdf.

Zuo:2004:SFT

[ZZ04] Zhihong Zuo and Mingtian Zhou. Some further theoretical results about computer viruses. The Computer Journal, 47(6):627–633, November 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_47/Issue_06/470627.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_47/Issue_06/pdf/
470627.pdf.

Zhang:2001:CSG

[ZZC01] Da-Qian Zhang, Kang Zhang, and Jiannong Cao. A contextsensitive graph grammar formalism for the specification of visual languages. The Computer Journal, 44(3):186– 200, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_03/440186.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_03/pdf/
440186.pdf.